### v1

# Random dynamics and invariant measures for a class of non-Newtonian fluids of differential type on 2D and 3D Poincaré domains

Abstract.
In this article, we consider a class of incompressible stochastic third-grade fluids (non-Newtonian fluids) equations on two- as well as three-dimensional Poincaré domains $\mathscr{O}$ (which may be bounded or unbounded). Our aims are to study the well-posedness and asymptotic analysis for the solutions of the underlying system. Firstly, we prove that the underlying system defined on $\mathscr{O}$ has a unique weak solution (in the analytic sense) under Dirichlet boundary condition and it also generates random dynamical system $\Psi$. Secondly, we consider the underlying system on bounded domains. Using the compact Sobolev embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$, we prove the existence of a unique random attractor for the underlying system on bounded domains with external forcing in $\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$. Thirdly, we consider the underlying system on unbounded Poincaré domains with external forcing in $\mathbb{L}^{2}(\mathscr{O})$ and show the existence of a unique random attractor. In order to obtain the existence of a unique random attractor on unbounded domains, due to the lack of compact Sobolev embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$, we use the uniform-tail estimates method which helps us to demonstrate the asymptotic compactness of $\Psi$. Note that due to the presence of several nonlinear terms in the underlying system, we are not able to use the energy equality method to obtain the asymptotic compactness of $\Psi$ in unbounded domains, which makes the analysis of this work in unbounded domains more difficult and interesting. Finally, as a consequence of the existence of random attractors, we address the existence of invariant measures for underlying system. To the best of authors’ knowledge, this is the first work which consider the 2D as well as 3D incompressible stochastic third-grade fluids equations and establish the existence of random attractor in bounded as well as unbounded domains. In addition, this is the first work which address the existence of invariant measures for underlying system on unbounded domains.
Keywords: Stochastic third-grade fluids, unbounded domains, random attractors, asymptotic compactness, uniform-tail estimates, invariant measures.
MSC 2020: 76A05, 35R60, 35Q35, 37L30, 35B40.
1. Introduction
In this work, we are concerned with the well-posedness and the existence of random attractors for a stochastic version of a class of incompressible non-Newtonian fluids filling a two- or three-dimensional domain (which may be bounded or unbounded) under Dirichlet boundary condition. The majority of research on fluid dynamics has focused on Newtonian fluids, which are defined by a linear relationship between shear stress and strain rate. As a result, these fluids are modeled by the Navier-Stokes equations, which have been the subject of substantial mathematical and physical study. Nonetheless, many actual and industrial fluids exhibit nonlinear viscoelastic behaviour that defies Newton’s equation and, as a result, are not amenable to the description provided by the traditional viscous Newtonian fluid models. These fluids include natural biological fluids such as blood, geological flows and others, see the works [30, 33, 60] and their references. As such, consideration of more broad fluid models is required. Non-Newtonian viscoelastic fluids of differential type have received a lot of attention lately; see for instance [24]. However, in order to comprehend and explain the properties of various nanofluids, a number of simulation studies have been conducted using third-grade fluid models (see for instance [54, 55] and the references therein). Nanofluids are engineered colloidal suspensions of nanoparticles in a base fluid, such as water, ethylene glycol, or oil, and they exhibit enhanced thermal conductivity in comparison to the base fluid, which shows great promise for use in technology and microelectronics. Thus, in order to comprehend the behaviors of these fluids, it is crucial to perform a mathematical analysis of third-grade fluids equations.
1.1. The underlying system
Let us now briefly recall how to obtain the fluids equations for non-Newtonian fluids of differential type, for more details about Kinematics of such fluids we refer to [24]. Denote the velocity field of the fluid by $\boldsymbol{v}$ and introduce the Rivlin-Ericksen kinematic tensors $\mathrm{A}_{n},n\geq 1$, (see [56]), defined by
$\mathrm{A}_{1}(\boldsymbol{v})$ $=\nabla\boldsymbol{v}+(\nabla\boldsymbol{v})^{T};\,\mathrm{A}_{n}(\boldsymbol{v})=\dfrac{\mathrm{d}}{\mathrm{d}t}\mathrm{A}_{n-1}(\boldsymbol{v})+\mathrm{A}_{n-1}(\boldsymbol{v})(\nabla\boldsymbol{v})+(\nabla\boldsymbol{v})^{T}\mathrm{A}_{n-1}(\boldsymbol{v}),\quad n=2,3,\cdots.$
The constitutive law of fluids of grade $n$ reads $\mathbb{T}=-p\mathrm{I}+F(\mathrm{A}_{1},\cdots,\mathrm{A}_{n}),$ where $\mathbb{T}$ is the Cauchy stress tensor, $p$ is the pressure and $F$ is an isotropic polynomial function of degree $n$, subject to the usual requirement of material frame indifference, see [24]. The constitutive law of third-grade fluids $(n=3)$ is given by the following equation
$\mathbb{T}=-p\mathrm{I}+\nu\mathrm{A}_{1}+\alpha_{1}\mathrm{A}_{2}+\alpha_{2}\mathrm{A}_{1}^{2}+\beta_{1}\mathrm{A}_{3}+\beta_{2}(\mathrm{A}_{1}\mathrm{A}_{2}+\mathrm{A}_{2}\mathrm{A}_{1})+\beta_{3}\mathrm{Tr}(\mathrm{A}_{1}^{2})\mathrm{A}_{1},$
where $\nu$ is the viscosity and $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2},\beta_{3}$ are material moduli. We recall that the momentum equations, established by the Newton’s second law, are given by
$\dfrac{\mathrm{D}\boldsymbol{v}}{\mathrm{D}t}=\dfrac{\mathrm{d}\boldsymbol{v}}{\mathrm{d}t}+\boldsymbol{v}\cdot\nabla\boldsymbol{v}=\text{div}(\mathbb{T}).$
If $\beta_{1}=\beta_{2}=\beta_{3}=0$, the constitutive equations correspond to second grade fluids. It has been shown that the Clausius-Duhem inequality and the assumption that the Helmholtz free energy is a minimum in equilibrium requires the viscosity and material moduli to satisfy
$\nu\geq 0,\quad\alpha_{1}+\alpha_{2}=0,\quad\alpha_{1}\geq 0.$ (1.1)
Although second grade fluids are mathematically more treatable, dealing with several non-Newtonian fluids, the rheologists have not confirmed these restrictions (1.1), thus give the conclusion that the fluids that have been tested are not fluids of second grade but are fluids that are characterized by a different constitutive structure, we refer to [33] and references therein for more details. Following [33], in order to allow the motion of the fluid to be compatible with thermodynamic, it should be imposed that
$\nu\geq 0,\quad\alpha_{1}\geq 0,\quad|\alpha_{1}+\alpha_{2}|\leq\sqrt{24\nu\beta},\quad\beta_{1}=\beta_{2}=0,\beta_{3}=\beta\geq 0.$ (1.2)
Consequently, the velocity field $\boldsymbol{v}$ satisfies the incompressible third-grade fluids equations
$\partial_{t}(z(\boldsymbol{v}))-\nu\Delta\boldsymbol{v}+(\boldsymbol{v}\cdot\nabla)z(\boldsymbol{v})&+\displaystyle\sum_{j=1}^{d}z(\boldsymbol{v})^{j}\nabla v^{j}-(\alpha_{1}+\alpha_{2})\text{div}((\mathrm{A}(\boldsymbol{v}))^{2})\\ -\beta\text{div}[\mathrm{Tr}(\mathrm{A}(\boldsymbol{v})\mathrm{A}(\boldsymbol{v})^{T})\mathrm{A}(\boldsymbol{v})]&=-\nabla\mathbf{P}+\boldsymbol{f},\\[5.69046pt] \text{div}\;\boldsymbol{v}&=0,\\[5.69046pt] z(\boldsymbol{v})&:=\boldsymbol{v}-\alpha_{1}\Delta\boldsymbol{v},\\[5.69046pt] \mathrm{A}(\boldsymbol{v})&:=\nabla\boldsymbol{v}+(\nabla\boldsymbol{v})^{T},\end{aligned}\right.$ (1.3)
where the viscosity $\nu$ and the material moduli $\alpha_{1},\alpha_{2}$, $\beta$ verify (1.2), $\mathbf{P}$ denotes the pressure and $\boldsymbol{f}$ denotes an external force. Notice that if $\alpha_{1}=\alpha_{2}=0$ and $\beta$=0, we recover the Navier Stokes equations (NSE). From mathematical point of view, fluids of grade $n$ constitute an hierarchy of fluids with increasing complexity and more nonlinear terms, then comparing with Newtonian (grade $1$) or second grade fluids, third-grade fluids are more complex and require more involved analysis.
Our aim in this work is to study the following stochastic version of system (1.3) with a subset of physical conditions (1.2), namely
$\nu>0,\quad\alpha_{1}=0,\ \alpha_{2}=\alpha,\quad|\alpha|<\sqrt{2\nu\beta},\quad\beta_{1}=\beta_{2}=0,\beta_{3}=\beta>0,$ (1.4)
in the presence of infinite-dimensional Wiener noise:
$\mathrm{d}\boldsymbol{v}&=[\boldsymbol{f}-\nabla\textbf{P}+\nu\Delta\boldsymbol{v}-(\boldsymbol{v}\cdot\nabla)\boldsymbol{v}+\alpha\text{div}((\mathrm{A}(\boldsymbol{v}))^{2})+\beta\text{div}(|\mathrm{A}(\boldsymbol{v})|^{2}\mathrm{A}(\boldsymbol{v}))]\mathrm{d}t&&\\ &\quad+\mathrm{d}\mathrm{W}&&\text{in }\mathscr{O}\times(0,\infty),\\ \text{div}\;\boldsymbol{v}&=0\quad&&\text{in }\mathscr{O}\times[0,\infty),\\ \boldsymbol{v}&=\boldsymbol{0}&&\text{on }\partial\mathscr{O}\times[0,\infty),\\ \boldsymbol{v}(x,0)&=\boldsymbol{x}\quad&&\text{in }\mathscr{O},\end{aligned}\right.$ (1.5)
where $\{\mathrm{W}(t)\}_{t\in\mathbb{R}}$ is a two-sided cylindrical Wiener process in $\mathbb{H}$ (see Subsection 2.1 below for the function spaces) with its Cameron-Martin space/Reproducing Kernel Hilbert Space (RKHS) $\mathrm{K}$ satisfying Hypothesis 1.1 below, defined on some probability space $(\Omega,\mathscr{F},\mathbb{P})$.
Remark 1.1.
Note that (1.4) allow us to define $\varepsilon_{0}:=1-\frac{\alpha^{2}}{2\nu\beta}\in(0,1)$.
1.2. Literature
We first recall some initial works which deals with system (1.3) with $\alpha_{1}>0$. In [2], the authors established the existence of local (in time) solution in $\mathbb{H}^{3}(\mathscr{O})$ to system (1.3) on bounded domains with Dirichlet boundary condition. Later the authors in [15] proved the global (in time) existence in $\mathbb{R}^{d},$ $d\in\{2,3\}$ for $\mathbb{H}^{2}$-valued solution and uniqueness in $\mathbb{R}^{2}$. Note that the uniqueness result in 3D for $\mathbb{H}^{2}$-valued solution is still an open question. In [16], the authors considered system (1.3) with a Navier-slip boundary condition and demonstrated the existence of a global (in time) solution for initial conditions in $\mathbb{H}^{2}(\mathbb{R}^{d})$, $d\in\{2,3\}$. In addition, they also proved that uniqueness holds in 2D. We highlight that the complex nonlinearities in (1.3) make it difficult to develop a solution with less regular initial data, and that in order to obtain any results, one must impose an additional limitation on the parameters $\alpha_{1},\alpha_{2},\beta$, and $\nu$. In fact, the author in [53] demonstrated the existence of a global weak solution for (1.3) in $\mathbb{R}^{d}$, $d\in\{2,3\}$, under some additional restrictions on the parameters, which permits the application of a monotonicity techniques when the initial data belong only to $\mathbb{H}^{1}(\mathbb{R}^{d})$ and $\alpha_{1}>0$ ([17] for the stationary case). Moreover, the author established a weak-strong uniqueness result and the validity of the energy equality.
Now, let us focus on the literature related to system (1.5) (that is, (1.3) with $\alpha_{1}=0$). Firstly, we draw reader’s attention to the groundbreaking work [47], where the author introduced a new system for incompressible viscous fluids with the viscosity depending on the velocity gradient. This in fact has similar nonlinear terms as system (1.5). In [34], the authors proved the global well-posedness to system (1.5) in $\mathbb{R}^{3}$ with divergence-free initial data belongs to $\mathbb{L}^{2}(\mathbb{R}^{3})$ using a monotonicity method and some extra restriction on the parameters $\nu$, $\alpha$ and $\beta$. We refer readers to the works [36, 53], etc. for the analysis to system (1.5). Some of the stochastic versions of system (1.3) (for $\alpha_{1}>$ as well as $\alpha=0$) have been studied in [23, 52, 60], etc.
Examining the asymptotic behavior of dynamical systems is one of the most significant and thorough areas of mechanics and mathematical physics. The concept of attractors is essential to the study of deterministic infinite-dimensional dynamical systems theory (see the books by Temam [62], Robinson [57], Chepyzhov and Vishik [21] and many others). A key topic in the study of random dynamics of SPDEs is obtaining confirmation that an SPDE generates a stochastic flow or random dynamical system (RDS). It is widely known in the literature that RDS can be generated for Itô stochastic ordinary differential equations and a wide family of PDEs with stationary random coefficients (see [3, 45], etc.). Another crucial area of research into the qualitative characteristics of SPDEs is the examination of infinite-dimensional RDS (see the works [9, 27, 25], etc.).
The existence of a random attractor in bounded domains depends on the compactness of the embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$, which facilitates analysis (see [26, 40], etc.). In the case of unbounded domains, however, the embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$ is no longer compact. Consequently, we cannot use the compactness requirement to demonstrate the existence of random attractors. This challenge (in unbounded domains) was addressed in the deterministic case using several approaches; see [1, 58], etc., for the autonomous case, and [19, 20], etc., for the non-autonomous case. A number of authors have also generalized the deterministic case approaches for SPDEs (see, for instance, [5, 12, 66], etc.). This idea has since been used to demonstrate the existence of random attractors for a number of SPDEs, including the stochastic NSE on the 2D unit sphere ([10]), the stochastic $g$-NSE ([32]), the stochastic reaction-diffusion equations ([5]), the 3D stochastic Benjamin-Bona-Mahony equations ([65]), the 2D and 3D stochastic convective Brinkman-Forchheimer equations (CBFE) [38, 37], the 3D globally modified stochastic NSE [43] and references therein.
1.3. Scope of the article, approaches and assumptions
Broadly, this article is divided into three main parts. Our first aim is to establish the existence and uniqueness of weak solutions (in the analytic sense) of the stochastic third-grade fluids equations (1.5). We begin by defining an Ornstein-Uhlenbeck process which takes values in $\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})$. In order to define a suitable Ornstein-Uhlenbeck process, one of the technical difficulties related to the noise is resolved by using the corresponding RKHS. In particular, motivated from [11, 12, 13] for 2D NSE on unbounded Poincaré domains, we assume that the RKHS $\mathrm{K}$ satisfies the following hypothesis:
Hypothesis 1.1.
$\mathrm{K}\subset\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})$ is a Hilbert space such that for some $\delta\in(0,1/2),$
[…]
We refer readers to Section 2 for necessary functional settings.
Remark 1.2.
1. Let us discuss an example which satisfies Hypothesis 1.1. Since $\mathrm{D}(\mathscr{A}^{1+\frac{r}{2}})\hookrightarrow\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})$ for any $r>0$, one can consider $\mathrm{K}\subset\mathrm{D}(\mathscr{A}^{1+\frac{r}{2}})$ for some $r>0$ as an example, and Hypothesis 1.1 can be reformulated in the following way (see [11]). Let us choose $r>0$ and fix it. $\mathrm{K}$ is a Hilbert space such that $\mathrm{K}\subset\mathrm{D}(\mathscr{A}^{1+\frac{r}{2}})$ and, for some $\delta\in(0,1/2)$, the map
[…]
The condition (1.7) also says that the mapping $\mathscr{A}^{-\delta-1-\frac{r}{2}}:\mathrm{K}\to\mathbb{H}$ is Hilbert-Schmidt. If $\mathscr{O}$ is a bounded domain, then $\mathscr{A}^{-s}:\mathbb{H}\to\mathbb{H}$ is Hilbert-Schmidt if and only if $\sum_{j=1}^{\infty}\lambda_{j}^{-2s}<+\infty,$ where $\mathscr{A}e_{j}=\lambda_{j}e_{j}$, $j\in\mathbb{N}$ and $\{e_{j}\}_{j\in\mathbb{N}}$ is an orthogonal basis of $\mathbb{H}$. In bounded domains, it is well known that $\lambda_{j}\sim j^{\frac{2}{d}},$ for large $j$ (growth of eigenvalues) and hence $\mathscr{A}^{-s}$ is Hilbert-Schmidt if and only if $s>\frac{d}{4}.$ In other words, with $\mathrm{K}=\mathrm{D}(\mathscr{A}^{s+1+\frac{r}{2}}),$ the embedding $\mathrm{K}\hookrightarrow\mathrm{D}(\mathscr{A}^{1+\frac{r}{2}})\hookrightarrow\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})$ is $\gamma$-radonifying if and only if $s>\frac{d}{4}.$ Thus, Hypothesis 1.1 is satisfied for any $\delta>0.$ In fact, the condition (1.6) holds if and only if the operator $\mathscr{A}^{-(s+1+\frac{r}{2}+\delta)}:\mathbb{H}\to\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})$ is $\gamma$-radonifying.
2. The requirement of $\delta<\frac{1}{2}$ in Hypothesis 1.1 is necessary because we need (see Subsection 3.2 below) the corresponding Ornstein-Uhlenbeck process to take values in $\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})$.
Making use of the Ornstein-Uhlenbeck process, we define a Doss-Sussman transformation [29, 59] (see (3.17) below) and obtain an equivalent pathwise deterministic system (see system (3.18) below) to system (1.5). Next, we show the existence of a unique weak solution to the equivalent pathwise deterministic system (3.18) via a monotonicity argument and a Minty-Browder technique. The following monotonicity result of the operator $\mathscr{G}(\cdot):=\nu\mathscr{A}\cdot+\mathscr{B}(\cdot+\boldsymbol{z})+\alpha\mathscr{J}(\cdot+\boldsymbol{z})+\beta\mathscr{K}(\cdot+\boldsymbol{z})$ plays a crucial role:
$\mathscr{G}(\boldsymbol{y}_{1})-\mathscr{G}(\boldsymbol{y}_{2}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>+\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})\|_{{4}}^{2}\|\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\|_{2}^{2}$
$\geq\frac{\nu\varepsilon_{0}}{4}\|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{4}\|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|^{4}_{4}\geq 0,$
for any $\boldsymbol{y}_{1},\boldsymbol{y}_{2}\in\mathbb{X}$ and $\boldsymbol{z}\in\mathbb{W}^{1,4}(\mathscr{O})$ (see Lemma 2.2 for complete details). The global solvability of system (3.18) helps to establish the existence and uniqueness of weak solutions to system (1.5). We point out here that there is no earlier paper which discuss stochastic third-grade fluids equations on unbounded domains.
Secondly, we are interested in the large-time behavior of solutions of system (1.5), that is, the existence of random attractors for system (1.5). Similar to the case of NSE, it is a challenging problem to study the long-time behavior of third-grade fluids equations on the whole space $\mathbb{R}^{d}$. Nevertheless, one can try these kinds of analysis on Poincaré domains which may be bounded or unbounded, see for instance [12, 13, 43, 67], etc. By a Poincaré domain, we mean a domain in which the Poincaré inequality is satisfied. More precisely, we consider the following hypothesis on the domain $\mathscr{O}$:
Hypothesis 1.2.
Let $\mathscr{O}$ be an open, connected and unbounded subset of $\mathbb{R}^{3}$, the boundary of which is uniformly of class $\mathrm{C}^{3}$ (see [35]). We assume that, there exists a positive constant $\lambda$ such that the following Poincaré inequality is satisfied:
$\lambda\int_{\mathscr{O}}|\psi(x)|^{2}\mathrm{d}x\leq\int_{\mathscr{O}}|\nabla\psi(x)|^{2}\mathrm{d}x,\ \text{ for all }\ \psi\in\mathbb{H}^{1}_{0}(\mathscr{O}).$ (1.8)
Example 1.3.
A typical example of a unbounded Poincaré domain in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ are $\mathscr{O}=\mathbb{R}\times(-L,L)$ and $\mathscr{O}=\mathbb{R}^{2}\times(-L,L)$ with $L>0$, respectively, see [62, p.306] and [57, p.117].
Remark 1.3.
When $\mathscr{O}$ is a bounded domain, Poincaré inequality is satisfied automatically with $\lambda=\lambda_{1}$, where $\lambda_{1}$ is the first eigenvalue of the Stokes operator defined on bounded domains.
In this work, we apply the abstract theory established in [13] to demonstrate the existence of random attractors for system (1.5). In order to use the abstract result, we first show that system (1.5) generates random dynamical system $\Psi$ (see Section 3). Next, we need to show that $\Psi$ has an random absorbing set and it is asymptotically compact. We show the existence of random absorbing set using the uniform energy estimates. On the other hand, on bounded domains, we use the compact Sobolev embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$ to obtain asymptotic compactness of $\Psi$. In unbounded domains, due to lack of compact Sobolev embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$, the commonly use methods in literature are the following:
  • 1.
    the energy equality method introduced by Ball [4];
  • 2.
    the uniform-tail estimate method introduced by Wang [64].
Note that unlike the Navier-Stokes equations (see [12, 13]), convective Brinkman-Forchheimer equations (see [38, 41], etc.), etc. we are not able to show that the solutions of third-grade fluids equations (1.5) satisfy the weak continuity property with respect to the initial data. Due to this reason, we are not able to use the energy equality method to obtain the asymptotic compactness of $\Psi$ on unbounded domains. Therefore, for unbounded domains, we explore the uniform-tail estimate method to prove the asymptotic compactness of $\Psi$ (Theorem 5.5).
Remark 1.4.
In the context of attractor when $\mathscr{O}$ is a bounded domains, we have considered $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$. But for unbounded domain case, we need to restrict the forcing term to be in a regular space, that is, $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$. If one proves the weak continuity of the solutions with respect to the initial data of the third-grade fluids equation, it will be possible to work with the same regularity as in bounded domains via energy equality method (see the works [13, 38], etc. for Navier-Stokes equations and related models). Unfortunately, we are not able to prove the weak continuity of the solutions of third-grade fluids equations with respect to the initial data, so the method of energy equality is not applied.
Essentially unlike the parabolic or hyperbolic equations as considered in [18, 48, 64, 68], etc. the fluids equations like (1.5) contain the pressure term P. When we derive these uniform tail-estimates, the pressure term P can not be simply treated by the divergence theorem (does not vanish). However, by taking the divergence in (1.5) and using the incompressibility condition, we get the rigorous expression of the pressure term
P $=(-\Delta)^{-1}\nabla\cdot\left[\nabla\cdot(\boldsymbol{v}\otimes\boldsymbol{v})\right]+\alpha(-\Delta)^{-1}\nabla\cdot\left[\nabla\cdot\left\{(\mathrm{A}(\boldsymbol{v}))^{2}\right\}\right]$
$\quad+\beta(-\Delta)^{-1}\nabla\cdot\left[\nabla\cdot\left\{|\mathrm{A}(\boldsymbol{v})|^{2}\mathrm{A}(\boldsymbol{v})\right\}\right]-(-\Delta)^{-1}\left[\nabla\cdot\boldsymbol{f}\right].$
Then it is possible to derive uniform-tail estimates by estimating all nonlinear terms in system (1.5) and pressure P carefully, see Lemma 5.4 below. As a result of uniform-tail estimates, the asymptotic compactness of $\Psi$ on unbounded domains follows (see Theorem 5.5).
Our final aim is to demonstrate the existence of an invariant measure for the SGMNS equations (1.5). Crauel et. al. in [27] demonstrated that a sufficient condition for the existence of invariant measures is the existence of a compact invariant random set. They implemented this idea to demonstrate the existence of invariant measures in bounded domains for 2D stochastic NSE and reaction-diffusion equations. Moreover, in unbounded domains, this idea was also used to establish the existence of invariant measures for 2D stochastic NSE by Brzézniak et. al. ([12]) and Kinra et. al. ([42]), for 2D and 3D stochastic CBFE by Kinra et. al. ([39, 38]), for stochastic globally modified NSE by Kinra at. al. [43]. Since we show the existence of a random attractor, the existence of an invariant measures is guaranteed as the random attractor itself is a compact invariant set.
Let us now state our main results whose proofs follow from Theorems 3.12, 4.3, 5.5 and 6.2.
Theorem 1.4.
Under (1.4) and Hypotheses 1.1-1.2,
  1. (i)
    there exists a unique weak solution (in the analytic sense) for the stochastic third-grade fluids equations (1.5) with external forcing $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$,
  2. (ii)
    on bounded domains, there exists a unique random attractor for the stochastic third-grade fluids equations (1.5) with external forcing $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$,
  3. (iii)
    on unbounded domains, there exists a unique random attractor for the stochastic third-grade fluids equations (1.5) with external forcing $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$,
  4. (iv)
    on bounded and unbounded domains, there exists an invariant measure for the stochastic third-grade fluids equations (1.5) with external forcing $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$, respectively.
We remark that to the best of authors’ knowledge, the well-posedness and the large-time behavior of the solutions to the stochastic third-grade fluids equations have not been investigated in bounded/unbounded domains till the date. In addition, the existence of invariant measures for the stochastic third-grade fluids equations on unbounded domains is being studied here for the first time. Unfortunately, we are unable to demonstrate the uniqueness of invariant measures, which is a subject we plan to tackle in the future.
1.4. Organization of the article
The organization of this article is as follows: In the next section, we provide necessary function spaces, some important inequalities, linear and nonlinear operators with their properties for abstract formulation of system (1.5) and further analysis and abstract formulation of system (1.5). The metric dynamical system (MDS) and RDS corresponding to stochastic third-grade fluids equations are constructed in Section 3 using a random transformation (known as Doss-Sussman transformation, [29, 59]). The existence and uniqueness of a weak solution satisfying the energy equality to the transformed stochastic third-grade fluids equations (see system (3.19) below) by using monotonicity arguments and a Minty-Browder technique has also been established (Theorem 3.8) in the same section. Section 4 provides the existence of a unique random attractor for stochastic third-grade fluids equations on bounded domains via compact Sobolev embeddings and abstract theory developed in [13] (see Lemma 4.2 and Theorem 4.3). In Section 5, we consider the stochastic third-grade fluids equations on unbounded domains. In order to do this, we first present Lemmas 5.1 and 5.2, which provides us the energy estimates for the solution of system (3.19) and helps us to prove the existence of random absorbing set. Then, we present Lemma 5.3 which is one of the key lemma to prove the $\mathfrak{DK}$-asymptotic compactness of RDS. Next we prove the uniform-tail estimates for the solution of system (3.19) which is another key result to prove the $\mathfrak{DK}$-asymptotic compactness of RDS (see Lemma 5.4). Finally, we establish the main result of this section, that is, the existence of unique random attractor for system (1.5) on unbounded domains using Lemmas 5.2, 5.3 and 5.4, and abstract theory developed in [13], see Theorem 5.5. In the final and smallest section, we address the existence of invariant measures for system (1.5) as a consequence of the existence of random attractor (see Theorem 6.2)
2. Mathematical formulation and some auxiliary results
In this section, we first provide necessary function spaces, inequalities and notations to the analysis of this article. Then, we define linear and nonlinear operators with their properties. Finally, using the operators, we provide the abstract formulation of system (1.5) and the definition of weak solution in the analytic sense.
2.1. Notations and the functional setting
Let $T>0$, for a Banach space $E$, we define
$(E)^{k}:=\{(f_{1},\cdots,f_{k}):f_{l}\in E,\quad l=1,\cdots,k\}\;\text{ for positive integer }k.$
Let $m\in\mathbb{N}^{*}$ and $1\leq p<\infty$, we denote by $\mathbb{W}^{m,p}(\mathscr{O})$ (resp. $\mathrm{W}^{m,p}(\mathscr{O})$) the standard Sobolev space of matrix/vector-valued (resp. scalar-valued) functions whose weak derivative up to order $m$ belong to the Lebesgue space $\mathbb{L}^{p}(\mathscr{O})$ (resp. $\mathrm{L}^{p}(\mathscr{O})$) and set $\mathbb{H}^{m}(\mathscr{O})=\mathbb{W}^{m,2}(\mathscr{O})$ and $\mathbb{H}^{0}(\mathscr{O})=\mathbb{L}^{2}(\mathscr{O})$.
Let us introduce the following spaces:
$\begin{array}[]{ll}\mathscr{V}&:=\{\boldsymbol{v}\in(\mathrm{C}^{\infty}_{c}(\mathscr{O}))^{d}\,:\text{ div }\boldsymbol{v}=0\},\\[2.84526pt] \mathbb{H}&:=\text{The closure of }\mathscr{V}\text{ in }\mathbb{L}^{2}(\mathscr{O})=\{\boldsymbol{v}\in\mathbb{L}^{2}(\mathscr{O})\,:\text{ div }\boldsymbol{v}=0\text{ in }\mathscr{O}\;\;\text{and}\;\;\boldsymbol{v}\cdot\boldsymbol{n}=0\text{ on }\partial\mathscr{O}\},\\[2.84526pt] \mathbb{V}&:=\text{The closure of }\mathscr{V}\text{ in }\mathbb{H}^{1}(\mathscr{O})=\{\boldsymbol{v}\in\mathbb{H}^{1}_{0}(\mathscr{O})\,:\text{ div }\boldsymbol{v}=0\text{ in }\mathscr{O}\}\\ \mathbb{V}_{s}&:=\text{the closure of }\ \mathscr{V}\ \text{ in the Sobolev space }\mathbb{H}^{s}(\mathscr{O}),\ \text{ for }s>1.\end{array}$
Now, we recall the Leray-Helmholtz projector $\mathscr{P}:\mathbb{L}^{2}(\mathscr{O})\to\mathbb{H}$, which is a linear bounded operator characterized by the following $\mathbb{L}^{2}$-orthogonal decomposition $\boldsymbol{v}=\mathscr{P}\boldsymbol{v}+\nabla\varphi,\;\varphi\in\mathrm{H}^{1}(\mathscr{O})$, (cf. [61]).
Now, let us introduce the scalar product between two matrices $A:B=\mathrm{Tr}(AB^{T})$ and denote $|A|^{2}:=A:A.$ The divergence of a matrix $A\in\mathscr{M}_{d\times d}(E)$ is given by $(\text{div}(A)_{i})_{i=1}^{i=d}=(\displaystyle\sum_{j=1}^{d}\partial_{j}a_{ij})_{i=1}^{i=d}.$ The space $\mathbb{H}$ is endowed with the $\mathbb{L}^{2}$-inner product (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) and the associated norm $\|\cdot\|_{2}$. We recall that
$(\boldsymbol{u},\boldsymbol{v})$ $=\sum_{i=1}^{d}\int_{\mathscr{O}}\boldsymbol{u}_{i}\boldsymbol{v}_{i}\mathrm{d}x,\quad\text{ for all }\boldsymbol{u},\boldsymbol{v}\in\mathbb{L}^{2}(\mathscr{O})$
and
$(A,B)$ $=\int_{\mathscr{O}}A:B\mathrm{d}x;\quad\text{ for all }A,B\in\mathscr{M}_{d\times d}(L^{2}(\mathscr{O})).$
On the functional space $\mathbb{V}$, we will consider the following inner product
$\begin{array}[]{ll}(\boldsymbol{u},\boldsymbol{v})_{\mathbb{V}}&:=(\boldsymbol{u},\boldsymbol{v})+(\nabla\boldsymbol{u},\nabla\boldsymbol{v}),\end{array}$
and denote by $\|\cdot\|_{\mathbb{V}}$ the corresponding norm. The usual norms on the classical Lebesgue and Sobolev spaces $\mathrm{L}^{p}(\mathscr{O})$ (resp. $\mathbb{L}^{p}(\mathscr{O})$) and $\mathrm{W}^{m,p}(\mathscr{O})$ (resp. $\mathbb{W}^{m,p}(\mathscr{O})$) will be denoted by $\|\cdot\|_{p}$ and $\|\cdot\|_{\mathrm{W}^{m,p}}$ (resp. $\|\cdot\|_{\mathbb{W}^{m,p}}$), respectively. In addition, given a Banach space $E$, we will denote by $E^{\prime}$ its dual.
Let us introduce the following Banach space $(\mathbb{X},\|\cdot\|_{\mathbb{X}})$
$\mathbb{X}=\mathbb{W}^{1,4}(\mathscr{O})\cap\mathbb{V},\quad\text{ with \;\; }\|\cdot\|_{\mathbb{X}}:=\|\cdot\|_{\mathbb{W}^{1,4}}+\|\cdot\|_{\mathbb{V}}.$
Indeed, we recall that $\mathbb{W}^{1,4}(\mathscr{O})$ endowed with $\|\cdot\|_{\mathbb{W}^{1,4}}$-norm is Banach space, where
$\|\boldsymbol{w}\|_{\mathbb{W}^{1,4}}^{4}=\int_{\mathscr{O}}|\boldsymbol{w}|^{4}\mathrm{d}x+\int_{\mathscr{O}}|\nabla\boldsymbol{w}|^{4}\mathrm{d}x.$
[…]
Let us recall an embedding result from [44, Subsection 2.4].
Lemma 2.1.
For $p>d$, there exists a constant $C=C(p)>0$ such that for all $\boldsymbol{w}\in\mathbb{W}_{0}^{1,p}(\mathscr{O})$
$\|\boldsymbol{w}\|_{\infty}\leq C_{S,3}\|\nabla\boldsymbol{w}\|_{p}.$
Remark 2.1.
Let us provide some inequalities which will be useful in the sequel.
  1. (i)
    Gagliardo-Nirenberg inequality ([51, Theorem 1]) gives (for $d=2,3$)
    $\|\boldsymbol{v}\|_{{4}}\leq C\|\boldsymbol{v}\|^{\frac{4}{4+d}}_{{2}}\|\nabla\boldsymbol{v}\|^{\frac{d}{4+d}}_{{4}},\ \boldsymbol{v}\in\mathbb{L}^{2}(\mathscr{O})\cap\mathbb{W}_{0}^{1,4}(\mathscr{O}).$ (2.1)
  2. (ii)
    Korn-type inequality (see [28, Theorem 2 (ii)] and [46, Theorem 4, p.90]) gives that there exists a constant $C_{K}>0$ such that
    $\|\nabla\boldsymbol{w}\|_{4}\leq C_{K}\|\mathrm{A}(\boldsymbol{w})\|_{4},\quad\text{ for all }\;\boldsymbol{w}\in\mathbb{W}_{0}^{1,4}(\mathscr{O}).$ (2.2)
For the sake of simplicity, we do not distinguish between scalar, vector or matrix-valued notations when it is clear from the context. In particular, $\|\cdot\|_{E}$ should be understood as follows
  • $\|f\|_{E}^{2}=\|f_{1}\|_{E}^{2}+\cdots+\|f_{d}\|_{E}^{2}$ for any $f=(f_{1},\cdots,f_{d})\in(E)^{d}$.
  • $\|f\|_{E}^{2}=\displaystyle\sum_{i,j=1}^{d}\|f_{ij}\|_{E}^{2}$ for any $f\in\mathscr{M}_{d\times d}(E)$.
Throughout the article, we denote by $C$ generic constant, which may vary from line to line.
2.2. Linear and nonlinear operators
Let us define linear operator by
$\mathscr{A}\boldsymbol{v}:=-\mathscr{P}\Delta\boldsymbol{v},\ \boldsymbol{v}\in\mathrm{D}(\mathscr{A}):=\mathbb{V}\cap\mathbb{H}^{2}(\mathscr{O}).$
Remember that the operator $\mathscr{A}$ is a non-negative, self-adjoint operator in $\mathbb{H}$ and
$\mathscr{A}\boldsymbol{v},\boldsymbol{v}\right>=\|\boldsymbol{v}\|_{\mathbb{V}}^{2},\ \textrm{ for all }\ \boldsymbol{v}\in\mathbb{V},\ \text{ so that }\ \|\mathscr{A}\boldsymbol{v}\|_{\mathbb{V}^{\prime}}\leq\|\boldsymbol{v}\|_{\mathbb{V}}.$ (2.3)
Next, we define the trilinear form $b(\cdot,\cdot,\cdot):\mathbb{V}\times\mathbb{V}\times\mathbb{V}\to\mathbb{R}$ by
$b(\boldsymbol{u},\boldsymbol{v},\boldsymbol{w})=\int_{\mathscr{O}}(\boldsymbol{u}(x)\cdot\nabla)\boldsymbol{v}(x)\cdot\boldsymbol{w}(x)\mathrm{d}x=\sum_{i,j=1}^{d}\int_{\mathscr{O}}u_{i}(x)\frac{\partial v_{j}(x)}{\partial x_{i}}w_{j}(x)\mathrm{d}x.$
If $\boldsymbol{u},\boldsymbol{v}$ are such that the linear map $b(\boldsymbol{u},\boldsymbol{v},\cdot)$ is continuous on $\mathbb{V}$, the corresponding element of $\mathbb{V}^{\prime}$ is denoted by $\mathscr{B}(\boldsymbol{u},\boldsymbol{v})$. We represent $\mathscr{B}(\boldsymbol{v})=\mathscr{B}(\boldsymbol{v},\boldsymbol{v})=\mathscr{P}(\boldsymbol{v}\cdot\nabla)\boldsymbol{v}$. Using an integration by parts, it is immediate that
{$b(\boldsymbol{u},\boldsymbol{v},\boldsymbol{v})&=0,\ \text{ for all }\ \boldsymbol{u},\boldsymbol{v}\in\mathbb{V},\\ b(\boldsymbol{u},\boldsymbol{v},\boldsymbol{w})&=-b(\boldsymbol{u},\boldsymbol{w},\boldsymbol{v}),\ \text{ for all }\ \boldsymbol{u},\boldsymbol{v},\boldsymbol{w}\in\mathbb{V}.\end{aligned}\right.$ (2.4)
The trilinear map $b:\mathbb{V}\times\mathbb{V}\times\mathbb{V}\to\mathbb{R}$ has a unique extension to a bounded trilinear map from $\mathbb{L}^{4}(\mathscr{O})\times\mathbb{W}^{1,4}(\mathscr{O})\times\mathbb{H}$ to $\mathbb{R}$. For any $\boldsymbol{u}\in\mathbb{W}^{1,4}(\mathscr{O})$ and $\boldsymbol{v}\in\mathbb{H}$, we have
$|b(\boldsymbol{u},\boldsymbol{u},\boldsymbol{v})|\leq\|\boldsymbol{u}\|_{4}\|\nabla\boldsymbol{u}\|_{4}\|\boldsymbol{v}\|_{2},$
which implies
$\|\mathscr{B}(\boldsymbol{u})\|_{2}\leq\|\boldsymbol{u}\|_{4}\|\nabla\boldsymbol{u}\|_{4}.$
Therefore, $\mathscr{B}$ maps $\mathbb{W}^{1,4}(\mathscr{O})$ into $\mathbb{H}$.
Let us now define the operator $\mathscr{J}(\boldsymbol{v}):=-\mathscr{P}\mathrm{div}(\mathrm{A}(\boldsymbol{v})\mathrm{A}(\boldsymbol{v}))$. Note that for $\boldsymbol{v}\in\mathbb{W}^{1,4}(\mathscr{O})$, we have
$\|\mathscr{J}(\boldsymbol{v})\|_{\mathbb{V}^{\prime}}\leq C\|\mathrm{A}(\boldsymbol{v})\|^{2}_{4},$
and hence the map $\mathscr{J}(\cdot):\mathbb{W}^{1,4}(\mathscr{O})\to\mathbb{V}^{\prime}$ is well-defined.
Finally, we define the operator $\mathscr{K}(\boldsymbol{v}):=-\mathscr{P}\mathrm{div}(|\mathrm{A}(\boldsymbol{v})|^{2}\mathrm{A}(\boldsymbol{v}))$. It is immediate that
$\langle\mathscr{K}(\boldsymbol{v}),\boldsymbol{v}\rangle=\frac{1}{2}\|\mathrm{A}(\boldsymbol{v})\|_{4}^{4}.$
Furthermore, the operator $\mathscr{K}(\cdot):\mathbb{W}^{1,4}(\mathscr{O})\to\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ is well-defined.
Lemma 2.2.
Assume that $\boldsymbol{z}\in\mathbb{W}^{1,4}(\mathscr{O})$ and $\varepsilon_{0}:=1-\sqrt{\frac{\alpha^{2}}{2\beta\nu}}\in(0,1)$. Then, the operator $\mathscr{G}:\mathbb{X}\to\mathbb{X}^{\prime}$ given by
$\mathscr{G}(\cdot):=\nu\mathscr{A}\cdot+\mathscr{B}(\cdot+\boldsymbol{z})+\alpha\mathscr{J}(\cdot+\boldsymbol{z})+\beta\mathscr{K}(\cdot+\boldsymbol{z})$ (2.5)
satisfies
$\mathscr{G}(\boldsymbol{y}_{1})-\mathscr{G}(\boldsymbol{y}_{2}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>+\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})\|_{{4}}^{2}\|\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\|_{2}^{2}$
$\geq\frac{\nu\varepsilon_{0}}{4}\|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{4}\|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|^{4}_{4}\geq 0,$ (2.6)
for any $\boldsymbol{y}_{1},\boldsymbol{y}_{2}\in\mathbb{X}.$
Proof.
Firstly, note that $\varepsilon_{0}=1-\sqrt{\frac{\alpha^{2}}{2\beta\nu}}$ implies $\frac{\alpha^{2}}{4\nu(1-\varepsilon_{0})}=\frac{\beta(1-\varepsilon_{0})}{2}$. We know that
$\nu\mathscr{A}\boldsymbol{y}_{1}-\nu\mathscr{A}\boldsymbol{y}_{2},\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>=\frac{\nu}{2}\|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|_{2}^{2}.$ (2.7)
Now, using (2.4), Hölder’s inequality, Lemma 2.1, (2.2) and Young’s inequality, we estimate
|$\mathscr{B}(\boldsymbol{y}_{1}+\boldsymbol{z})-\mathscr{B}(\boldsymbol{y}_{2}+\boldsymbol{z}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>|$ |
$\leq|b(\boldsymbol{y}_{1}-\boldsymbol{y}_{2},\boldsymbol{y}_{1}-\boldsymbol{y}_{2},\boldsymbol{y}_{2}+\boldsymbol{z})|+|b(\boldsymbol{y}_{2}+\boldsymbol{z},\boldsymbol{y}_{1}-\boldsymbol{y}_{2},\boldsymbol{y}_{1}-\boldsymbol{y}_{2})|$
$\leq 2\|\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\|_{2}\|\nabla(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|_{2}\|\boldsymbol{y}_{2}+\boldsymbol{z}\|_{{\infty}}$
$\leq C_{S,3}C_{K}\|\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\|_{2}\|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|_{2}\|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})\|_{{4}}$
$\leq\frac{\nu\varepsilon_{0}}{4}\|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|_{2}^{2}+\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})\|_{{4}}^{2}\|\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\|_{2}^{2}$
which implies
$\mathscr{B}(\boldsymbol{y}_{1}+\boldsymbol{z})-\mathscr{B}(\boldsymbol{y}_{2}+\boldsymbol{z}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>$
$\geq-\frac{\nu\varepsilon_{0}}{4}\|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|^{2}_{2}-\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})\|_{{4}}^{2}\|\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\|_{2}^{2}.$ (2.8)
From [34, Equation (2.21)], we have
$|\alpha\left<\mathscr{J}(\boldsymbol{y}_{1}+\boldsymbol{z})-\mathscr{J}(\boldsymbol{y}_{2}+\boldsymbol{z}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>|$
$\leq\frac{\nu(1-\varepsilon_{0})}{2}\|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|_{2}^{2}+\frac{\alpha^{2}}{4\nu(1-\varepsilon_{0})}\int_{\mathscr{O}}|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})|^{2}(|\mathrm{A}(\boldsymbol{y}_{1}+\boldsymbol{z})|^{2}+|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})|^{2})$
$\leq\frac{\nu(1-\varepsilon_{0})}{2}\|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|_{2}^{2}+\frac{\beta(1-\varepsilon_{0})}{2}\int_{\mathscr{O}}|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})|^{2}(|\mathrm{A}(\boldsymbol{y}_{1}+\boldsymbol{z})|^{2}+|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})|^{2}),$
which implies
$\alpha\left<\mathscr{J}(\boldsymbol{y}_{1}+\boldsymbol{z})-\mathscr{J}(\boldsymbol{y}_{2}+\boldsymbol{z}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>$
$\geq-\frac{\nu(1-\varepsilon_{0})}{2}\|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})\|_{2}^{2}-\frac{\beta(1-\varepsilon_{0})}{2}\int_{\mathscr{O}}|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})|^{2}(|\mathrm{A}(\boldsymbol{y}_{1}+\boldsymbol{z})|^{2}+|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})|^{2}).$
Now, from [34, Equation (2.13)], we have
$\beta\left<\mathscr{K}(\boldsymbol{y}_{1}+\boldsymbol{z})-\mathscr{K}(\boldsymbol{y}_{2}+\boldsymbol{z}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>$
$=\frac{\beta}{2}\int_{\mathscr{O}}(|\mathrm{A}(\boldsymbol{y}_{1}+\boldsymbol{z})|^{2}-|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})|^{2})^{2}+\frac{\beta}{2}\int_{\mathscr{O}}|\mathrm{A}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})|^{2}(|\mathrm{A}(\boldsymbol{y}_{1}+\boldsymbol{z})|^{2}+|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})|^{2}).$
In view of (2.7) and (2.2), we obtain (2.2). This completes the proof. ∎
Lemma 2.3.
Assume that $\boldsymbol{z}\in\mathbb{W}^{1,4}(\mathscr{O})$. Then, the operator $\mathscr{G}:\mathbb{X}\to\mathbb{X}^{\prime}$ is demicontinuous.
Proof.
Let us take a sequence $\boldsymbol{y}_{m}\to\boldsymbol{y}$ in $\mathbb{X}$, that is, $\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{X}}=\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{V}}+\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{W}^{1,4}}\to 0$ as $m\to\infty$. For any $\boldsymbol{\phi}\in\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\mathbb{X}}$, we consider
$\langle\mathscr{G}(\boldsymbol{y}_{m})-\mathscr{G}(\boldsymbol{y}),\boldsymbol{\phi}\rangle$
$=\nu\langle\mathscr{A}\boldsymbol{y}_{m}-\mathscr{A}\boldsymbol{y},\boldsymbol{\phi}\rangle+\langle\mathscr{B}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{B}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle+\alpha\langle\mathscr{J}(\boldsymbol{y}_{n}+\boldsymbol{z})-\mathscr{J}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle$
$\quad+\beta\langle\mathscr{K}(\boldsymbol{y}_{n}+\boldsymbol{z})-\mathscr{K}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle.$ (2.9)
First, we pick $\langle\mathscr{A}\boldsymbol{y}_{m}-\mathscr{A}\boldsymbol{y},\boldsymbol{\phi}\rangle$ from (2.2) and estimate as follows:
$|\langle\mathscr{A}\boldsymbol{y}_{m}-\mathscr{A}\boldsymbol{y},\boldsymbol{\phi}\rangle|=|(\nabla(\boldsymbol{y}_{m}-\boldsymbol{y}),\nabla\boldsymbol{\phi})|\leq\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{V}}\|\boldsymbol{\phi}\|_{\mathbb{V}}\to 0,\ \text{ as }\ m\to\infty,$ (2.10)
since $\boldsymbol{y}_{n}\to\boldsymbol{y}$ in $\mathbb{V}$. We estimate the term $\langle\mathscr{B}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{B}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle$ from (2.2) using (2.4), Hölder’s inequality and Lemma 2.1 as
$|\langle\mathscr{B}(\boldsymbol{y}_{n}+\boldsymbol{z})-\mathscr{B}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle|$
$\leq|b(\boldsymbol{y}_{m}-\boldsymbol{y},\boldsymbol{\phi},\boldsymbol{y}_{n}-\boldsymbol{y})|+|b(\boldsymbol{y}_{m}-\boldsymbol{y},\boldsymbol{\phi},\boldsymbol{y}+\boldsymbol{z})|+|b(\boldsymbol{y}+\boldsymbol{z},\boldsymbol{\phi},\boldsymbol{y}_{m}-\boldsymbol{y})|$
$\leq\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{2}\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{{\infty}}\|\boldsymbol{\phi}\|_{\mathbb{V}}+\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{2}\|\boldsymbol{y}+\boldsymbol{z}\|_{{\infty}}\|\boldsymbol{\phi}\|_{\mathbb{V}}+\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{2}\|\boldsymbol{y}+\boldsymbol{z}\|_{{\infty}}\|\boldsymbol{\phi}\|_{\mathbb{V}}$
$\leq C\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{2}\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{W}^{1,4}}\|\boldsymbol{\phi}\|_{\mathbb{V}}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}C}\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{2}\|\boldsymbol{y}+\boldsymbol{z}\|_{\mathbb{W}^{1,4}}\|\boldsymbol{\phi}\|_{\mathbb{V}}$
$\to 0,\ \text{ as }\ m\to\infty,$ (2.11)
since $\boldsymbol{y}_{m}\to\boldsymbol{y}$ in $\mathbb{X}$. We estimate the term $\langle\mathscr{J}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{J}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle$ from (2.2) using Hölder’s inequality as
$|\alpha\langle\mathscr{J}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{J}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{y}+\boldsymbol{z}}),\boldsymbol{\phi}\rangle|$
$=|\alpha||\langle\mathrm{A}^{2}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathrm{A}^{2}(\boldsymbol{y}+\boldsymbol{z}),\nabla\boldsymbol{\phi}\rangle|$
$\leq|\alpha||\langle\mathrm{A}^{2}(\boldsymbol{y}_{m}-\boldsymbol{y}),\nabla\boldsymbol{\phi}\rangle|+|\alpha||\langle\mathrm{A}(\boldsymbol{y}_{m}-\boldsymbol{y})\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}),\nabla\boldsymbol{\phi}\rangle|+|\alpha||\langle\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\mathrm{A}(\boldsymbol{y}_{m}-\boldsymbol{y}),\nabla\boldsymbol{\phi}\rangle|$
$\leq|\alpha|\|\boldsymbol{y}_{m}-\boldsymbol{y}\|^{2}_{\mathbb{W}^{1,4}}\|\boldsymbol{\phi}\|_{\mathbb{V}}+2|\alpha|\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{W}^{1,4}}\|\boldsymbol{y}+\boldsymbol{z}\|_{\mathbb{W}^{1,4}}\|\boldsymbol{\phi}\|_{\mathbb{V}}$
$\to 0,\ \text{ as }\ m\to\infty,$ (2.12)
since $\boldsymbol{y}_{m}\to\boldsymbol{y}$ in $\mathbb{W}^{1,4}(\mathscr{O})$. We estimate the term $\langle\mathscr{K}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{K}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle$ from (2.2) using Hölder’s inequality as
$|\beta\langle\mathscr{K}(\boldsymbol{y}_{m}+\boldsymbol{z})-\mathscr{K}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{\phi}\rangle|$
$=\beta|\langle|\mathrm{A}(\boldsymbol{y}_{m}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}_{m}+\boldsymbol{z})-|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}),\nabla\boldsymbol{\phi}\rangle|$
$\leq\beta|\langle|\mathrm{A}(\boldsymbol{y}_{m}-\boldsymbol{y})|^{2}\mathrm{A}(\boldsymbol{y}_{m}-\boldsymbol{y}),\nabla\boldsymbol{\phi}\rangle|+\beta|\langle|\mathrm{A}(\boldsymbol{y}_{m}-\boldsymbol{y})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}),\nabla\boldsymbol{\phi}\rangle|$
$\quad+2\beta|\langle[\mathrm{A}(\boldsymbol{y}_{m}-\boldsymbol{y}):\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})]\mathrm{A}(\boldsymbol{y}_{m}-\boldsymbol{y}),\nabla\boldsymbol{\phi}\rangle|+2\beta|\langle[\mathrm{A}(\boldsymbol{y}_{m}-\boldsymbol{y}):\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})]\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}),\nabla\boldsymbol{\phi}\rangle|$
$\quad+\beta|\langle|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}_{m}-\boldsymbol{y}),\nabla\boldsymbol{\phi}\rangle|$
$\leq\beta\|\boldsymbol{y}_{m}-\boldsymbol{y}\|^{3}_{\mathbb{W}^{1,4}}\|\boldsymbol{\phi}\|_{\mathbb{W}^{1,4}}+3\beta\|\boldsymbol{y}_{m}-\boldsymbol{y}\|^{2}_{\mathbb{W}^{1,4}}\|\boldsymbol{y}+\boldsymbol{z}\|_{\mathbb{W}^{1,4}}\|\boldsymbol{\phi}\|_{\mathbb{W}^{1,4}}$
$\quad+3\beta\|\boldsymbol{y}_{m}-\boldsymbol{y}\|_{\mathbb{W}^{1,4}}\|\boldsymbol{y}+\boldsymbol{z}\|^{2}_{\mathbb{W}^{1,4}}\|\boldsymbol{\phi}\|_{\mathbb{W}^{1,4}}$
$\to 0,\ \text{ as }\ m\to\infty,$ (2.13)
since $\boldsymbol{y}_{m}\to\boldsymbol{y}$ in $\mathbb{W}^{1,4}(\mathscr{O})$.
From the above convergences, it is immediate that $\langle\mathscr{G}(\boldsymbol{y}_{m})-\mathscr{G}(\boldsymbol{y}),\boldsymbol{\phi}\rangle\to 0$, for all $\boldsymbol{\phi}\in\mathbb{X}$. Hence the operator $\mathscr{G}:\mathbb{X}\to\mathbb{X}^{\prime}$ is demicontinuous, which implies that, in addition, the operator $\mathscr{G}(\cdot)$ is hemicontinuous also. ∎
Finally, we provide the result of locally Lipschitz continuity of the operator $\mathscr{G}(\cdot)$.
Lemma 2.4.
Assume that $\boldsymbol{z}\in\mathbb{W}^{1,4}(\mathscr{O})$. Then, the operator $\mathscr{G}:\mathbb{X}\to\mathbb{X}^{\prime}$ is locally Lipschitz.
Proof.
For $\boldsymbol{u},\boldsymbol{v},\boldsymbol{w}\in\mathbb{X}$, we have
$\mathscr{A}\boldsymbol{u}-\mathscr{A}\boldsymbol{v},\boldsymbol{w}\right>$ $=(\nabla(\boldsymbol{u}-\boldsymbol{v}),\nabla\boldsymbol{w})$
$\leq\|\boldsymbol{u}-\boldsymbol{v}\|_{\mathbb{V}}\|\boldsymbol{w}\|_{\mathbb{V}}$ (2.14)
$\langle\mathscr{B}(\boldsymbol{u}+\boldsymbol{z})-\mathscr{B}(\boldsymbol{v}+\boldsymbol{z}),\boldsymbol{w}\rangle$ $=b(\boldsymbol{u}+\boldsymbol{z},\boldsymbol{u}-\boldsymbol{v},\boldsymbol{w})+b(\boldsymbol{u}-\boldsymbol{v},\boldsymbol{v}+\boldsymbol{z},\boldsymbol{w})$
$\leq\big{[}\|\boldsymbol{u}+\boldsymbol{z}\|_{4}\|\nabla(\boldsymbol{u}-\boldsymbol{v})\|_{4}+\|\boldsymbol{u}-\boldsymbol{v}\|_{4}\|\nabla(\boldsymbol{v}+\boldsymbol{z})\|_{4}\big{]}\|\boldsymbol{w}\|_{2}$
$\leq C\big{[}\|\boldsymbol{u}+\boldsymbol{z}\|_{4}\|\mathrm{A}(\boldsymbol{u}-\boldsymbol{v})\|_{4}+\|\boldsymbol{u}-\boldsymbol{v}\|_{4}\|\mathrm{A}(\boldsymbol{v}+\boldsymbol{z})\|_{4}\big{]}\|\boldsymbol{w}\|_{\mathbb{V}},$ (2.15)
$\langle\mathscr{J}(\boldsymbol{u}+\boldsymbol{z})-\mathscr{J}(\boldsymbol{v}+\boldsymbol{z}),\boldsymbol{w}\rangle$ $=\frac{1}{2}\int_{\mathscr{O}}[\mathrm{A}(\boldsymbol{u}-\boldsymbol{v})\mathrm{A}(\boldsymbol{u}+\boldsymbol{z})+\mathrm{A}(\boldsymbol{v}+\boldsymbol{z})\mathrm{A}(\boldsymbol{u}-\boldsymbol{v})]:\mathrm{A}(\boldsymbol{w})\mathrm{d}x$
$\leq C\big{[}\|\mathrm{A}(\boldsymbol{u}+\boldsymbol{z})\|_{4}+\|\mathrm{A}(\boldsymbol{v}+\boldsymbol{z})\|_{4}\big{]}\|\mathrm{A}(\boldsymbol{u}-\boldsymbol{v})\|_{4}\|\mathrm{A}(\boldsymbol{w})\|_{2},$ (2.16)
$\langle\mathscr{K}(\boldsymbol{u})-\mathscr{K}(\boldsymbol{v}),\boldsymbol{w}\rangle$ $=\frac{1}{2}\int_{\mathscr{O}}[\mathrm{A}(\boldsymbol{u}-\boldsymbol{v}):\mathrm{A}(\boldsymbol{u}+\boldsymbol{z})+\mathrm{A}(\boldsymbol{v}+\boldsymbol{z}):\mathrm{A}(\boldsymbol{u}-\boldsymbol{v})]\mathrm{A}(\boldsymbol{u}+\boldsymbol{z}):\mathrm{A}(\boldsymbol{w})\mathrm{d}x$
$\quad+\frac{1}{2}\int_{\mathscr{O}}|\mathrm{A}(\boldsymbol{v}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{u}-\boldsymbol{v}):\mathrm{A}(\boldsymbol{w})\mathrm{d}x$
$\leq C\big{[}\|\mathrm{A}(\boldsymbol{u}+\boldsymbol{z})\|_{4}^{2}+\|\mathrm{A}(\boldsymbol{v}+\boldsymbol{z})\|_{4}^{2}\big{]}\|\mathrm{A}(\boldsymbol{u}-\boldsymbol{v})\|_{4}\|\mathrm{A}(\boldsymbol{w})\|_{4}.$ (2.17)
Hence, the operator $\mathscr{K}(\cdot):\mathbb{X}\to\mathbb{X}^{\prime}$ is locally Lipschitz. ∎
2.3. Stochastic third-grade fluids equations
Here, we provide an abstract formulation of system (1.5). On taking projection $\mathscr{P}$ onto the first equation in (1.5), we obtain
[…]
where $\boldsymbol{x}\in\mathbb{H},\ \boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\{\mathrm{W}(t)\}_{t\in\mathbb{R}}$ is a two-sided cylindrical Wiener process in $\mathbb{H}$ with its RKHS $\mathrm{K}$ which satisfies Hypothesis 1.1. Now we present the definition of weak solution (in the analytic sense) to system (2.18) with the initial data $\boldsymbol{x}\in\mathbb{H}$ at the initial time $s\in\mathbb{R}.$
Definition 2.1.
[…]
$(\boldsymbol{v}(t),\boldsymbol{\phi})-(\boldsymbol{v}(s),\boldsymbol{\phi})$
$=-\int_{s}^{t}\langle\nu\mathscr{A}\boldsymbol{v}(\xi)+\mathscr{B}(\boldsymbol{v}(\xi))+\alpha\mathscr{J}(\boldsymbol{v}(\xi))+\beta\mathscr{K}(\boldsymbol{v}(\xi))-\boldsymbol{f},\boldsymbol{\phi}\rangle\mathrm{d}\xi+\int_{s}^{t}(\boldsymbol{\phi},\mathrm{d}\mathrm{W}(\xi)).$
The well-posedness of system (2.18) is discussed in Theorem 3.12 below.
3. RDS generated by stochastic third-grade fluids equations
In this section, we prove the existence of RDS generated by 3D stochastic third-grade fluids equations using a random transformation (known as Doss-Sussman transformation, [29, 59]). The process of generation of RDS has been adopted from the works [12, 13].
3.1. Metric dynamical system
Let us denote $\mathfrak{X}=\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})$. Let $\mathrm{E}$ denote the completion of $\mathscr{A}^{-\delta}(\mathfrak{X})$ with respect to the graph norm $\|x\|_{\mathrm{E}}=\|\mathscr{A}^{-\delta}x\|_{\mathfrak{X}},\text{ for }x\in\mathfrak{X}$, where $\|\cdot\|_{\mathfrak{X}}=\|\cdot\|_{2}+\|\cdot\|_{\mathbb{W}^{1,4}}$. Note that $\mathrm{E}$ is a separable Banach spaces (cf. [8]).
For $\xi\in(0,1/2)$, let us set
$\|\omega\|_{C^{\xi}_{1/2}(\mathbb{R};\mathrm{E})}=\sup_{t\neq s\in\mathbb{R}}\frac{\|\omega(t)-\omega(s)\|_{\mathrm{E}}}{|t-s|^{\xi}(1+|t|+|s|)^{1/2}}.$
Furthermore, we define
$C^{\xi}_{1/2}(\mathbb{R};\mathrm{E})$ =$=\left\{\omega\in C(\mathbb{R};\mathrm{E}):\omega(0)=\boldsymbol{0},\ \|\omega\|_{C^{\xi}_{1/2}(\mathbb{R};\mathrm{E})}<\infty\right\},$
$\Omega(\xi,\mathrm{E})$ =$=\overline{\{\omega\in C^{\infty}_{0}(\mathbb{R};\mathrm{E}):\omega(0)=0\}}^{C^{\xi}_{1/2}(\mathbb{R};\mathrm{E})}.$
The space $\Omega(\xi,\mathrm{E})$ is a separable Banach space. We also define
$C_{1/2}(\mathbb{R};\mathrm{E})=\left\{\omega\in C(\mathbb{R};\mathrm{E}):\omega(0)=0,\|\omega\|_{C_{1/2}(\mathbb{R};\mathrm{E})}=\sup_{t\in\mathbb{R}}\frac{\|\omega(t)\|_{\mathrm{E}}}{1+|t|^{\frac{1}{2}}}<\infty\right\}.$
Let us denote by $\mathscr{F}$, the Borel $\sigma$-algebra on $\Omega(\xi,\mathrm{E}).$ For $\xi\in(0,1/2)$, there exists a Borel probability measure $\mathbb{P}$ on $\Omega(\xi,\mathrm{E})$ (cf. [6]) such that the canonical process $\{w_{t},\ t\in\mathbb{R}\}$ defined by
$w_{t}(\omega):=\omega(t),\ \ \ \omega\in\Omega(\xi,\mathrm{E}),$ (3.1)
is an $\mathrm{E}$-valued two-sided Wiener process such that the RKHS of the Gaussian measure $\mathscr{L}(w_{1})$ on $\mathrm{E}$ is $\mathrm{K}$. For $t\in\mathbb{R},$ let $\mathscr{F}_{t}:=\sigma\{w_{s}:s\leq t\}.$ Then there exists a unique bounded linear map $\mathrm{W}(t):\mathrm{K}\to\mathrm{L}^{2}(\Omega(\xi,\mathrm{E}),\mathscr{F}_{t},\mathbb{P}).$ Moreover, the family $\{\mathrm{W}(t)\}_{t\in\mathbb{R}}$ is a $\mathrm{K}$-cylindrical Wiener process on a filtered probability space $(\Omega(\xi,\mathrm{E}),\mathscr{F},\{\mathscr{F}_{t}\}_{t\in\mathbb{R}},\mathbb{P})$ (cf. [14] for more details).
We consider a flow $\theta=(\theta_{t})_{t\in\mathbb{R}}$ on the space $C_{1/2}(\mathbb{R};\mathrm{E}),$ defined by
$\theta_{t}\omega(\cdot)=\omega(\cdot+t)-\omega(t),\ \ \ \omega\in C_{1/2}(\mathbb{R};\mathrm{E}),\ \ t\in\mathbb{R}.$
This flow keeps the spaces $C^{\xi}_{1/2}(\mathbb{R};\mathrm{E})$ and $\Omega(\xi,\mathrm{E})$ invariant and preserves $\mathbb{P}.$
Summing up, we have the following result:
Proposition 3.1 ([12, Proposition 6.13]).
The quadruple $(\Omega(\xi,\mathrm{E}),\mathscr{F},\mathbb{P},{\theta})$ is an MDS.
3.2. Ornstein-Uhlenbeck process
Let us first recall some analytic preliminaries from [12] which will help us to define an Ornstein-Uhlenbeck process. All the results of this subsection are valid for the space $\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})$ replaced by $\Omega(\xi,\mathbb{Y}).$
Proposition 3.2 ([12, Proposition 2.11]).
Let $-\mathbb{A}$ be the generator of an analytic semigroup $\{e^{t\mathbb{A}}\}_{t\geq 0}$ on a separable Banach space $\mathbb{Y}$ such that for some $C>0\ \text{and}\ \gamma>0$
$\|\mathbb{A}^{1+\delta}e^{-t\mathbb{A}}\|_{\mathfrak{L}(\mathbb{Y})}\leq C_{\delta}t^{-1-\delta}e^{-\gamma t},\ \ t>0,$ (3.2)
where $\mathfrak{L}(\mathbb{Y})$ denotes the space of all bounded linear operators from $\mathbb{Y}$ to $\mathbb{Y}$. For $\xi\in(\delta,1/2)$ and $\widetilde{\omega}\in\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y}),$ define
$\hat{\boldsymbol{z}}(t)=\hat{\boldsymbol{z}}(\mathbb{A};\widetilde{\omega})(t):=\int_{-\infty}^{t}\mathbb{A}^{1+\delta}e^{-(t-r)\mathbb{A}}(\widetilde{\omega}(t)-\widetilde{\omega}(r))\mathrm{d}r,\ \ t\in\mathbb{R}.$ (3.3)
If $t\in\mathbb{R},$ then $\hat{\boldsymbol{z}}(t)$ is a well-defined element of $\mathbb{Y}$ and the mapping
$\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})\ni\widetilde{\omega}\mapsto\hat{\boldsymbol{z}}(t)\in\mathbb{Y}$
is continuous. Moreover, the map $\hat{\boldsymbol{z}}:\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})\to\mathrm{C}_{1/2}(\mathbb{R};\mathbb{Y})$ is well defined, linear and bounded. In particular, there exists a constant $C>0$ such that for any $\widetilde{\omega}\in\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})$
$\|\hat{\boldsymbol{z}}(\widetilde{\omega})(t)\|_{\mathbb{Y}}\leq C(1+|t|^{1/2})\|\widetilde{\omega}\|_{\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})},\ \ \ t\in\mathbb{R}.$ (3.4)
Furthermore, under the same assumption, the following results hold (Corollaries 6.4, 6.6 and 6.8 in [12]):
  • 1.
    For all $-\infty<a<b<\infty$ and $t\in\mathbb{R}$, the map
    $\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y})\ni\widetilde{\omega}\mapsto(\hat{\boldsymbol{z}}(\widetilde{\omega})(t),\hat{\boldsymbol{z}}(\widetilde{\omega}))\in\mathbb{Y}\times\mathrm{L}^{q}(a,b;\mathbb{Y}),$ (3.5)
    where $q\in[1,\infty]$, is continuous.
  • 2.
    For any $\omega\in\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y}),$
    $\hat{\boldsymbol{z}}(\theta_{s}\omega)(t)=\hat{\boldsymbol{z}}(\omega)(t+s),\ \ t,s\in\mathbb{R}.$ (3.6)
  • 3.
    For $\zeta\in\mathrm{C}_{1/2}(\mathbb{R};\mathbb{Y}),$ if we put $\uptau_{s}(\zeta(t))=\zeta(t+s),\ t,s\in\mathbb{R},$ then, for $t\in\mathbb{R},\ \uptau_{s}\circ\hat{\boldsymbol{z}}=\hat{\boldsymbol{z}}\circ\theta_{s}$, that is,
    $\uptau_{s}\big{(}\hat{\boldsymbol{z}}(\omega)\big{)}=\hat{\boldsymbol{z}}\big{(}\theta_{s}(\omega)\big{)},\ \ \ \omega\in\mathrm{C}^{\xi}_{1/2}(\mathbb{R};\mathbb{Y}).$ (3.7)
Next, we define the Ornstein-Uhlenbeck process under Hypothesis 1.1. For $\delta$ as in Hypothesis 1.1, $\nu>0,\ \chi\geq 0,\ \xi\in(\delta,1/2)$ and $\omega\in C^{\xi}_{1/2}(\mathbb{R};\mathrm{E})$ (so that $(\nu\mathscr{A}+\chi\mathrm{I})^{-\delta}\omega\in C^{\xi}_{1/2}(\mathbb{R};\mathfrak{X})$), we define
$\boldsymbol{z}_{\chi}(\omega):=\hat{\boldsymbol{z}}((\nu\mathscr{A}+\chi\mathrm{I});(\nu\mathscr{A}+\chi\mathrm{I})^{-\delta}\omega)\ \in C_{1/2}(\mathbb{R};\mathfrak{X}),$
that is, for any $t\geq 0,$
$\boldsymbol{z}_{\chi}(\omega)(t)$ $=\int_{-\infty}^{t}(\nu\mathscr{A}+\chi\mathrm{I})^{1+\delta}e^{-(t-\tau)(\nu\mathscr{A}+\chi\mathrm{I})}((\nu\mathscr{A}+\chi\mathrm{I})^{-\delta}\theta_{\tau}\omega)(t-\tau)\mathrm{d}\tau.$ (3.8)
For $\omega\in C^{\infty}_{0}(\mathbb{R};\mathrm{E})$ with $\omega(0)=\boldsymbol{0},$ using integration by parts, we obtain
$\frac{\mathrm{d}\boldsymbol{z}_{\chi}(t)}{\mathrm{d}t}$ $=-(\nu\mathscr{A}+\chi\mathrm{I})\int_{-\infty}^{t}(\nu\mathscr{A}+\chi\mathrm{I})^{1+\delta}e^{-(t-r)(\nu\mathscr{A}+\chi\mathrm{I})}[(\nu\mathscr{A}+\chi\mathrm{I})^{-\delta}\omega(t)$
$\qquad\qquad-(\nu\mathscr{A}+\chi\mathrm{I})^{-\delta}\omega(r)]\mathrm{d}r+\frac{\mathrm{d}\omega(t)}{\mathrm{d}t}.$
Thus $\boldsymbol{z}_{\chi}(\cdot)$ is the solution of the following equation:
$\frac{\mathrm{d}\boldsymbol{z}_{\chi}(t)}{\mathrm{d}t}+(\nu\mathscr{A}+\chi\mathrm{I})\boldsymbol{z}_{\chi}(t)=\frac{\mathrm{d}\omega(t)}{\mathrm{d}t},\ \ t\in\mathbb{R}.$ (3.9)
Therefore, from the definition of the space $\Omega(\xi,\mathrm{E}),$ we have
Corollary 3.3.
If $\chi_{1},\chi_{2}\geq 0,$ then the difference $\boldsymbol{z}_{\chi_{1}}-\boldsymbol{z}_{\chi_{2}}$ is a solution to
$\frac{\mathrm{d}(\boldsymbol{z}_{\chi_{1}}-\boldsymbol{z}_{\chi_{2}})(t)}{\mathrm{d}t}+\nu\mathscr{A}(\boldsymbol{z}_{\chi_{1}}-\boldsymbol{z}_{\chi_{2}})(t)=-(\chi_{1}\boldsymbol{z}_{\chi_{1}}-\chi_{2}\boldsymbol{z}_{\chi_{2}})(t),\ \ \ t\in\mathbb{R}.$ (3.10)
According to the definition (3.1) of Wiener process $\{w_{t},\ t\in\mathbb{R}\},$ one can view the formula (3.8) as a definition of a process $\{\boldsymbol{z}_{\chi}(t),\ t\in\mathbb{R}\}$ on the probability space $(\Omega(\xi,\mathrm{E}),\mathscr{F},\mathbb{P})$. Equation (3.9) clearly tells that the process $\boldsymbol{z}_{\chi}(\cdot)$ is an Ornstein-Uhlenbeck process. Furthermore, the following results hold for $\boldsymbol{z}_{\chi}(\cdot)$.
Proposition 3.4 ([12, Proposition 6.10]).
The process $\{\boldsymbol{z}_{\chi}(t),\ t\in\mathbb{R}\},$ is a stationary Ornstein-Uhlenbeck process on $(\Omega(\xi,\mathrm{E}),\mathscr{F},\mathbb{P})$. It is a solution of the equation
$\mathrm{d}\boldsymbol{z}_{\chi}(t)+(\nu\mathscr{A}+\chi\mathrm{I})\boldsymbol{z}_{\chi}(t)\mathrm{d}t=\mathrm{d}\mathrm{W}(t),\ \ t\in\mathbb{R},$ (3.11)
[…]
$\boldsymbol{z}_{\chi}(t)=\int_{-\infty}^{t}e^{-(t-\xi)(\nu\mathscr{A}+\chi\mathrm{I})}\mathrm{d}\mathrm{W}(\xi),$ (3.12)
where the integral is an Itô integral on the M-type 2 Banach space $\mathfrak{X}$ (cf. [7]). In particular, for some $C$ depending on $\mathfrak{X}$,
$\mathbb{E}\left[\|\boldsymbol{z}_{\chi}(t)\|^{2}_{\mathfrak{X}}\right]$ $=\mathbb{E}\left[\left\|\int_{-\infty}^{t}e^{-(t-\xi)(\nu\mathscr{A}+\chi\mathrm{I})}\mathrm{d}\mathrm{W}(\xi)\right\|^{2}_{\mathfrak{X}}\right]\leq C\int_{-\infty}^{t}\|e^{-(t-\xi)(\nu\mathscr{A}+\chi\mathrm{I})}\|^{2}_{\gamma(\mathrm{K},\mathfrak{X})}\mathrm{d}\xi$
$=C\int_{0}^{\infty}e^{-2\chi\xi}\|e^{-\nu\xi\mathscr{A}}\|^{2}_{\gamma(\mathrm{K},\mathfrak{X})}\mathrm{d}\xi.$ (3.13)
Moreover, $\mathbb{E}\left[\|\boldsymbol{z}_{\chi}(t)\|^{2}_{\mathfrak{X}}\right]\to 0$ as $\chi\to\infty.$
Remark 3.1.
Note that (3.11) is the projected form of an equation of the following type:
[…]
where $p$ is a scalar field associated with projected equation (3.11).
Let us now provide some consequences of the previous discussion which will be used in the sequel.
Lemma 3.5.
For each $\omega\in\Omega$ and $c>0$, we obtain
$\lim_{t\to-\infty}\|\boldsymbol{z}_{\chi}(\omega)(t)\|^{2}_{2}\ e^{ct}=0.$
Proof.
Let us fix $\omega\in\Omega$. Because of (3.4), there exists a $\rho_{1}=\rho_{1}(\omega)\geq 0$ such that
$\frac{\|\boldsymbol{z}_{\chi}(t)\|_{2}}{|t|}\leq\rho_{1},\ \frac{\|\boldsymbol{z}_{\chi}(t)\|_{4}}{|t|}\leq\rho_{1}\ \text{ and }\ \frac{\|\boldsymbol{z}_{\chi}(t)\|_{\mathbb{W}^{1,4}}}{|t|}\leq\rho_{1}\ \text{ for }\ t\leq t_{0}.$ (3.15)
Therefore, we have, for every $\omega\in\Omega,$
$\lim_{t\to-\infty}\|\boldsymbol{z}_{\chi}(\omega)(t)\|^{2}_{2}\ e^{ct}\leq$ $\rho_{1}^{2}\lim_{t\to-\infty}|t|^{2}e^{ct}=0,$
which completes the proof. ∎
Lemma 3.6.
For each $\omega\in\Omega$ and $c>0$, we get
$\int_{-\infty}^{0}\bigg{\{}1+\|\boldsymbol{z}_{\chi}(t)\|^{2}_{2}+\|\boldsymbol{z}_{\chi}(t)\|^{4}_{\mathbb{W}^{1,4}}\bigg{\}}e^{ct}\mathrm{d}t<\infty.$
Proof.
Note that for $t_{0}\leq 0$,
$\int_{t_{0}}^{0}\bigg{\{}1+\|\boldsymbol{z}_{\chi}(t)\|^{2}_{2}+\|\boldsymbol{z}_{\chi}(t)\|^{4}_{\mathbb{W}^{1,4}}\bigg{\}}e^{ct}\mathrm{d}t<\infty.$
Therefore, we only need to show that the integral
$\int_{-\infty}^{t_{0}}\bigg{\{}1+\|\boldsymbol{z}_{\chi}(t)\|^{2}_{2}+\|\boldsymbol{z}_{\chi}(t)\|^{4}_{\mathbb{W}^{1,4}}\bigg{\}}e^{ct}\mathrm{d}t<\infty.$
In view of (3.15), we obtain
$\int_{-\infty}^{t_{0}}\bigg{\{}\|\boldsymbol{z}_{\chi}(t)\|^{2}_{2}+\|\boldsymbol{z}_{\chi}(t)\|^{4}_{\mathbb{W}^{1,4}}\bigg{\}}e^{ct}\mathrm{d}t\leq\int_{-\infty}^{t_{0}}\big{\{}\rho^{2}_{1}|t|^{2}+\rho^{4}_{1}|t|^{4}\big{\}}e^{ct}\mathrm{d}t<\infty,$
which completes the proof. ∎
Definition 3.1.
A function $\kappa:\Omega\to(0,\infty)$ belongs to the class $\mathfrak{K}$ if and only if
$\lim_{t\to\infty}[\kappa(\theta_{-t}\omega)]^{2}e^{-ct}=0,$ (3.16)
for all $c>0$.
Let us denote the class of all closed and bounded random sets $D$ on $\mathbb{H}$ by $\mathfrak{DK}$ such that the radius function $\Omega\ni\omega\mapsto\kappa(D(\omega)):=\sup\{\|x\|_{2}:x\in D(\omega)\}$ belongs to the class $\mathfrak{K}.$ It is straight forward that the constant functions belong to $\mathfrak{K}$. It is clear by Definition 3.1 that the class $\mathfrak{K}$ is closed with respect to sum, multiplication by a constant and if $\kappa\in\mathfrak{K},0\leq\bar{\kappa}\leq\kappa,$ then $\bar{\kappa}\in\mathfrak{K}.$
Proposition 3.7.
We define functions $\kappa_{i}:\Omega\to(0,\infty),i=1,\ldots,4,$ by the following formulae: for $\omega\in\Omega,$
$[\kappa_{1}(\omega)]^{2}$ $:=\|\boldsymbol{z}_{\chi}(\omega)(0)\|_{2},$ $[\kappa_{2}(\omega)]^{2}:=\sup_{s\leq 0}\|\boldsymbol{z}_{\chi}(\omega)(s)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s},$
$[\kappa_{3}(\omega)]^{2}$ $:=\int_{-\infty}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\|\boldsymbol{z}_{\chi}(\omega)(t)\|^{2}_{2}\ \mathrm{d}t,$ $[\kappa_{4}(\omega)]^{2}:=\int_{-\infty}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\|\boldsymbol{z}_{\chi}(\omega)(t)\|^{4}_{\mathbb{W}^{1,4}}\ \mathrm{d}t.$
Then all these functions belongs to class $\mathfrak{K}.$
Proof.
Let us recall from (3.6) that $\boldsymbol{z}_{\chi}(\theta_{-t}\omega)(s)=\boldsymbol{z}_{\chi}(\omega)(s-t)$. For any given $c>0$, we consider
$\lim_{t\to\infty}[\kappa_{1}(\theta_{-t}\omega)]^{2}e^{-ct}=$ $\lim_{t\to\infty}\|\boldsymbol{z}_{\chi}(\theta_{-t}\omega)(0)\|^{2}_{2}e^{-ct}=\lim_{t\to\infty}\|\boldsymbol{z}_{\chi}(\omega)(-t)\|^{2}_{2}e^{-ct}.$
Using Lemma 3.5, we have, $\kappa_{1}\in\mathfrak{K}.$
Next, for any given $c>0$, we define $c_{1}:=\min\{\frac{c}{2},\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\}$ and consider
$[\kappa_{2}(\theta_{-t}\omega)]^{2}e^{-ct}$ $=e^{-ct}\sup_{s\leq 0}\|\boldsymbol{z}_{\chi}(\omega)(s-t)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}$
$=e^{-ct}\sup_{\sigma\leq-t}\|\boldsymbol{z}_{\chi}(\omega)(\sigma)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(\sigma+t)}$
$\leq e^{-ct}\sup_{\sigma\leq-t}\|\boldsymbol{z}_{\chi}(\omega)(\sigma)\|^{2}_{2}\ e^{c_{1}(\sigma+t)}$
$=e^{-(c-c_{1})t}\sup_{\sigma\leq-t}\|\boldsymbol{z}_{\chi}(\omega)(\sigma)\|^{2}_{2}\ e^{c_{1}\sigma},$
which implies
$\lim_{t\to\infty}[\kappa_{2}(\theta_{-t}\omega)]^{2}e^{-ct}\leq\lim_{t\to\infty}e^{-(c-c_{1})t}\lim_{t\to\infty}\left[\sup_{\sigma\leq-t}\|\boldsymbol{z}_{\chi}(\omega)(\sigma)\|^{2}_{2}\ e^{c_{1}\sigma}\right]=0,$
where we have used Lemma 3.5. This implies that $\kappa_{2}\in\mathfrak{K}.$
Finally, from Lemma 3.6 and absolute continuity of Lebesgue integrals, we obtain
$\bigg{\{}[\kappa_{3}(\theta_{-t}\omega)]^{2}+[\kappa_{4}(\theta_{-t}\omega)]^{2}\bigg{\}}e^{-ct}$ $=e^{-ct}\int_{-\infty}^{0}\bigg{\{}\|\boldsymbol{z}_{\chi}(\theta_{-t}\omega)(\sigma)\|^{2}_{2}+\|\boldsymbol{z}_{\chi}(\theta_{-t}\omega)(\sigma)\|^{2}_{\mathbb{W}^{1,4}}\bigg{\}}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\sigma}\mathrm{d}\sigma$
$=e^{-ct}\int_{-\infty}^{0}\bigg{\{}\|\boldsymbol{z}_{\chi}(\omega)(\sigma-t)\|^{2}_{2}+\|\boldsymbol{z}_{\chi}(\omega)(\sigma-t)\|^{2}_{\mathbb{W}^{1,4}}\bigg{\}}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\sigma}\mathrm{d}\sigma$
$=e^{-ct}\int_{-\infty}^{-t}\bigg{\{}\|\boldsymbol{z}_{\chi}(\omega)(\sigma)\|^{2}_{2}+\|\boldsymbol{z}_{\chi}(\omega)(\sigma)\|^{2}_{\mathbb{W}^{1,4}}\bigg{\}}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(\sigma+t)}\mathrm{d}\sigma$
$\leq e^{-ct}\int_{-\infty}^{-t}\bigg{\{}\|\boldsymbol{z}_{\chi}(\omega)(\sigma)\|^{2}_{2}+\|\boldsymbol{z}_{\chi}(\omega)(\sigma)\|^{2}_{\mathbb{W}^{1,4}}\bigg{\}}e^{c_{1}(\sigma+t)}\mathrm{d}\sigma$
$\leq e^{-(c-c_{1})t}\int_{-\infty}^{-t}\bigg{\{}\|\boldsymbol{z}_{\chi}(\omega)(\sigma)\|^{2}_{2}+\|\boldsymbol{z}_{\chi}(\omega)(\sigma)\|^{2}_{\mathbb{W}^{1,4}}\bigg{\}}e^{c_{1}\sigma}\mathrm{d}\sigma$
$\to 0\ \text{ as }\ t\to\infty.$
This implies that $\kappa_{3},\kappa_{4}\in\mathfrak{K}$, which completes the proof. ∎
3.3. Random dynamical system
Remember that Hypothesis 1.1 is satisfied and that $\delta$ has the property stated there. Let us fix $\nu>0$, and the parameters $\chi\geq 0$ and $\xi\in(\delta,1/2)$.
Using a random transformation (known as Doss-Sussman transformation, [29, 59]), we get a random partial differential equation which equivalent to system (2.18). Let us define
$\boldsymbol{y}^{\chi}:=\boldsymbol{v}-\boldsymbol{z}_{\chi}(\omega).$ (3.17)
For convenience, we write $\boldsymbol{y}^{\chi}(t)=\boldsymbol{y}(t)$ and $\boldsymbol{z}_{\chi}(\omega)(t)=\boldsymbol{z}(t)$. Then $\boldsymbol{y}(\cdot)$ satisfies the following system:
$\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}&=-\underbrace{\nabla(\textbf{P}-{p})}_{:=\nabla\widehat{\textbf{P}}}+\nu\Delta\boldsymbol{y}-\big{(}(\boldsymbol{y}+\boldsymbol{z})\cdot\nabla\big{)}(\boldsymbol{y}+\boldsymbol{z})+\alpha\text{div}((\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2})&&\\ &\quad+\beta\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))+\chi\boldsymbol{z}+\boldsymbol{f}&&\text{in}\mathscr{O}\times(0,\infty),\\ \text{div}\;\boldsymbol{y}&=0\quad&&\text{in }\mathscr{O}\times[0,\infty),\\ \boldsymbol{y}&=\boldsymbol{0}&&\text{on }\partial\mathscr{O}\times[0,\infty),\\ \boldsymbol{y}(x,0)&=\boldsymbol{x}-\boldsymbol{z}_{\chi}(\omega)(0)=:\boldsymbol{y}_{0}\quad&&\text{in }\mathscr{O},\end{aligned}\right.$ (3.18)
where $p$ is the scalar appearing in (3.14), and the following projected system:
[…]
Since $\boldsymbol{z}_{\chi}(\omega)\in C_{1/2}(\mathbb{R};\mathfrak{X}),$ then $\boldsymbol{z}_{\chi}(\omega)(0)$ is a well defined element of $\mathbb{H}$. For each fixed $\omega\in\Omega$, system (3.19) is a deterministic system. Let us now provide the definition of weak solution (in the deterministic sense, for each fixed $\omega$) for (3.19).
Definition 3.2.
Assume that $\boldsymbol{y}_{0}\in\mathbb{H}$, $\boldsymbol{z}\in\mathrm{L}^{2}_{\mathrm{loc}}([0,\infty);\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}([0,\infty);\mathbb{W}^{1,4}(\mathscr{O}))$ and $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$. A function $\boldsymbol{y}$ is called a weak solution of system (3.19) on the time interval $[0,\infty)$, if
$\boldsymbol{y}$ $\in\mathrm{C}_{w}([0,\infty);\mathbb{H})\cap\mathrm{L}^{2}_{\mathrm{loc}}(0,\infty;\mathbb{V})\cap\mathrm{L}^{4}_{\mathrm{loc}}(0,\infty;\mathbb{W}^{1,4}(\mathscr{O})),$
$\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}$ $\in\mathrm{L}^{2}_{\mathrm{loc}}(0,\infty;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}_{\mathrm{loc}}(0,\infty;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})),$
and it satisfies
  • (i)
    for any $\boldsymbol{\phi}\in\mathbb{V},$
    $\frac{\mathrm{d}\boldsymbol{y}(t)}{\mathrm{d}t},\boldsymbol{\phi}\right>$
    $=-\left\langle\nu\mathscr{A}\boldsymbol{y}(t)+\mathscr{B}(\boldsymbol{y}(t)+\boldsymbol{z}(t))+\alpha\mathscr{J}(\boldsymbol{y}(t)+\boldsymbol{z}(t))+\beta\mathscr{K}(\boldsymbol{y}(t)+\boldsymbol{z}(t))-\chi\boldsymbol{z}(t)-\boldsymbol{f},\boldsymbol{\phi}\right\rangle,$
    for a.e. $t\in[0,\infty);$
  • (ii)
    the initial data:
    $\boldsymbol{y}(0)=\boldsymbol{y}_{0}\ \text{ in }\ \mathbb{H}.$
Theorem 3.8.
Assume that (1.4) is satisfied, $\chi\geq 0$, $\boldsymbol{v}_{0}\in\mathbb{H}$, $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\boldsymbol{z}\in\mathrm{L}^{2}_{\mathrm{loc}}([0,\infty);\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}([0,\infty);\mathbb{W}^{1,4}(\mathscr{O}))$. Then, there exists a unique weak solution $\boldsymbol{y}(\cdot)$ to system (3.19) in the sense of Definition 3.2 which satisfies the following energy equality:
$\|\boldsymbol{y}(t)\|_{2}^{2}+\nu\int_{0}^{t}\|\mathrm{A}(\boldsymbol{y}(s))\|_{2}^{2}\mathrm{d}s+\beta\int_{0}^{t}\|\mathrm{A}(\boldsymbol{y}(s)+\boldsymbol{z}(s))\|_{4}^{4}\mathrm{d}s+\alpha\int_{0}^{t}\mathrm{Tr}\big{(}[\mathrm{A}(\boldsymbol{y}(s)+\boldsymbol{z}(s))]^{3}\big{)}\mathrm{d}s$
$=\|\boldsymbol{y}_{0}\|_{2}^{2}+2\int_{0}^{t}\langle\mathscr{B}(\boldsymbol{y}(s)+\boldsymbol{z}(s))+\alpha\mathscr{J}(\boldsymbol{y}(s)+\boldsymbol{z}(s))+\beta\mathscr{K}(\boldsymbol{y}(s)+\boldsymbol{z}(s)),\boldsymbol{z}(s)\rangle\mathrm{d}s$
$\quad+2\int_{0}^{t}\langle\chi\boldsymbol{z}(s)+\boldsymbol{f},\boldsymbol{y}(s)\rangle\mathrm{d}s,$ (3.20)
for all $t\in[0,\infty)$. Consequently, $\boldsymbol{y}\in\mathrm{C}([0,\infty);\mathbb{H})$.
Proof of Theorem 3.8.
Let us fix $T>0$. Note that it is enough to prove the result on the interval $[0,T]$. The proof is divided into the following four steps. Step I. Finite-dimensional approximation and weak limits. Let $\{\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{m},\ldots\}$ be a complete orthonormal system in $\mathbb{H}$ belonging to $\mathbb{V}_{s}\subset\mathbb{X}$ (where $s>2$) and let $\mathbb{H}_{m}:=\mathrm{span}\{\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{m}\}$. Let $\mathrm{P}_{m}$ denote the orthogonal projection of $\mathbb{X}^{\prime}$ to $\mathbb{H}_{m}$, that is,
$\mathrm{P}_{m}\boldsymbol{x}=\sum_{i=1}^{m}\langle\boldsymbol{x},\boldsymbol{e}_{i}\rangle\boldsymbol{e}_{i}.$ (3.21)
Since every element $\boldsymbol{x}\in\mathbb{H}$ induces a functional $\boldsymbol{x}^{*}\in\mathbb{H}$ by the formula $\langle\boldsymbol{x}^{*},\widetilde{\boldsymbol{x}}\rangle=(\boldsymbol{x},\widetilde{\boldsymbol{x}})$, $\widetilde{\boldsymbol{x}}\in\mathbb{X}$, then $\mathrm{P}_{m}\big{|}_{\mathbb{H}}$, the orthogonal projection of $\mathbb{H}$ onto $\mathbb{H}_{m}$ is given by $\mathrm{P}_{m}\boldsymbol{x}=\sum_{i=1}^{m}\left(\boldsymbol{x},\boldsymbol{e}_{i}\right)\boldsymbol{e}_{i}$. Hence in particular, $\mathrm{P}_{m}$ is the orthogonal projection from $\mathbb{H}$ onto $\mathbb{H}_{m}$.
Let us consider the following system of ODEs on the finite dimensional space $\mathbb{H}_{n}$:
[…]
where
$\mathscr{A}_{m}\boldsymbol{y}^{m}=\mathrm{P}_{m}\mathscr{A}\boldsymbol{y}^{m},\quad\mathscr{B}_{m}(\boldsymbol{y}^{m}+\boldsymbol{z})=\mathrm{P}_{m}\mathscr{B}(\boldsymbol{y}^{m}+\boldsymbol{z}),\quad\mathscr{J}_{m}(\boldsymbol{y}^{m}+\boldsymbol{z})=\mathrm{P}_{m}\mathscr{J}(\boldsymbol{y}^{m}+\boldsymbol{z}),$
$\mathscr{K}_{m}(\boldsymbol{y}^{m}+\boldsymbol{z})=\mathrm{P}_{m}\mathscr{K}(\boldsymbol{y}^{m}+\boldsymbol{z}),\quad\boldsymbol{z}_{m}=\mathrm{P}_{m}\boldsymbol{z},\quad\boldsymbol{f}_{m}=\mathrm{P}_{m}[\mathscr{P}\boldsymbol{f}]\quad\text{ and }\quad{\boldsymbol{y}_{0}}_{m}=\mathrm{P}_{m}[\boldsymbol{y}_{0}].$
Since $\mathscr{A}_{n}\cdot+\mathscr{B}_{n}(\cdot+\boldsymbol{z})+\alpha\mathscr{J}_{n}(\cdot+\boldsymbol{z})+\beta\mathscr{K}_{n}(\cdot+\boldsymbol{z})$ is locally Lipschitz (see Subsection 2.2), system (3.22) has a unique local solution $\boldsymbol{y}^{m}\in\mathrm{C}([0,T^{*}];\mathbb{H}_{n})$, for some $0<T^{*}<T$. The following uniform estimates show that the time $T^{*}$ can be extended to time $T$. Taking the inner product with $\boldsymbol{y}^{m}(\cdot)$ to the first equation of (3.22) and using (2.4), we obtain
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\boldsymbol{y}^{m}(t)\|^{2}_{2}+\frac{\nu}{2}\|\mathrm{A}(\boldsymbol{y}^{m}(t))\|^{2}_{2}+\frac{\beta}{2}\|\mathrm{A}(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t))\|^{4}_{4}$
$=-b(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t),\boldsymbol{y}^{m}(t),\boldsymbol{z}(t))+\frac{\alpha}{2}\int_{\mathscr{O}}(\mathrm{A}(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t)))^{2}:\mathrm{A}(\boldsymbol{y}^{m}(t))\mathrm{d}x$
$\quad+\frac{\beta}{2}\int_{\mathscr{O}}|\mathrm{A}(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t))|^{2}[\mathrm{A}(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t)):\mathrm{A}(\boldsymbol{z}(t))]\mathrm{d}x+\chi(\boldsymbol{z}(t),\boldsymbol{y}^{m}(t))+\langle\boldsymbol{f},\boldsymbol{y}^{m}(t)\rangle,$ (3.23)
for all $t\in[0,T]$. Next, using the Hölder’s inequality, Lemma 2.1, Korn-type inequality (2.2), (1.4) (see Remark 1.1 also), Gagliardo-Nirenberg inequality (2.1) and Young’s inequality, we estimate the terms of the right hand side of (3.3) as
$|b(\boldsymbol{y}^{m}+\boldsymbol{z},\boldsymbol{y}^{m},\boldsymbol{z})|$ $\leq|b(\boldsymbol{y}^{m},\boldsymbol{y}^{m}+\boldsymbol{z},\boldsymbol{z})|+|b(\boldsymbol{z},\boldsymbol{y}^{m},\boldsymbol{z})|$
$\leq\|\boldsymbol{y}^{m}\|_{2}\|\nabla(\boldsymbol{y}^{m}+\boldsymbol{z})\|_{4}\|\boldsymbol{z}\|_{4}+\|\boldsymbol{z}\|_{4}^{2}\|\nabla\boldsymbol{y}^{m}\|_{2}$
$\leq C\|\mathrm{A}(\boldsymbol{y}^{m})\|_{2}\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|_{4}\|\boldsymbol{z}\|_{4}+C\|\boldsymbol{z}\|_{4}^{2}\|\mathrm{A}(\boldsymbol{y}^{m})\|_{2}$
$\leq\frac{\nu\varepsilon_{0}}{24}\|\mathrm{A}(\boldsymbol{y}^{m})\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{12}\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{4}_{4}+C\|\boldsymbol{z}\|_{\mathbb{W}^{1,4}}^{4},$ (3.24)
$|\displaystyle\left|\frac{\alpha}{2}\int_{\mathscr{O}}(\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z}))^{2}:\mathrm{A}(\boldsymbol{y}^{m})\mathrm{d}x\right|$$\leq\frac{\left|\alpha\right|}{2}\int_{\mathscr{O}}\left|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\right|^{2}\left|\mathrm{A}(\boldsymbol{y}^{m})\right|\mathrm{d}x$
$\leq\frac{\left|\alpha\right|}{2}\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{2}_{4}\|\mathrm{A}(\boldsymbol{y}^{m})\|_{2}$
$\leq\frac{\nu(1-\varepsilon_{0})}{4}\|\mathrm{A}(\boldsymbol{y}^{m})\|^{2}_{2}+\frac{\alpha^{2}}{4\nu(1-\varepsilon_{0})}\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{4}_{4}$
$\leq\frac{\nu(1-\varepsilon_{0})}{4}\|\mathrm{A}(\boldsymbol{y}^{m})\|^{2}_{2}+\frac{\beta(1-\varepsilon_{0})}{2}\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{4}_{4}$ (3.25)
$\left|\frac{\beta}{2}\int_{\mathscr{O}}|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})|^{2}[\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z}):\mathrm{A}(\boldsymbol{z})]\mathrm{d}x\right|$ $\leq\frac{\beta}{2}\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{3}_{4}\|\mathrm{A}(\boldsymbol{z})\|_{4}$
$\leq\frac{\beta\varepsilon_{0}}{12}\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{4}_{4}+C\|\mathrm{A}(\boldsymbol{z})\|^{4}_{4},$ (3.26)
$|\chi(\boldsymbol{z},\boldsymbol{y}^{m})|$ $\leq\chi\|\boldsymbol{z}\|_{2}\|\boldsymbol{y}^{m}\|_{2}\leq C\|\boldsymbol{z}\|_{2}\|\mathrm{A}(\boldsymbol{y}^{m})\|_{2}$
$\leq\frac{\nu\varepsilon_{0}}{24}\|\mathrm{A}(\boldsymbol{y}^{m})\|^{2}_{2}+C\|\boldsymbol{z}\|^{2}_{2},$ (3.27)
and
$\big{|}\langle\boldsymbol{f},\boldsymbol{y}^{m}\rangle\big{|}$
$=\big{|}\langle\boldsymbol{f}_{1}+\boldsymbol{f}_{2},\boldsymbol{y}^{m}\rangle\big{|}$
$\leq\|\boldsymbol{f}_{1}\|_{\mathbb{H}^{-1}}\|\boldsymbol{y}^{m}\|_{\mathbb{V}}+\|\boldsymbol{f}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}\|\boldsymbol{y}^{m}\|_{\mathbb{W}^{1,4}}$
$\leq\|\boldsymbol{f}_{1}\|_{\mathbb{H}^{-1}}\|\boldsymbol{y}^{m}\|_{\mathbb{V}}+C\|\boldsymbol{f}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}(\|\boldsymbol{y}^{m}\|_{{4}}+\|\nabla\boldsymbol{y}^{m}\|_{{4}})$
$\leq\|\boldsymbol{f}_{1}\|_{\mathbb{H}^{-1}}\|\boldsymbol{y}^{m}\|_{\mathbb{V}}+C\|\boldsymbol{f}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}(\|\mathrm{A}(\boldsymbol{y}^{m})\|^{\frac{d}{4+d}}_{{4}}\|\boldsymbol{y}^{m}\|_{2}^{\frac{4}{4+d}}+\|\mathrm{A}(\boldsymbol{y}^{m})\|_{{4}})$
$\leq C\|\boldsymbol{f}_{1}\|_{\mathbb{H}^{-1}}\|\mathrm{A}(\boldsymbol{y}^{m})\|_{2}+C\|\boldsymbol{f}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}(\|\mathrm{A}(\boldsymbol{y}^{m})\|^{\frac{d}{4+d}}_{{4}}\|\mathrm{A}(\boldsymbol{y}^{m})\|_{2}^{\frac{4}{4+d}}+\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|_{{4}}+\|\mathrm{A}(\boldsymbol{z})\|_{{4}})$
$\leq\frac{\nu\varepsilon_{0}}{24}\|\mathrm{A}(\boldsymbol{y}^{m})\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{12}\|\mathrm{A}(\boldsymbol{y}^{m}+\boldsymbol{z})\|^{4}_{4}+C\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\mathrm{A}(\boldsymbol{z})\|^{4}_{4}\bigg{]}.$ (3.28)
Combining (3.3)-(3.28), we deduce
$\frac{\mathrm{d}}{\mathrm{d}t}\|\boldsymbol{y}^{m}(t)\|^{2}_{2}+\frac{\nu}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\|\mathrm{A}(\boldsymbol{y}^{m}(t))\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{2}\|\mathrm{A}(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t))\|^{4}_{4}$
$\leq C\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}\bigg{]},$ (3.29)
which gives for all $t\in[0,T]$
$\|\boldsymbol{y}^{m}(t)\|^{2}_{2}+\frac{\nu}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\int_{0}^{t}\|\mathrm{A}(\boldsymbol{y}^{m}(s))\|^{2}_{2}\mathrm{d}s+\frac{\beta\varepsilon_{0}}{2}\int_{0}^{t}\|\mathrm{A}(\boldsymbol{y}^{m}(s))\|^{4}_{4}\mathrm{d}s$
$\leq\|\boldsymbol{y}^{m}(0)\|^{2}_{2}+C\int_{0}^{t}\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{z}(s)\|^{2}_{2}+\|\boldsymbol{z}(s)\|^{4}_{\mathbb{W}^{1,4}}\bigg{]}\mathrm{d}s.$ (3.30)
Hence, an application of Gronwall’s inequality and the fact that $\|\boldsymbol{y}^{m}(0)\|_{2}\leq\|\boldsymbol{y}_{0}\|_{2}$, $\boldsymbol{f}_{1}\in\mathbb{H}^{-1}(\mathscr{O})$, $\boldsymbol{f}_{2}\in\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\boldsymbol{z}\in\mathrm{L}^{2}(0,T;\mathbb{H})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$, we have from (3.3) that
$\{\boldsymbol{y}^{m}\}_{m\in\mathbb{N}}\text{ is a bounded sequence in }\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})).$ (3.31)
For any arbitrary element $\boldsymbol{\phi}\in\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$, using Hölder’s inequality and Lemma 2.1, we have from (3.22) that
$\left|\int_{0}^{T}\left\langle\frac{\mathrm{d}\boldsymbol{y}^{m}(t)}{\mathrm{d}t},\boldsymbol{\phi}(t)\right\rangle\mathrm{d}t\right|$
$\leq\int_{0}^{T}\bigg{[}\nu\left|(\nabla\boldsymbol{y}^{m}(t),\mathrm{P}_{m}\nabla\boldsymbol{\phi}(t))\right|+\left|b(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t),\mathrm{P}_{m}\boldsymbol{\phi}(t),\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t))\right|+\chi|(\boldsymbol{z}(t),\mathrm{P}_{m}\boldsymbol{\phi}(t))|$
$\qquad+|\langle\boldsymbol{f},\mathrm{P}_{m}\boldsymbol{\phi}(t)\rangle|\bigg{]}\mathrm{d}t+|\alpha|\int_{0}^{T}\int_{\mathscr{O}}|\mathrm{A}(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t))|^{2}|\mathrm{P}_{m}\nabla\boldsymbol{\phi}(t)|\mathrm{d}x\mathrm{d}t$
$\qquad+\beta\int_{0}^{T}\int_{\mathscr{O}}|\mathrm{A}(\boldsymbol{y}^{m}(t)+\boldsymbol{z}(t))|^{3}|\mathrm{P}_{m}\nabla\boldsymbol{\phi}(t)|\mathrm{d}x\mathrm{d}t$
$\leq C\bigg{[}\|\boldsymbol{y}^{m}\|_{\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4})}+T^{\frac{1}{2}}\|\boldsymbol{f}_{1}\|_{\mathbb{H}^{-1}}+T^{\frac{3}{4}}\|\boldsymbol{f}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{\boldsymbol{z}}\|_{\mathrm{L}^{2}(0,T;\mathbb{H})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4})}\bigg{]}$
×$\qquad\times\|\boldsymbol{\phi}\|_{\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4})}$
which implies that the sequences
$\{\frac{\mathrm{d}\boldsymbol{y}^{m}}{\mathrm{d}t}\}_{m\in\mathbb{N}}$ and $\{\mathscr{G}(\boldsymbol{y}^{m})\}_{m\in\mathbb{N}}$ are bounded in $\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})),$ (3.32)
where $\mathscr{G}(\cdot)$ given by (2.5). Taking (3.31)-(3.32) into account and using the Banach-Alaoglu theorem, we obtain the existence of an element $\boldsymbol{y}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ with $\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}\in\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$ and $\mathscr{G}_{0}\in\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$ such that
[…]
along a subsequence (still denoted by the same symbol). Note that $\boldsymbol{f}_{m}\to\mathscr{P}\boldsymbol{f}$ in $\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\boldsymbol{z}_{m}\to\boldsymbol{z}$ in $\mathrm{L}^{2}(0,T;\mathbb{H})$. Therefore, on passing to limit as $m\to\infty$ in (3.22), the limit $\boldsymbol{y}(\cdot)$ satisfies:
$\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}=-\mathscr{G}_{0}+\chi\boldsymbol{z}+\mathscr{P}\boldsymbol{f},$ (3.37)
in $\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$. Note that the embedding of $\mathbb{H}\subset\mathbb{V}^{\prime}$ is continuous and $\boldsymbol{y}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})$ implies $\boldsymbol{y}\in\mathrm{L}^{\infty}(0,T;\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$. Thus, we get $\boldsymbol{y},\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}\in\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$ and then invoking [31, Theorem 2, Section 5.9.2], it is immediate that $\boldsymbol{y}\in\mathrm{C}([0,T];\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$. Since $\mathbb{H}$ is reflexive, using [22, Proposition 1.7.1], we obtain $\boldsymbol{y}\in\mathrm{C}_{w}([0,T];\mathbb{H})$ and the map $t\mapsto\|\boldsymbol{y}(t)\|_{2}$ is bounded. Thus the condition (ii) in the Definition (3.2) makes sense.
Next, note that [21, Chapter II, Theorem 1.8], $\boldsymbol{y}\in\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ and $\frac{\mathrm{d}\boldsymbol{y}}{\mathrm{d}t}\in\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$ imply $\boldsymbol{y}\in\mathrm{C}([0,T];\mathbb{H})$, the real-valued function $t\mapsto\|\boldsymbol{y}(t)\|_{2}^{2}$ is absolutely continuous and the following equality is satisfied:
$\frac{\mathrm{d}}{\mathrm{d}t}\|\boldsymbol{y}(t)\|_{2}^{2}=2\left<\frac{\mathrm{d}\boldsymbol{y}(t)}{\mathrm{d}t},\boldsymbol{y}(t)\right>,\ \ \ \ \text{ for a.e. }t\in[0,T].$ (3.38)
Step II. Minty-Browder technique: From (3.38), for a measurable function $\eta(t)\geq 0$, we have the following equality:
$e^{-2\eta(t)}\|\boldsymbol{y}(t)\|_{2}^{2}+2\int_{0}^{t}e^{-2\eta(s)}\left\langle\mathscr{G}_{0}(s)-\chi\boldsymbol{z}(s)-\boldsymbol{f}+\eta^{\prime}(s)\boldsymbol{y}(s),\boldsymbol{y}(s)\right\rangle\mathrm{d}s=\|\boldsymbol{y}(0)\|_{2}^{2},$ (3.39)
for all $t\in[0,T]$. Similar to (3.39), for system (3.22), we obtain the following energy equality:
$e^{-2\eta(t)}\|\boldsymbol{y}^{m}(t)\|_{2}^{2}+2\int_{0}^{t}e^{-2\eta(s)}\left\langle\mathscr{G}(\boldsymbol{y}^{m}(s))-\chi\boldsymbol{z}_{m}(s)-\boldsymbol{f}_{m}+\eta^{\prime}(s)\boldsymbol{y}^{m}(s),\boldsymbol{y}^{m}(s)\right\rangle\mathrm{d}s=\|\boldsymbol{y}^{m}(0)\|_{2}^{2},$ (3.40)
for all $t\in[0,T]$. Remember that $\boldsymbol{y}^{m}(0)=\mathrm{P}_{m}\boldsymbol{v}(0)$, and hence the initial value $\boldsymbol{y}^{m}(0)$ converges strongly in $\mathbb{H}$, that is, we have
$\lim_{m\to\infty}\|\boldsymbol{y}^{m}(0)-\boldsymbol{y}(0)\|_{2}=0.$ (3.41)
For any $\boldsymbol{\phi}\in\mathrm{L}^{\infty}(0,T;\mathbb{H}_{n})$ with $n<m$, we define
$\eta(t)=\frac{(C_{S}C_{K})^{2}}{\nu\varepsilon_{0}}\int_{0}^{t}\|\mathrm{A}(\boldsymbol{\phi}(s)+\boldsymbol{z}(s))\|_{{4}}^{2}\mathrm{d}s,\mbox{ for all }t\in[0,T].$
Using (2.2), we obtain
$\int_{0}^{T}e^{-2\eta(t)}\{\left<\mathscr{G}(\boldsymbol{\phi}(t))-\mathscr{G}(\boldsymbol{y}^{m}(t)),\boldsymbol{\phi}(t)-\boldsymbol{y}^{m}(t)\right>+\eta^{\prime}(t)(\boldsymbol{\phi}(t)-\boldsymbol{y}^{m}(t),\boldsymbol{\phi}(t)-\boldsymbol{y}^{m}(t))\}\mathrm{d}t\geq 0.$ (3.42)
Making use of (3.40) in (3.42), we obtain
$\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}(\boldsymbol{\phi}(t))+\eta^{\prime}(t)\boldsymbol{\phi}(t),\boldsymbol{\phi}(t)-\boldsymbol{y}^{m}(t)\right>\mathrm{d}t$
$\geq\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}(\boldsymbol{y}^{m}(t))+\eta^{\prime}(t)\boldsymbol{y}^{m}(t),\boldsymbol{\phi}(t)-\boldsymbol{y}^{m}(t)\right>\mathrm{d}t$
$=\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}(\boldsymbol{y}^{m}(t))+\eta^{\prime}(t)\boldsymbol{y}^{m}(t),\boldsymbol{\phi}(t)\right>\mathrm{d}t+\frac{1}{2}\bigg{[}e^{-2\eta(T)}\|\boldsymbol{y}^{m}(T)\|_{2}^{2}-\|\boldsymbol{y}^{m}(0)\|_{2}^{2}\bigg{]}$
$\qquad-\chi\int_{0}^{T}e^{-2\eta(t)}\left(\boldsymbol{z}_{m}(t),\boldsymbol{y}^{m}(t)\right)\mathrm{d}t-\int_{0}^{T}e^{-2\eta(t)}\left\langle\boldsymbol{f}_{m},\boldsymbol{y}^{m}(t)\right\rangle\mathrm{d}t.$ (3.43)
Taking limit infimum on both sides of (3.3), we deduce
$\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}(\boldsymbol{\phi}(t))+\eta^{\prime}(t)\boldsymbol{\phi}(t),\boldsymbol{\phi}(t)-\boldsymbol{y}(t)\right>\mathrm{d}t$
$\geq\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}_{0}(t)+\eta^{\prime}(t)\boldsymbol{y}(t),\boldsymbol{\phi}(t)\right>\mathrm{d}t+\frac{1}{2}\liminf_{n\to\infty}\bigg{[}e^{-2\eta(T)}\|\boldsymbol{y}^{m}(T)\|_{2}^{2}-\|\boldsymbol{y}^{m}(0)\|_{2}^{2}\bigg{]}$
$\qquad-\chi\int_{0}^{T}e^{-2\eta(t)}\left(\boldsymbol{z}(t),\boldsymbol{y}(t)\right)\mathrm{d}t-\int_{0}^{T}e^{-2\eta(t)}\left\langle\boldsymbol{f},\boldsymbol{y}(t)\right\rangle\mathrm{d}t$
$\geq\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}_{0}(t)+\eta^{\prime}(t)\boldsymbol{y}(t),\boldsymbol{\phi}(t)\right>\mathrm{d}t+\frac{1}{2}\bigg{[}e^{-2\eta(T)}\|\boldsymbol{y}(T)\|_{2}^{2}-\|\boldsymbol{y}(0)\|_{2}^{2}\bigg{]}$
$\qquad-\chi\int_{0}^{T}e^{-2\eta(t)}\left(\boldsymbol{z}(t),\boldsymbol{y}(t)\right)\mathrm{d}t-\int_{0}^{T}e^{-2\eta(t)}\left\langle\boldsymbol{f},\boldsymbol{y}(t)\right\rangle\mathrm{d}t,$ (3.44)
where we have used the weak lower semicontinuity property of the $\mathbb{H}$-norm and the strong convergence of the initial data (3.41) in the final inequality. Now, using the equality (3.39) in (3.3), we further have
$\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}(\boldsymbol{\phi}(t))+\eta^{\prime}(t)\boldsymbol{\phi}(t),\boldsymbol{\phi}(t)-\boldsymbol{y}(t)\right>\mathrm{d}t$
$\geq\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}_{0}(t)+\eta^{\prime}(t)\boldsymbol{y}(t),\boldsymbol{\phi}(t)\right>\mathrm{d}t-\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}_{0}(t)+\eta^{\prime}(t)\boldsymbol{y}(t),\boldsymbol{y}(t)\right>\mathrm{d}t$
$\geq\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}_{0}(t)+\eta^{\prime}(t)\boldsymbol{y}(t),\boldsymbol{\phi}(t)-\boldsymbol{y}(t)\right>\mathrm{d}t.$ (3.45)
Note that the estimate (3.3) holds true for any $\boldsymbol{\phi}\in\mathrm{L}^{\infty}(0,T;\mathbb{H}_{m})$, $m\in\mathbb{N}$, since the inequality given in (3.3) is independent of both $m$ and $n$. Using a density argument, one can show that the inequality (3.3) remains true for any $\boldsymbol{\phi}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})).$ In fact, for any $\boldsymbol{\phi}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})),$ there exists a strongly convergent subsequence $\boldsymbol{\phi}_{m}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})),$ that satisfies the inequality (3.3).
Taking $\boldsymbol{\phi}=\boldsymbol{y}+r\boldsymbol{w}$, $r>0$, where $\boldsymbol{w}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})),$ and substituting for $\boldsymbol{\phi}$ in (3.3), we get
$\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}(\boldsymbol{y}(t)+r\boldsymbol{w}(t))-\mathscr{G}_{0}(t)+r\eta(t)\boldsymbol{w}(t),r\boldsymbol{w}(t)\right>\mathrm{d}t\geq 0.$ (3.46)
Dividing the inequality (3.46) by $r$, using the hemicontinuity property of the operator $\mathscr{G}(\cdot)$ (see Lemma 2.3), and then passing $r\to 0$, we find
$\int_{0}^{T}e^{-2\eta(t)}\left<\mathscr{G}(\boldsymbol{y}(t))-\mathscr{G}_{0}(t),\boldsymbol{w}(t)\right>\mathrm{d}t\geq 0,$ (3.47)
for any $\boldsymbol{w}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})).$ Therefore, from (3.47), we deduce that $\mathscr{G}(\boldsymbol{y}(\cdot))=\mathscr{G}_{0}(\cdot).$ In addition, $\boldsymbol{y}(\cdot)$ satisfies the energy equality (3.8) for all $t\in[0,T]$. Step III. Uniqueness: Define $\mathscr{Y}=\boldsymbol{y}_{1}-\boldsymbol{y}_{2}$, where $\boldsymbol{y}_{1}$ and $\boldsymbol{y}_{2}$ are two weak solutions of system (3.19) in the sense of Definition 3.2. Then $\mathscr{Y}\in\mathrm{C}([0,T];\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ and satisfies
$\frac{\mathrm{d}\mathscr{Y}(t)}{\mathrm{d}t}&=-\left[\mathscr{G}(\boldsymbol{y}_{1}(t))-\mathscr{G}(\boldsymbol{y}_{2}(t))\right],\\ \mathscr{Y}(0)&=\textbf{0},\end{aligned}\right.$ (3.48)
in the weak sense. Therefore, we have
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\mathscr{Y}(t)\|^{2}_{2}$ =$\mathscr{G}(\boldsymbol{y}_{1}(t))-\mathscr{G}(\boldsymbol{y}_{2}(t)),\boldsymbol{y}_{1}(t)-\boldsymbol{y}_{2}(t)\right>$
$\leq\frac{(C_{S}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})(t)\|_{{4}}^{2}\;\|\mathscr{Y}(t)\|_{2}^{2},$ (3.49)
for a.e. $t\in[0,T]$, where we have used (2.2). An application of the variation of constants formula and the fact that $\mathscr{Y}(0)=\textbf{0}$ give $\boldsymbol{y}_{1}(t)=\boldsymbol{y}_{2}(t)$, for all $t\in[0,T]$ in $\mathbb{H}$, which proves the uniqueness. ∎
Next, we show that the weak solution of system (3.19) is continuous with respect to given data (particularly $\boldsymbol{x}$, $\boldsymbol{f}$ and $\boldsymbol{z}$).
Lemma 3.9.
Let (1.4) be satisfied. For some $T>0$ fixed, assume that $\boldsymbol{x}_{m}\to\boldsymbol{x}$ in $\mathbb{H}$, $\boldsymbol{f}_{m}\to\boldsymbol{f}\ \text{ in }\ \mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\boldsymbol{z}_{m}\to\boldsymbol{z}\ \text{ in }\ \mathrm{L}^{2}(0,T;\mathbb{H})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$. Let us denote by $\boldsymbol{y}(t,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}),$ the solution of system (3.19) and by $\boldsymbol{y}(t,\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m}),$ the solution of system (3.19) with $\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}$ being replaced by $\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m}$. Then
$\boldsymbol{y}(\cdot,\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m})\to\boldsymbol{y}(\cdot,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x})\ \text{ in }\ \mathrm{C}([0,T];\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})).$ (3.50)
In particular, $\boldsymbol{y}(T,\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m})\to\boldsymbol{y}(T,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x})$ in $\mathbb{H}$.
Proof.
Let us introduce the following notations which help us to simplify the proof:
$\boldsymbol{y}_{m}(\cdot)$ $:=\boldsymbol{y}(\cdot,\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m}),\ \ \boldsymbol{y}(\cdot):=\boldsymbol{y}(\cdot,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}),\ \ \mathfrak{F}_{m}(\cdot):=\boldsymbol{y}_{m}(\cdot)-\boldsymbol{y}(\cdot),$
$\hat{\boldsymbol{z}}_{m}(\cdot)$ $:=\boldsymbol{z}_{m}(\cdot)-\boldsymbol{z}(\cdot),\ \ \hat{\boldsymbol{f}}_{m}(\cdot):=\mathscr{P}\boldsymbol{f}_{m}(\cdot)-\mathscr{P}\boldsymbol{f}(\cdot).$
Then $\mathfrak{F}_{m}$ satisfies the following system:
[…]
Multiplying by $\mathfrak{F}_{m}(\cdot)$ to the first equation in (3.51), integrating over $\mathscr{O}$ and using (2.2), we obtain
$\frac{1}{2}$ $\frac{\mathrm{d}}{\mathrm{d}t}\|\mathfrak{F}_{m}(t)\|^{2}_{2}$
=$\mathscr{G}(\boldsymbol{y}_{m}(t))-\mathscr{G}(\boldsymbol{y}(t)),\mathfrak{F}_{m}(t)\right>+\chi(\hat{\boldsymbol{z}}_{m}(t),\ \mathfrak{F}_{m}(t))+\langle\hat{\boldsymbol{f}}_{m},\mathfrak{F}_{m}(t)\rangle$
$\leq-\frac{\nu\varepsilon_{0}}{4}\|\mathrm{A}(\mathfrak{F}_{m}(t))\|^{2}_{2}-\frac{\beta\varepsilon_{0}}{4}\|\mathrm{A}(\mathfrak{F}_{m}(t))\|^{4}_{4}+\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|_{{4}}^{2}\|\mathfrak{F}_{m}(t)\|_{2}^{2}$
$\quad+\chi(\hat{\boldsymbol{z}}_{m}(t),\ \mathfrak{F}_{m}(t))+\langle\hat{\boldsymbol{f}}_{m},\mathfrak{F}_{m}(t)\rangle,$ (3.52)
for a.e. $t\in[0,T]$. Using Hölder’s inequality, (1.8) and Young’s inequality, we have
$|\chi(\hat{\boldsymbol{z}}_{m},\ \mathfrak{F}_{m})|$ $\leq\chi\|\mathfrak{F}_{m}\|_{2}\|\hat{\boldsymbol{z}}_{m}\|_{2}\leq C\|\mathrm{A}(\mathfrak{F}_{m})\|_{2}\|\hat{\boldsymbol{z}}_{m}\|_{2}\leq\frac{\nu\varepsilon_{0}}{16}\|\mathrm{A}(\mathfrak{F}_{m})\|^{2}_{2}+C\|\hat{\boldsymbol{z}}_{m}\|^{2}_{2},.$ (3.53)
Similar to (3.28), we find for $\boldsymbol{f}_{m}=\boldsymbol{f}_{1,m}+\boldsymbol{f}_{2,m}$, $\boldsymbol{f}=\boldsymbol{f}_{1}+\boldsymbol{f}_{2}$ and $\hat{\boldsymbol{f}}_{m}=\hat{\boldsymbol{f}}_{1,m}+\hat{\boldsymbol{f}}_{2,m}$
$\big{|}\langle\hat{\boldsymbol{f}}_{m},\mathfrak{F}_{m}\rangle\big{|}$ $\leq\frac{\nu\varepsilon_{0}}{16}\|\mathrm{A}(\mathfrak{F}_{m})\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{8}\|\mathrm{A}(\mathfrak{F}_{m})\|^{4}_{4}+C\bigg{[}\|\hat{\boldsymbol{f}}_{1,m}\|^{2}_{\mathbb{H}^{-1}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}\bigg{]},$ (3.54)
Combining (3.3) and (3.54), we deduce
$\frac{\mathrm{d}}{\mathrm{d}t}\|\mathfrak{F}_{m}(t)\|^{2}_{2}+\frac{\nu\varepsilon_{0}}{4}\|\mathrm{A}(\mathfrak{F}_{m}(t))\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{4}\|\mathrm{A}(\mathfrak{F}_{m}(t))\|^{4}_{4}$
$\leq\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|_{{4}}^{2}\|\mathfrak{F}_{m}(t)\|_{2}^{2}$
$\qquad+C\|\hat{\boldsymbol{z}}_{m}(t)\|^{2}_{2}+C\bigg{[}\|\hat{\boldsymbol{f}}_{1,m}\|^{2}_{\mathbb{H}^{-1}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}\bigg{]},$
for a.e. $t\in[0,T]$. In view of Gronwall’s inequality, we obtain
$\|\mathfrak{F}_{m}(t)\|^{2}_{2}+\frac{\nu\varepsilon_{0}}{4}\int_{0}^{t}\|\mathrm{A}(\mathfrak{F}_{m}(s))\|^{2}_{2}\mathrm{d}s+\frac{\beta\varepsilon_{0}}{4}\int_{0}^{t}\|\mathrm{A}(\mathfrak{F}_{m}(s))\|^{4}_{4}\mathrm{d}s$
$\leq\biggl{\{}\|\mathfrak{F}_{m}(0)\|^{2}_{2}+C\int_{0}^{t}\bigg{[}\|\hat{\boldsymbol{z}}_{m}(s)\|^{2}_{2}+\|\hat{\boldsymbol{f}}_{1,m}\|^{2}_{\mathbb{H}^{-1}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}\bigg{]}\mathrm{d}s\biggr{\}}$
×$\qquad\times\exp\left\{\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\int_{0}^{t}\|\mathrm{A}(\boldsymbol{y}(s)+\boldsymbol{z}(s))\|_{{4}}^{2}\mathrm{d}s\right\},$ (3.55)
for all $t\in[0,T]$. Since, $\lim\limits_{m\to\infty}\|\mathfrak{F}_{m}(0)\|_{2}=\lim\limits_{m\to\infty}\|\boldsymbol{x}_{m}-\boldsymbol{x}\|_{2}=0$,
$\lim\limits_{m\to\infty}\int_{0}^{T}\bigg{[}\|\hat{\boldsymbol{z}}_{m}(s)\|^{2}_{2}+\|\hat{\boldsymbol{f}}_{1,m}\|^{2}_{\mathbb{H}^{-1}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4}{3}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\hat{\boldsymbol{f}}_{2,m}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}\bigg{]}\mathrm{d}s=0,$
and $\boldsymbol{y}+\boldsymbol{z}\in\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$, then (3.3) asserts that
$\|\mathfrak{F}_{m}(t)\|^{2}_{2}+\frac{\nu\varepsilon_{0}}{4}\int_{0}^{t}\|\mathrm{A}(\mathfrak{F}_{m}(s))\|^{2}_{2}\mathrm{d}s+\frac{\beta\varepsilon_{0}}{4}\int_{0}^{t}\|\mathrm{A}(\mathfrak{F}_{m}(s))\|^{4}_{4}\mathrm{d}s\to 0$
as $m\to\infty$ uniformly in $t\in[0,T].$ Since $\boldsymbol{y}_{m}(\cdot)$ and $\boldsymbol{y}(\cdot)$ are continuous, we further have
$\boldsymbol{y}(\cdot,\boldsymbol{z}_{m},\boldsymbol{f}_{m},\boldsymbol{x}_{m})\to\boldsymbol{y}(\cdot,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x})\ \text{ in }\ \mathrm{C}([0,T];\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})),$
as $m\to\infty$. This completes the proof. ∎
Definition 3.3.
We define a map $\Psi_{\chi}:[0,\infty)\times\Omega\times\mathbb{H}\to\mathbb{H}$ by
$(t,\omega,\boldsymbol{x})\mapsto\boldsymbol{y}^{\chi}(t)+\boldsymbol{z}_{\chi}(\omega)(t)\in\mathbb{H},$ (3.56)
where $\boldsymbol{y}^{\chi}(t)=\boldsymbol{y}(t,\boldsymbol{z}_{\chi}(\omega)(t),\boldsymbol{f},\boldsymbol{x}-\boldsymbol{z}_{\chi}(\omega)(0))$ is a solution to system (3.19) with the initial condition $\boldsymbol{x}-\boldsymbol{z}_{\chi}(\omega)(0).$
Proposition 3.10.
If $\chi_{1},\chi_{2}\geq 0$, then $\Psi_{\chi_{1}}=\Psi_{\chi_{2}}.$
Proof.
Let us fix $\boldsymbol{x}\in\mathbb{H}.$ We need to prove that
$\boldsymbol{y}^{\chi_{1}}(t)+\boldsymbol{z}_{\chi_{1}}(t)=\boldsymbol{y}^{\chi_{2}}(t)+\boldsymbol{z}_{\chi_{2}}(t),\ \ \ t\geq 0,$
where $\boldsymbol{z}_{\chi}$ is defined by (3.8) and $\boldsymbol{y}^{\chi}$ is a solution to system (3.19).
Since $\boldsymbol{z}_{\chi_{1}},\boldsymbol{z}_{\chi_{2}}\in\mathrm{C}([0,T];\mathbb{H}\cap\mathbb{L}^{4}(\mathscr{O}))$ and $\boldsymbol{z}_{\chi_{2}}-\boldsymbol{z}_{\chi_{1}}$ satisfies (3.10), it implies from [49, Vol. II, Theorem 3.2, p.22] that
$\boldsymbol{z}_{\chi_{2}}-\boldsymbol{z}_{\chi_{1}}\in\mathrm{C}([0,T];\mathbb{H}\cap\mathbb{W}^{1,4}(\mathscr{O}))\cap\mathrm{L}^{2}(0,T;\mathbb{V})$ (3.57)
with
$\partial_{t}(\boldsymbol{z}_{\chi_{2}}-\boldsymbol{z}_{\chi_{1}})\in\mathrm{L}^{2}(0,T,\mathbb{V}^{\prime}).$ (3.58)
From (3.19), we infer that $\boldsymbol{y}^{\chi_{1}}(0)-\boldsymbol{y}^{\chi_{2}}(0)=-(\boldsymbol{z}_{\chi_{1}}(0)-\boldsymbol{z}_{\chi_{2}}(0))$ and
$\frac{\mathrm{d}(\boldsymbol{y}^{\chi_{1}}(t)-\boldsymbol{y}^{\chi_{2}}(t))}{\mathrm{d}t}$
$=-\nu\mathscr{A}(\boldsymbol{y}^{\chi_{1}}(t)-\boldsymbol{y}^{\chi_{2}}(t))+\big{[}\chi_{1}\boldsymbol{z}_{\chi_{1}}(t)-\chi_{2}\boldsymbol{z}_{\chi_{2}}(t)\big{]}-[\mathscr{B}(\boldsymbol{y}^{\chi_{1}}(t)+\boldsymbol{z}_{\chi_{1}}(t))-\mathscr{B}(\boldsymbol{y}^{\chi_{2}}(t)+\boldsymbol{z}_{\chi_{2}}(t))]$
$\quad-\alpha[\mathscr{J}(\boldsymbol{y}^{\chi_{1}}(t)+\boldsymbol{z}_{\chi_{1}}(t))-\mathscr{J}(\boldsymbol{y}^{\chi_{2}}(t)+\boldsymbol{z}_{\chi_{2}}(t))]-\beta[\mathscr{K}(\boldsymbol{y}^{\chi_{1}}(t)+\boldsymbol{z}_{\chi_{1}}(t))-\mathscr{K}(\boldsymbol{y}^{\chi_{2}}(t)+\boldsymbol{z}_{\chi_{2}}(t))],$
in $\mathbb{X}^{\prime}$. It implies from Theorem 3.8 that
$\boldsymbol{y}^{\chi_{1}}-\boldsymbol{y}^{\chi_{2}}\in\mathrm{C}([0,T];\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ (3.59)
with
$\partial_{t}(\boldsymbol{y}^{\chi_{1}}-\boldsymbol{y}^{\chi_{2}})\in\mathrm{L}^{2}(0,T,\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})).$ (3.60)
Adding the equation (3.10) to the above equation, we obtain
$\frac{\mathrm{d}(\boldsymbol{v}^{\chi_{1}}(t)-\boldsymbol{v}^{\chi_{2}}(t))}{\mathrm{d}t}$ $=-\nu\mathscr{A}(\boldsymbol{v}^{\chi_{1}}(t)-\boldsymbol{v}^{\chi_{2}}(t))-[\mathscr{B}(\boldsymbol{v}^{\chi_{1}}(t))-\mathscr{B}(\boldsymbol{v}^{\chi_{2}}(t))]$
$\quad-\alpha[\mathscr{J}(\boldsymbol{v}^{\chi_{1}}(t))-\mathscr{J}(\boldsymbol{v}^{\chi_{2}}(t))]-\beta[\mathscr{K}(\boldsymbol{v}^{\chi_{1}}(t))-\mathscr{K}(\boldsymbol{v}^{\chi_{2}}(t))],$ (3.61)
in $\mathbb{X}^{\prime}$, where $\boldsymbol{v}^{\chi_{1}}(t)=\boldsymbol{y}^{\chi_{1}}(t)+\boldsymbol{z}_{\chi_{1}}(t),\boldsymbol{v}^{\chi_{2}}(t)=\boldsymbol{y}^{\chi_{2}}(t)+\boldsymbol{z}_{\chi_{2}}(t),\ t\geq 0$ and $\boldsymbol{v}^{\chi_{1}}(0)-\boldsymbol{v}^{\chi_{2}}(0)=\boldsymbol{0}.$ In view of (3.57)-(3.58) and (3.59)-(3.60), we infer that
$\boldsymbol{v}^{\chi_{1}}-\boldsymbol{v}^{\chi_{2}}\in\mathrm{C}([0,T];\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$
with
$\partial_{t}(\boldsymbol{v}^{\chi_{1}}-\boldsymbol{v}^{\chi_{2}})\in\mathrm{L}^{2}(0,T,\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})).$
Therefore, from [21, Chapter II, Theorem 1.8], we have
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\boldsymbol{v}^{\chi_{1}}(t)-\boldsymbol{v}^{\chi_{2}}(t)\|^{2}_{2}$
$=-\big{\langle}\nu\mathscr{A}(\boldsymbol{v}^{\chi_{1}}(t)-\boldsymbol{v}^{\chi_{2}}(t))+[\mathscr{B}(\boldsymbol{v}^{\chi_{1}}(t))-\mathscr{B}(\boldsymbol{v}^{\chi_{2}}(t))]-\alpha[\mathscr{J}(\boldsymbol{v}^{\chi_{1}}(t))-\mathscr{J}(\boldsymbol{v}^{\chi_{2}}(t))]$
$\qquad\qquad\qquad-\beta[\mathscr{K}(\boldsymbol{v}^{\chi_{1}}(t))-\mathscr{K}(\boldsymbol{v}^{\chi_{2}}(t))],\boldsymbol{v}^{\chi_{1}}(t)-\boldsymbol{v}^{\chi_{2}}(t)\big{\rangle},$
for a.e. $t\in[0,T]$. Applying similar steps as we have performed for (2.2), we arrive at
$\frac{\mathrm{d}}{\mathrm{d}t}\|\boldsymbol{v}^{\chi_{1}}(t)-\boldsymbol{v}^{\chi_{2}}(t)\|^{2}_{2}\leq\frac{2(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{v}^{\chi_{2}}(t))\|_{{4}}^{2}\|\boldsymbol{v}^{\chi_{1}}(t)-\boldsymbol{v}^{\chi_{2}}(t)\|^{2}_{\mathbb{H}},$ (3.62)
for a.e. $t\in[0,T]$. Since $\|\boldsymbol{v}^{\chi_{1}}(0)-\boldsymbol{v}^{\chi_{2}}(0)\|^{2}_{2}=0$ and $\boldsymbol{v}^{\chi_{2}}\in\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$, by applying Gronwall inequality, we deduce that $\|\boldsymbol{v}^{\chi_{1}}(t)-\boldsymbol{v}^{\chi_{2}}(t)\|^{2}_{2}=0$, for all $t\geq 0$, which completes the proof. ∎
It is proved in Proposition 3.10 that the map $\Psi_{\chi}$ does not depend on $\chi$ and hence, from now onward, it will be denoted by $\Psi$. A proof of the following result is similar to that in [12, Theorem 6.15] which is based on the uniqueness of the solutions to system (3.19). Hence we omit it here.
Theorem 3.11.
$(\Psi,\theta)$ is an RDS.
Now we are ready to present the solution to system (2.18) with the initial data $\boldsymbol{x}\in\mathbb{H}$ at the initial time $s\in\mathbb{R}.$ The following theorem is a consequence of our previous discussion.
Theorem 3.12.
In the framework of Definition 2.1, suppose that $\boldsymbol{v}(t)=\boldsymbol{y}^{\chi}(t)+\boldsymbol{z}_{\chi}(t),t\geq s,$ where $\boldsymbol{y}^{\chi}$ is the unique solution to system (3.19) with initial data $\boldsymbol{x}-\boldsymbol{z}_{\chi}(s)$ at time $s$. If the process $\{\boldsymbol{v}(t),\ t\geq s\},$ has trajectories in $\mathrm{C}([s,\infty);\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}([s,\infty);\mathbb{W}^{1,4}(\mathscr{O}))$, then it is a solution to system (2.18). Vice-versa, if a process $\{\boldsymbol{v}(t),t\geq s\},$ with trajectories in $\mathrm{C}([s,\infty);\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}([s,\infty);\mathbb{W}^{1,4}(\mathscr{O}))$ is a solution to system (2.18), then for any $\chi\geq 0,$ a process $\{\boldsymbol{y}^{\chi}(t),t\geq s\},$ defined by $\boldsymbol{y}^{\chi}(t)=\boldsymbol{v}(t)-\boldsymbol{z}_{\chi}(t),t\geq s,$ is a solution to (3.19) on $[s,\infty).$
4. Random attractors for stochastic third-grade fluids equations: Bounded domains
In this section, we prove the existence of unique random attractor for stochastic third-grade fluids equation 2.18 on unbounded domains. In order to prove the result of this section, we assume that external forcing term $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$. Here, the RDS $\Psi$ is considered over the MDS $(\Omega,{\mathscr{F}},{\mathbb{P}},{\theta})$. The results that we have obtained in the previous sections provide a unique solution to system (2.18), which is continuous with respect to the data (particularly $\boldsymbol{x}$ and $\boldsymbol{f}$). Furthermore, if we define, for $\boldsymbol{x}\in\mathbb{H},\ \omega\in\Omega,$ and $t\geq s,$
$\boldsymbol{v}(t,\omega;s,\boldsymbol{x}):=\Psi(t-s,\theta_{s}\omega,\boldsymbol{x})=\boldsymbol{y}\big{(}t,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s)\big{)}+\boldsymbol{z}(t),$ (4.1)
then the process $\{\boldsymbol{v}(t):\ t\geq s\}$ is a solution to equation (2.18)1italic-(2.18subscriptitalic-)$_{1}$ for $t>s$, for each $s\in\mathbb{R}$ and each $\boldsymbol{x}\in\mathbb{H}$.
The following lemma helps us to prove the existence of random $\mathfrak{DK}$-absorbing set.
Lemma 4.1.
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Suppose that $\boldsymbol{y}$ solves system (3.19) on the time interval $[a,\infty)$ with $\boldsymbol{z}\in\mathrm{L}^{2}_{\mathrm{loc}}(\mathbb{R};\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}(\mathbb{R};\mathbb{W}^{1,4}(\mathscr{O}))$, $\chi\geq 0$ and $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ (that is, $\boldsymbol{f}=\boldsymbol{f}_{1}+\boldsymbol{f}_{2}$ with $\boldsymbol{f}_{1}\in\mathbb{H}^{-1}(\mathscr{O})$ and $\boldsymbol{f}_{2}\in\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$). Then, for any $t\geq\tau\geq a,$
$\|\boldsymbol{y}(t)\|^{2}_{2}$ $\leq\|\boldsymbol{y}(\tau)\|^{2}_{2}\ e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-\tau)}$
$\quad+C\int_{\tau}^{t}\bigg{[}\|\boldsymbol{z}(s)\|^{2}_{2}+\|\boldsymbol{z}(s)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+1\bigg{]}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-s)}\mathrm{d}s.$ (4.2)
Proof.
From (3.19), we obtain
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\boldsymbol{y}(t)\|^{2}_{2}$
$=-\frac{\nu}{2}\|\mathrm{A}(\boldsymbol{y}(t))\|^{2}_{2}-\left<\mathscr{B}(\boldsymbol{y}(t)+\boldsymbol{z}(t))+\alpha\mathscr{J}(\boldsymbol{y}(t)+\boldsymbol{z}(t))+\beta\mathscr{K}(\boldsymbol{y}(t)+\boldsymbol{z}(t)),\boldsymbol{y}(t)\right>+\chi(\boldsymbol{z}(t),\boldsymbol{y}(t))$
$\quad+\big{\langle}\boldsymbol{f},\boldsymbol{y}(t)\big{\rangle}$
$=-\frac{\nu}{2}\|\mathrm{A}(\boldsymbol{y}(t))\|^{2}_{2}-\frac{\beta}{2}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|^{4}_{4}+b(\boldsymbol{y}(t)+\boldsymbol{z}(t),\boldsymbol{y}(t),\boldsymbol{z}(t))-\alpha\left<\mathscr{J}(\boldsymbol{y}(t)+\boldsymbol{z}(t)),\boldsymbol{y}(t)\right>$
$\quad+\beta\left<\mathscr{K}(\boldsymbol{y}(t)+\boldsymbol{z}(t)),\boldsymbol{z}(t)\right>+\chi(\boldsymbol{z}(t),\boldsymbol{y}(t))+\big{\langle}\boldsymbol{f},\boldsymbol{y}(t)\big{\rangle},$ (4.3)
for a.e. $t\in[a,\infty)$. Next, using the Hölder’s inequality, Lemma 2.1, Korn-type inequality (2.2), Poincaré inequality (1.8), (1.4) (see Remark 1.1 also), Gagliardo-Nirenberg inequality (2.1) and Young’s inequality, we estimate the terms of the right hand side of (3.3) as (similar to (3.24)-(3.28))
$|b(\boldsymbol{y}+\boldsymbol{z},\boldsymbol{y},\boldsymbol{z})|$ $\leq\frac{\nu\varepsilon_{0}}{24}\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{12}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+C\|\boldsymbol{z}\|_{4}^{4},$ (4.4)
$\left|\alpha\left<\mathscr{J}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{y}\right>\right|$ $\leq\frac{\nu(1-\varepsilon_{0})}{4}\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\frac{\beta(1-\varepsilon_{0})}{2}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}$ (4.5)
$\left|\beta\left<\mathscr{K}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{z}\right>\right|$ $\leq\leq\frac{\beta\varepsilon_{0}}{12}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+C\|\mathrm{A}(\boldsymbol{z})\|^{4}_{4},$ (4.6)
$|\chi(\boldsymbol{z},\boldsymbol{y})|$ $\leq\frac{\nu\varepsilon_{0}}{24}\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+C\|\boldsymbol{z}\|^{2}_{2},$ (4.7)
and
$\big{|}\langle\boldsymbol{f},\boldsymbol{y}\rangle\big{|}$ $\leq\frac{\nu\varepsilon_{0}}{24}\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{16}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+C\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\mathrm{A}(\boldsymbol{z})\|^{4}_{4}+1\bigg{]}.$ (4.8)
Hence, from (4) and (1.8), we deduce
$\frac{\mathrm{d}}{\mathrm{d}t}\|\boldsymbol{y}(t)\|^{2}_{2}$ $\leq-\frac{\nu}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\|\mathrm{A}(\boldsymbol{y}(t))\|^{2}_{2}-\frac{\beta\varepsilon_{0}}{2}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|^{4}_{4}$
$\quad+C\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}+1\bigg{]}$
$=-\nu\left(1+\frac{\varepsilon_{0}}{2}\right)\|\nabla\boldsymbol{y}(t)\|^{2}_{2}-\frac{\beta\varepsilon_{0}}{2}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|^{4}_{4}$
$\quad+C\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}+1\bigg{]}$
$\leq-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\|\boldsymbol{y}(t)\|^{2}_{2}-\frac{\beta\varepsilon_{0}}{2}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|^{4}_{4}$
$\quad+C\bigg{[}\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}+1\bigg{]},$
for a.e. $t\in[a,\infty)$, and an application of the variation of constants formula yields (4.1). ∎
In the following lemma, we show that the RDS $\Psi$ is compact operator from $\mathbb{H}$ into itself.
Lemma 4.2.
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Then, the solution operator $\Psi$ given in Definition (4.1) is compact from $\mathbb{H}$ into itself.
Proof.
Consider the solution $\Psi(t-s,\theta_{s}\omega,\cdot)=\boldsymbol{y}\big{(}t,\omega;s,\cdot-\boldsymbol{z}(s)\big{)}+\boldsymbol{z}(t)$ of (2.18) for all $t>s$ and $s\in\mathbb{R}$. For simplicity, we write $\boldsymbol{y}\big{(}t,s;\omega,\cdot-\boldsymbol{z}(s)\big{)}=\boldsymbol{y}(t,s,\cdot)$. Assume that the sequence $\{\boldsymbol{x}_{m}\}_{m\in\mathbb{N}}\subset\mathbb{H}$ is bounded. We have that, for $\boldsymbol{z}\in\mathrm{C}([s,t];\mathbb{H})\cap\mathrm{L}^{4}(s,t;\mathbb{W}^{1,4}(\mathscr{O}))$, the sequences
$\{\boldsymbol{y}(\cdot,s,\boldsymbol{x}_{m})\}_{m\in\mathbb{N}}\subset\mathrm{C}([s,t];\mathbb{H})\cap\mathrm{L}^{2}(s,t;\mathbb{V})\cap\mathrm{L}^{4}(s,t;\mathbb{W}^{1,4}(\mathscr{O}))$
and
$\left\{\frac{\mathrm{d}}{\mathrm{d}t}[\boldsymbol{y}(\cdot,s,\boldsymbol{x}_{m})]\right\}_{m\in\mathbb{N}}\subset\mathrm{L}^{2}(s,t;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(s,t;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$
are bounded. Since $\mathrm{L}^{2}(s,t;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(s,t;\mathbb{W}^{-1,\frac{4}{3}})\subset\mathrm{L}^{\frac{4}{3}}(s,t;\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}})$, the above sequence is bounded in $\mathrm{L}^{\frac{4}{3}}(s,t;\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$. Note also that $\mathbb{V}\subset\mathbb{H}\subset\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}}$ and the embedding of $\mathbb{V}\subset\mathbb{H}$ is compact. By the Aubin-Lions compactness lemma, there exists a subsequence (keeping as it is) and $\widetilde{\boldsymbol{y}}\in\mathrm{L}^{2}(s,t;\mathbb{H})$ such that
$\boldsymbol{y}(\cdot,s,\boldsymbol{x}_{m})\to\widetilde{\boldsymbol{y}}(\cdot)\ \text{ strongly in }\ \mathrm{L}^{2}(s,t;\mathbb{H}),$ (4.9)
as $m\to\infty$. Again, choosing one more subsequence (again not relabeling), we infer from (4.9) that
$\boldsymbol{y}(\tau,s,\boldsymbol{x}_{m})\to\widetilde{\boldsymbol{y}}(\tau)\text{ in }\mathbb{H},\ \text{ for a.e. }\ \tau\in[s,t],$ (4.10)
as $m\to\infty$. For $s<t$, we obtain from (4.10) that there exists $\tau\in(s,t)$ such that (4.10) holds true for this particular $\tau$. Then by Lemma 3.9, we obtain
$\Psi(t-s,\theta_{s}\omega,\boldsymbol{x}_{m})$ $=\Psi(t-\tau,\theta_{\tau}\omega,\Psi(\tau-s,\theta_{s}\omega,\boldsymbol{x}_{m}))$
$=\Psi\big{(}t-\tau,\theta_{\tau}\omega,\boldsymbol{y}\big{(}\tau,\omega;s,\boldsymbol{x}_{m}-\boldsymbol{z}(s)\big{)}+\boldsymbol{z}(\tau)\big{)}$
$\to\Psi\big{(}t-\tau,\theta_{\tau}\omega,\widetilde{\boldsymbol{y}}(\tau)+\boldsymbol{z}(\tau)\big{)}$
as $m\to\infty$, in $\mathbb{H}$, which completes the proof. ∎
Let us now provide the main result of this section.
Theorem 4.3.
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Consider the MDS, $\Im=(\Omega,\mathscr{F},\mathbb{P},\theta)$ from Proposition 3.1, and the RDS $\Psi$ on $\mathbb{H}$ over $\Im$ generated by the SGMNS equations (2.18) with additive noise satisfying Hypothesis 1.1. Then, there exists a unique random $\mathfrak{DK}$-attractor for continuous RDS $\Psi$ in $\mathbb{H}$.
Proof.
Because of [13, Theorem 2.8], it is only needed to prove that there exists a $\mathfrak{DK}$-absorbing set $\textbf{B}\in\mathfrak{DK}$ and the RDS $\Psi$ is $\mathfrak{DK}$-asymptotically compact. Existence of $\mathfrak{DK}$-absorbing set $\textbf{B}\in\mathfrak{DK}$: Let $\mathrm{D}$ be a random set from the class $\mathfrak{DK}$. Let $\kappa_{\mathrm{D}}(\omega)$ be the radius of $\mathrm{D}(\omega)$, that is, $\kappa_{\mathrm{D}}(\omega):=\sup\{\|x\|_{2}:x\in\mathrm{D}(\omega)\}$ for $\omega\in\Omega$.
Let $\omega\in\Omega$ be fixed. For a given $s\leq 0$ and $\boldsymbol{x}\in\mathbb{H}$, let $\boldsymbol{y}$ be the solution of (3.19) on the time interval $[s,\infty)$ with the initial condition $\boldsymbol{y}(s)=\boldsymbol{x}-\boldsymbol{z}(s).$ Using (4.1) for $t=\xi\text{ and }\tau=s\leq-1$ with $\xi\in[-1,0]$, we obtain
$\|\boldsymbol{y}(\xi)\|^{2}_{2}$ $\leq 2\|\boldsymbol{x}\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}+2\|\boldsymbol{z}(s)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}$
$\quad+C\int_{s}^{0}\bigg{\{}\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+1\bigg{\}}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\mathrm{d}t.$ (4.11)
For $\omega\in\Omega,$ let us set
$[\kappa_{11}(\omega)]^{2}$ $=2+2\sup_{s\leq 0}\bigg{\{}\|\boldsymbol{z}(s)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}\bigg{\}}$
$\quad+C\int_{-\infty}^{0}\bigg{\{}\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}_{1}\|^{2}_{\mathbb{H}^{-1}}+\|\boldsymbol{f}_{2}\|^{\frac{4(4+d)}{8+3d}}_{\mathbb{W}^{-1,\frac{4}{3}}}+1\bigg{\}}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\mathrm{d}t,$ (4.12)
$\kappa_{12}(\omega)$ $=\|\boldsymbol{z}(\omega)(0)\|_{2}.$ (4.13)
In view of Lemma 3.6 and Proposition 3.7, we deduce that both $\kappa_{11},\kappa_{12}\in\mathfrak{K}$ and also that $\kappa_{11}+\kappa_{12}=:\kappa_{13}\in\mathfrak{K}$ as well. Therefore the random set B defined by
$\textbf{B}(\omega):=\{\boldsymbol{v}\in\mathbb{H}:\|\boldsymbol{v}\|_{2}\leq\kappa_{13}(\omega)\}$
is such that $\textbf{B}\in\mathfrak{DK}.$
Let us now show that B absorbs $\mathrm{D}$. Let $\omega\in\Omega$ be fixed. Since $\kappa_{\mathrm{D}}(\omega)\in\mathfrak{K}$, there exists $t_{\mathrm{D}}(\omega)\geq 1$ such that
$[\kappa_{\mathrm{D}}(\theta_{-t}\omega)]^{2}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}$ $\leq 1,\ \text{ for }\ t\geq t_{\mathrm{D}}(\omega).$
Thus, for $\omega\in\Omega$, if $\boldsymbol{x}\in\mathrm{D}(\theta_{s}\omega)$ and $s\leq-t_{\mathrm{D}}(\omega)\leq-1,$ then by (4), we obtain for $\xi\in[-1,0]$
$\|\boldsymbol{y}(\xi,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\|_{2}\leq\kappa_{11}(\omega),\ \text{ for }\ \omega\in\Omega.$ (4.14)
Thus, we conclude that, for $\omega\in\Omega$
$\|\boldsymbol{v}(0,\omega;s,\boldsymbol{x})\|_{2}\leq\|\boldsymbol{y}(0,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\|_{2}+\|\boldsymbol{z}(\omega)(0)\|_{2}\leq\kappa_{13}(\omega).$
The above inequality implies that for $\omega\in\Omega$, $\boldsymbol{v}(0,\omega;s,\boldsymbol{x})\in\textbf{B}(\omega)$, for all $s\leq-t_{\mathrm{D}}(\omega).$ This proves B absorbs $\mathrm{D}$. The RDS $\Psi$ is $\mathfrak{DK}$-asymptotically compact. In order to establish the $\mathfrak{DK}$-asymptotically compactness of $\Psi$, we use the compact Sobolev embedding $\mathbb{V}\subset\mathbb{H}$. Let us assume that $\mathrm{D}\in\mathfrak{DK}$ and $\textbf{B}\in\mathfrak{DK}$ be such that B absorbs $\mathrm{D}$. Let us fix $\omega\in\Omega$ and take a sequence of positive numbers $\{t_{m}\}^{\infty}_{m=1}$ such that $t_{1}\leq t_{2}\leq t_{3}\leq\cdots$ and $t_{m}\to\infty$. We take an arbitrary $\mathbb{H}$-valued sequence $\{\boldsymbol{x}_{m}\}^{\infty}_{m=1}$ such that $\boldsymbol{x}_{m}\in\mathrm{D}(\theta_{-t_{m}}\omega),$ for all $m\in\mathbb{N}.$
Since $t_{m}\to\infty$, there exists $M_{1}=M_{1}(\tau,\hat{K})\in\mathbb{N}$ such that $t_{m}\geq t_{\mathrm{D}}(\omega)$ for all $m\geq M_{1}.$ Therefore, in view of (4.14), the sequence
$\{\Psi(t_{m}-1,\theta_{-t_{m}}\omega,\boldsymbol{x}_{m})\}_{m\geq M_{1}}=\{\boldsymbol{v}(-1,\omega;-t_{m},\boldsymbol{x}_{m})\}_{m\geq M_{1}}\subset\mathbb{H},$ (4.15)
is bounded. Combining the fact of (4.15) and Lemma 4.2, we conclude that the sequence
$\{\Psi(t_{m},\theta_{-t_{m}}\omega,\boldsymbol{x}_{m})\}_{m\geq M_{1}}=\{\Psi(1,\theta_{-1}\omega,\Psi(t_{m}-1,\theta_{-t_{m}}\omega,\boldsymbol{x}_{m}))\}_{m\geq M_{1}},$ (4.16)
has a convergent subsequence. Hence, The RDS $\Psi$ is $\mathfrak{DK}$-asymptotically compact. This completes the proof. ∎
5. Random attractors for stochastic third-grade fluids equations: Unbounded domains
In this section, we prove the existence of unique random attractor for stochastic third-grade fluids equation 2.18 on unbounded domains. In order to prove the result of this section, we assume that external forcing term $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$. The following two lemmas help us to prove the existence of random $\mathfrak{DK}$-absorbing set.
Lemma 5.1.
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Suppose that $\boldsymbol{y}$ solves system (3.19) on the time interval $[a,\infty)$ with $\boldsymbol{z}\in\mathrm{L}^{2}_{\mathrm{loc}}(\mathbb{R};\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}(\mathbb{R};\mathbb{W}^{1,4}(\mathscr{O}))$, $\chi\geq 0$ and $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$. Then, for any $t\geq s\geq a,$
$\|\boldsymbol{y}(t)\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{2}\int_{s}^{t}\|\mathrm{A}(\boldsymbol{y}(\tau)+\boldsymbol{z}(\tau))\|^{4}_{4}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-\tau)}\mathrm{d}\tau$
$\leq\|\boldsymbol{y}(s)\|^{2}_{2}\ e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-s)}+C\int_{s}^{t}\bigg{[}\|\boldsymbol{z}(\tau)\|^{2}_{2}+\|\boldsymbol{z}(\tau)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{]}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-\tau)}\mathrm{d}\tau.$ (5.1)
Proof.
From (3.19), we obtain
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\boldsymbol{y}(t)\|^{2}_{2}$ $=-\frac{\nu}{2}\|\mathrm{A}(\boldsymbol{y}(t))\|^{2}_{2}-\frac{\beta}{2}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|^{4}_{4}+b(\boldsymbol{y}(t)+\boldsymbol{z}(t),\boldsymbol{y}(t),\boldsymbol{z}(t))$
$\quad-\alpha\left<\mathscr{J}(\boldsymbol{y}(t)+\boldsymbol{z}(t)),\boldsymbol{y}(t)\right>+\beta\left<\mathscr{K}(\boldsymbol{y}(t)+\boldsymbol{z}(t)),\boldsymbol{z}(t)\right>+(\chi\boldsymbol{z}(t)+\boldsymbol{f},\boldsymbol{y}(t)),$ (5.2)
for a.e. $t\in[a,\infty)$. Next, using the Hölder’s inequality, Lemma 2.1, Korn-type inequality (2.2), Poincaré inequality (1.8), (1.4) (see Remark 1.1 also), Gagliardo-Nirenberg inequality (2.1) and Young’s inequality, we estimate the terms of the right hand side of (3.3) as (see the proof of Lemma 4.1)
$|b(\boldsymbol{y}+\boldsymbol{z},\boldsymbol{y},\boldsymbol{z})|$ $\leq\frac{\nu\varepsilon_{0}}{16}\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{8}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+C\|\boldsymbol{z}\|_{4}^{4},$ (5.3)
$\left|\alpha\left<\mathscr{J}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{y}\right>\right|$ $\leq\frac{\nu(1-\varepsilon_{0})}{4}\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\frac{\beta(1-\varepsilon_{0})}{2}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}$ (5.4)
$\left|\beta\left<\mathscr{K}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{z}\right>\right|$ $\leq\frac{\beta\varepsilon_{0}}{8}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+C\|\mathrm{A}(\boldsymbol{z})\|^{4}_{4},$ (5.5)
$|(\chi\boldsymbol{z}+\boldsymbol{f},\boldsymbol{y})|$ $\leq\frac{\nu\varepsilon_{0}}{16}\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+C\|\boldsymbol{z}\|^{2}_{2}+C\|\boldsymbol{f}\|^{2}_{2}.$ (5.6)
Hence, from (5) and (1.8), we deduce
$\frac{\mathrm{d}}{\mathrm{d}t}\|\boldsymbol{y}(t)\|^{2}_{2}$
$\leq-\frac{\nu}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\|\mathrm{A}(\boldsymbol{y}(t))\|^{2}_{2}-\frac{\beta\varepsilon_{0}}{2}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|^{4}_{4}+C\bigg{[}\|\boldsymbol{f}\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}\bigg{]}$ (5.7)
$\leq-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\|\boldsymbol{y}(t)\|^{2}_{2}-\frac{\beta\varepsilon_{0}}{2}\|\mathrm{A}(\boldsymbol{y}(t)+\boldsymbol{z}(t))\|^{4}_{4}+C\bigg{[}\|\boldsymbol{f}\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}\bigg{]},$
for a.e. $t\in[a,\infty)$, and an application of the variation of constants formula yields (5.1). This completes the proof. ∎
Lemma 5.2.
Let (1.4) and Hypotheses 1.1-1.2 be satisfied, and $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$. Then, for every $\omega\in\Omega$ and $\mathrm{D}\in\mathfrak{DK}$, and for all $\xi\in[-1,0]$, there exists $t_{\mathrm{D}}(\omega)\geq 1$ and a constant $C_{ubd}>0$ such that for all $s\leq-t_{\mathrm{D}}(\omega),$
$e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\xi}\|\boldsymbol{y}(\xi,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{2}\int_{s}^{\xi}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s)))+\boldsymbol{z}(\tau)\|^{4}_{4}\mathrm{d}\tau$
$\leq 2+2\sup_{t\leq 0}\bigg{\{}\|\boldsymbol{z}(t)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\bigg{\}}+C_{ubd}\int_{-\infty}^{0}\bigg{\{}\|\boldsymbol{z}(\tau)\|^{2}_{2}+\|\boldsymbol{z}(\tau)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau,$ (5.8)
for any $\boldsymbol{x}\in\mathrm{D}(\theta_{s}\omega)$. In addition, there exist $t_{\mathrm{D}}^{\ast}(\omega)$ and $t_{\mathrm{D}}^{\ast\ast}(\omega)$ such that
$\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s)))\|^{2}_{2}\mathrm{d}\tau\leq C,\text{ for all }\;s\leq-t_{\mathrm{D}}^{\ast}(\omega),$ (5.9)
for any $\boldsymbol{x}\in\mathrm{D}(\theta_{s}\omega)$, and
$\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\|\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\|^{4}_{2}\mathrm{d}\tau\leq C,\text{ for all }\;s\leq-t_{\mathrm{D}}^{\ast\ast}(\omega),$ (5.10)
for any $\boldsymbol{x}\in\mathrm{D}(\theta_{s}\omega)$.
Proof.
Let $\mathrm{D}$ be a random set from the class $\mathfrak{DK}$. Let $\kappa_{\mathrm{D}}(\omega)$ be the radius of $\mathrm{D}(\omega)$, that is, $\kappa_{\mathrm{D}}(\omega):=\sup\{\|x\|_{2}:x\in\mathrm{D}(\omega)\}$ for $\omega\in\Omega$. Let $\omega\in\Omega$ be fixed. Since $\kappa_{\mathrm{D}}(\omega)\in\mathfrak{K}$, there exists $t_{\mathrm{D}}(\omega)\geq 1$ such that
$[\kappa_{\mathrm{D}}(\theta_{-t}\omega)]^{2}e^{-\frac{\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)t}\leq 1,$ (5.11)
for $t\geq t_{\mathrm{D}}(\omega).$ From (5.1), for $t=\xi\in[-1,0]$ and $s\leq-t_{\mathrm{D}}(\omega)$, we obtain
$e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\xi}\|\boldsymbol{y}(\xi,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\|^{2}_{2}+\frac{\beta\varepsilon_{0}}{2}\int_{s}^{\xi}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s)))+\boldsymbol{z}(\tau)\|^{4}_{4}\mathrm{d}\tau$
$\leq\|\boldsymbol{x}-\boldsymbol{z}(s)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}+C\int_{s}^{\xi}\bigg{[}\|\boldsymbol{z}(\tau)\|^{2}_{2}+\|\boldsymbol{z}(\tau)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{]}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\leq 2\|\boldsymbol{x}\|^{2}_{2}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}+2\|\boldsymbol{z}(s)\|^{2}_{2}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}$
$\quad+C_{ubd}\int_{-\infty}^{0}\bigg{\{}\|\boldsymbol{z}(\tau)\|^{2}_{2}+\|\boldsymbol{z}(\tau)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\leq 2+2\sup_{s\leq 0}\bigg{\{}\|\boldsymbol{z}(s)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}\bigg{\}}+C_{ubd}\int_{-\infty}^{0}\bigg{\{}\|\boldsymbol{z}(\tau)\|^{2}_{2}+\|\boldsymbol{z}(\tau)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau,$ (5.12)
where we have used (5.11). Note that, in view of Lemmas 3.5 and 3.6, the right hand side of (5) is finite. This completes the proof of (5.2).
Form (5.7), we have
$\frac{\mathrm{d}}{\mathrm{d}t}\|\boldsymbol{y}(t)\|^{2}_{2}+\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\|\boldsymbol{y}(t)\|^{2}_{2}+\frac{\nu}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\|\mathrm{A}(\boldsymbol{y}(t))\|^{2}_{2}$
$\leq C\bigg{[}\|\boldsymbol{f}\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{2}_{2}+\|\boldsymbol{z}(t)\|^{4}_{\mathbb{W}^{1,4}}\bigg{]}+\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\|\boldsymbol{y}(t)\|^{2}_{2},$
which implies
$\frac{\nu}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\int_{s}^{t}\|\mathrm{A}(\boldsymbol{y}(\tau))\|^{2}_{2}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-\tau)}\mathrm{d}\tau$
$\leq\|\boldsymbol{y}(s)\|^{2}_{2}\ e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-s)}+C\int_{s}^{t}\bigg{[}\|\boldsymbol{z}(\tau)\|^{2}_{2}+\|\boldsymbol{z}(\tau)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{]}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-\tau)}\mathrm{d}\tau$
$\quad+\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\int_{s}^{t}\|\boldsymbol{y}(\tau)\|^{2}_{2}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-\tau)}\mathrm{d}\tau$
$\leq\|\boldsymbol{y}(s)\|^{2}_{2}\ e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-s)}+C\int_{s}^{t}\bigg{[}\|\boldsymbol{z}(\tau)\|^{2}_{2}+\|\boldsymbol{z}(\tau)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{]}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-\tau)}\mathrm{d}\tau$
$\quad+\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\int_{s}^{t}\bigg{[}\|\boldsymbol{y}(s)\|^{2}_{2}\ e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(\tau-s)}$
$\qquad+C\int_{s}^{\tau}\bigg{\{}\|\boldsymbol{z}(\zeta)\|^{2}_{2}+\|\boldsymbol{z}(\zeta)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(\tau-\zeta)}\mathrm{d}\zeta\bigg{]}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(t-\tau)}\mathrm{d}\tau.$ (5.13)
where we have used (5.1) in the final inequality. Putting $t=0$, in (5), we get
$\frac{\nu}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s)))\|^{2}_{2}\mathrm{d}\tau$
$\leq\|\boldsymbol{x}-\boldsymbol{z}(s)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}+C\int_{s}^{0}\bigg{[}\|\boldsymbol{z}(\tau)\|^{2}_{2}+\|\boldsymbol{z}(\tau)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{]}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\quad+\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\int_{s}^{0}\bigg{[}\|\boldsymbol{x}-\boldsymbol{z}(s)\|^{2}_{2}\ e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(\tau-s)}$
$\qquad+C\int_{s}^{\tau}\bigg{\{}\|\boldsymbol{z}(\zeta)\|^{2}_{2}+\|\boldsymbol{z}(\zeta)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(\tau-\zeta)}\mathrm{d}\zeta\bigg{]}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\leq\|\boldsymbol{x}-\boldsymbol{z}(s)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}+C\int_{s}^{0}\bigg{[}\|\boldsymbol{z}(\tau)\|^{2}_{2}+\|\boldsymbol{z}(\tau)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{]}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\quad+\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\int_{s}^{0}\bigg{[}\|\boldsymbol{x}-\boldsymbol{z}(s)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}$
$\qquad+C\int_{s}^{\tau}\bigg{\{}\|\boldsymbol{z}(\zeta)\|^{2}_{2}+\|\boldsymbol{z}(\zeta)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\zeta}\mathrm{d}\zeta\bigg{]}\mathrm{d}\tau$
$\leq\|\boldsymbol{x}-\boldsymbol{z}(s)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}+C\int_{-\infty}^{0}\bigg{\{}\|\boldsymbol{z}(\tau)\|^{2}_{2}+\|\boldsymbol{z}(\tau)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\quad-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\|\boldsymbol{x}-\boldsymbol{z}(s)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}s$
$\qquad+C\int_{s}^{0}e^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\int_{s}^{\tau}\bigg{\{}\|\boldsymbol{z}(\zeta)\|^{2}_{2}+\|\boldsymbol{z}(\zeta)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\zeta}\mathrm{d}\zeta\mathrm{d}\tau$
$\leq 2\left[\|\boldsymbol{x}\|^{2}_{2}e^{\frac{\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)s}+\|\boldsymbol{z}(s)\|^{2}_{2}e^{\frac{\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)s}\right](1-s)e^{\frac{3\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)s}$
$\qquad+C\int_{-\infty}^{0}\bigg{\{}\|\boldsymbol{z}(\zeta)\|^{2}_{2}+\|\boldsymbol{z}(\zeta)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\zeta}\mathrm{d}\zeta.$ (5.14)
Using (5.11), Lemmas 3.5 and 3.6, and the fact that $\lim\limits_{s\to-\infty}(1-s)e^{\frac{3\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)s}=0,$ we have that there exists a $t^{\ast}_{\mathrm{D}}(\omega)\geq t_{\mathrm{D}}(\omega)$ such that for all $s\leq-t^{\ast}_{\mathrm{D}}(\omega)$
$\frac{\nu}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s)))\|^{2}_{2}\mathrm{d}\tau$
$\leq 2+2\sup_{s\leq 0}\bigg{\{}\|\boldsymbol{z}(s)\|^{2}_{2}\ e^{\frac{\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)s}\bigg{\}}+C\int_{-\infty}^{0}\bigg{\{}\|\boldsymbol{z}(\tau)\|^{2}_{2}+\|\boldsymbol{z}(\tau)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau.$ (5.15)
Again, in view of Lemmas 3.5 and 3.6, the right hand side of (5) is finite. This completes the proof of (5.9).
Finally, from (5.1), we have
$\int_{s}^{0}\|\boldsymbol{y}(\tau)\|^{4}_{2}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\leq\int_{s}^{0}\bigg{[}\|\boldsymbol{y}(s)\|^{2}_{2}\ e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(\tau-s)}$
$\qquad+C\int_{s}^{\tau}\bigg{\{}\|\boldsymbol{z}(\zeta)\|^{2}_{2}+\|\boldsymbol{z}(\zeta)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(\tau-\zeta)}\mathrm{d}\zeta\bigg{]}^{2}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\leq 2\int_{s}^{0}\|\boldsymbol{y}(s)\|^{4}_{2}\ e^{-2\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(\tau-s)}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\qquad+2C\int_{s}^{0}\bigg{[}\int_{s}^{\tau}\bigg{\{}\|\boldsymbol{z}(\zeta)\|^{2}_{2}+\|\boldsymbol{z}(\zeta)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(\tau-\zeta)}\mathrm{d}\zeta\bigg{]}^{2}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\leq-2s\|\boldsymbol{y}(s)\|^{4}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}$
$\qquad+2C\int_{s}^{0}\bigg{[}\int_{s}^{\tau}\bigg{\{}\|\boldsymbol{z}(\zeta)\|^{2}_{2}+\|\boldsymbol{z}(\zeta)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(\tau-\zeta)}\mathrm{d}\zeta\bigg{]}^{2}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau,$
which gives
$\int_{s}^{0}\|\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\|^{4}_{2}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\leq-2s\|\boldsymbol{x}-\boldsymbol{z}(s)\|^{4}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}$
$\qquad+2C\int_{s}^{0}\bigg{[}\int_{s}^{\tau}\bigg{\{}\|\boldsymbol{z}(\zeta)\|^{2}_{2}+\|\boldsymbol{z}(\zeta)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)(\tau-\zeta)}\mathrm{d}\zeta\bigg{]}^{2}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\leq-16s\left\{\|\boldsymbol{x}\|^{4}_{2}+\|\boldsymbol{z}(s)\|^{4}_{2}\right\}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}$
$\qquad+2C\int_{s}^{0}\bigg{[}\int_{s}^{\tau}\bigg{\{}\|\boldsymbol{z}(\zeta)\|^{2}_{2}+\|\boldsymbol{z}(\zeta)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{-\frac{\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)(\tau-\zeta)}\mathrm{d}\zeta\bigg{]}^{2}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\leq 16\left\{\|\boldsymbol{x}\|^{4}_{2}e^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)s}+\|\boldsymbol{z}(s)\|^{4}_{2}e^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)s}\right\}\ \left\{-se^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)s}\right\}$
$\qquad+2C\int_{s}^{0}\bigg{[}\int_{s}^{\tau}\bigg{\{}\|\boldsymbol{z}(\zeta)\|^{2}_{2}+\|\boldsymbol{z}(\zeta)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{\frac{\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)\zeta}\mathrm{d}\zeta\bigg{]}^{2}e^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau.$
Using (5.11), Lemmas 3.5 and 3.6, and the fact that $\lim\limits_{s\to-\infty}\left\{-se^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)s}\right\}=0,$ we have that there exists a $t^{\ast\ast}_{\mathrm{D}}(\omega)\geq t_{\mathrm{D}}(\omega)$ such that for all $s\leq-t^{\ast\ast}_{\mathrm{D}}(\omega)$
$\int_{s}^{0}\|\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\|^{4}_{2}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau$
$\leq 16+16\sup_{s\leq 0}\left\{\|\boldsymbol{z}(s)\|^{4}_{2}e^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)s}\right\}+C\bigg{[}\int_{-\infty}^{0}\bigg{\{}\|\boldsymbol{z}(\zeta)\|^{2}_{2}+\|\boldsymbol{z}(\zeta)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{\frac{\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)\zeta}\mathrm{d}\zeta\bigg{]}^{2},$
which is finite due to Lemmas 3.5 and 3.6. This completes the proof of (5.10). ∎
The following result is used to obtain the uniform-tail estimates for the solution of system (3.18) (see Lemma 5.4 below).
Lemma 5.3.
Let (1.4) and Hypotheses 1.1-1.2 be satisfied and let us denote by $\boldsymbol{y}(t,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}),$ the solution of system (3.19). Assume that $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\{\boldsymbol{x}_{m}\}_{m\in\mathbb{N}}$ be a bounded sequence in $\mathbb{H}$. Then, there exists $\tilde{\boldsymbol{y}}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ such that along a subsequence
$\boldsymbol{y}(\cdot,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}_{m})\ \to\tilde{\boldsymbol{y}}\ \text{ in }\ \mathrm{L}^{2}(0,T;\mathbb{L}^{2}_{\mathrm{loc}}(\mathscr{O})),$ (5.16)
for every $T>0$.
Proof.
Let $\boldsymbol{y}_{m}(\cdot)=\boldsymbol{y}(\cdot,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}_{m})$. Since $\{\boldsymbol{x}_{m}\}_{m\in\mathbb{N}}$ is a bounded sequence in $\mathbb{H}$, the sequence
$\{\boldsymbol{y}_{m}\}_{n\in\mathbb{N}}\ \text{ is bounded in }\ \mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})).$ (5.17)
Hence, there exists a subsequence $\{\boldsymbol{y}_{m^{\prime}}\}_{m^{\prime}\in\mathbb{N}}$ of $\{\boldsymbol{y}_{m}\}_{m\in\mathbb{N}}$ and
$\widetilde{\boldsymbol{y}}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})),$
such that, as $m^{\prime}\to\infty,$ (by the Banach-Alaoglu theorem)
[…]
From Theorem 3.8, we have $\big{\|}\frac{\mathrm{d}\boldsymbol{y}_{m}}{\mathrm{d}t}\big{\|}_{\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{V}^{\prime}+\mathbb{W}^{-1,\frac{4}{3}})}\leq C,$ for some $C>0$ and all $m\in\mathbb{N}.$ Therefore, by the Cauchy-Schwarz inequality, for all $0\leq t\leq t+a\leq T$, $m\in\mathbb{N}$ and $\boldsymbol{\varphi}\in\mathbb{V}\cap\mathbb{W}^{1,4}(\mathscr{O})$, we obtain
$|(\boldsymbol{y}_{m}(t+a)-\boldsymbol{y}_{m}(t),\boldsymbol{\varphi})|\leq\int_{t}^{t+a}\bigg{|}\bigg{\langle}\frac{\mathrm{d}\boldsymbol{y}_{m}(s)}{\mathrm{d}t},\boldsymbol{\varphi}\bigg{\rangle}\bigg{|}\mathrm{d}s\leq C(T)a^{\frac{1}{4}}\|\boldsymbol{\varphi}\|_{\mathbb{V}\cap\mathbb{W}^{1,4}}.$ (5.19)
Since $\boldsymbol{y}_{m}(t+a)-\boldsymbol{y}_{m}(t)\in\mathbb{V}\cap\mathbb{W}^{1,4}(\mathscr{O})$, for a.e. $t\in(0,T)$, choosing $\boldsymbol{\varphi}=\boldsymbol{y}_{m}(t+a)-\boldsymbol{y}_{m}(t)$ in (5.19), we obtain
$\|\boldsymbol{y}_{m}(t+a)-\boldsymbol{y}_{m}(t)\|_{2}^{2}\leq C(T)a^{\frac{1}{4}}\|\boldsymbol{y}_{m}(t+a)-\boldsymbol{y}_{m}(t)\|_{\mathbb{V}\cap\mathbb{W}^{1,4}}.$ (5.20)
Integrating from 00 to $T-a$, we further find
$\int_{0}^{T-a}\|\boldsymbol{y}_{m}(t+a)-\boldsymbol{y}_{m}(t)\|_{2}^{2}\mathrm{d}t$ $\leq C(T)a^{\frac{1}{4}}\int_{0}^{T-a}\|\boldsymbol{y}_{m}(t+a)-\boldsymbol{y}_{m}(t)\|_{\mathbb{V}\cap\mathbb{W}^{1,4}}\mathrm{d}t$
$\leq C(T)a^{\frac{1}{4}}\bigg{[}(T-a)^{1/2}\left(\int_{0}^{T-a}\|\boldsymbol{y}_{m}(t+a)-\boldsymbol{y}_{m}(t)\|_{\mathbb{V}}^{2}\mathrm{d}t\right)^{1/2}$
$\qquad+(T-a)^{\frac{3}{4}}\left(\int_{0}^{T-a}\|\boldsymbol{y}_{m}(t+a)-\boldsymbol{y}_{m}(t)\|_{\mathbb{W}^{1,4}}^{4}\mathrm{d}t\right)^{\frac{1}{4}}\bigg{]}$
$\leq\widetilde{C}(T)a^{\frac{1}{4}},$ (5.21)
where we have used the Hölder’s inequality and (5.17). Also, $\widetilde{C}(T)$ is an another positive constant independent of $m$. Furthermore, we have
$\lim_{a\to 0}\sup_{m}\int_{0}^{T-a}\|\boldsymbol{y}_{m}(t+a)-\boldsymbol{y}_{m}(t)\|_{2}^{2}\mathrm{d}t=0.$ (5.22)
Let us now consider a truncation function $\chi\in\mathrm{C}^{1}(\mathbb{R}^{+})$ with $\chi(s)=1$ for $s\in[0,1]$ and $\chi(s)=0$ for $s\in[4,\infty)$. For each $k>0$, let us define
$\boldsymbol{y}_{m,k}(x):=\chi\left(\frac{|x|^{2}}{k^{2}}\right)\boldsymbol{y}_{m}(x),\ \text{ for }x\in\mathscr{O}_{2k},$
where $\mathscr{O}_{k}=\{x\in\mathscr{O}:|x|<k\}$. It can easily be seen from (5.22) that
$\lim_{a\to 0}\sup_{m}\int_{0}^{T-a}\|\boldsymbol{y}_{m,k}(t+a)-\boldsymbol{y}_{m,k}(t)\|_{\mathbb{L}^{2}(\mathscr{O}_{2k})}^{2}\mathrm{d}t=0,$ (5.23)
for all $T,k>0$. Moreover, from (5.17), for all $T,k>0$, we infer
$\{\boldsymbol{y}_{m,k}\}_{m\in\mathbb{N}}\ \text{ is bounded in }\ \mathrm{L}^{\infty}(0,T;\mathbb{L}^{2}(\mathscr{O}_{2k}))\cap\mathrm{L}^{2}(0,T;\mathbb{H}^{1}_{0}(\mathscr{O}_{2k})).$ (5.24)
Since the injection $\mathbb{H}_{0}^{1}(\mathscr{O}_{2k})\subset\mathbb{L}^{2}{(\mathscr{O}_{2k})}$ is compact, we can apply [50, Theorem 16.3] (see [63, Theorem 13.3] also) to obtain
$\{\boldsymbol{y}_{m,k}\}_{m\in\mathbb{N}}\ \text{ is relatively compact in }\ \mathrm{L}^{2}(0,T;\mathbb{L}^{2}(\mathscr{O}_{2k})),$ (5.25)
for all $T,k>0$. From (5.25), we further infer
$\{\boldsymbol{y}_{m}\}_{m\in\mathbb{N}}\ \text{ is relatively compact in }\ \mathrm{L}^{2}(0,T;\mathbb{L}^{2}(\mathscr{O}_{k})),$ (5.26)
for all $T,k>0$. Using (5.18) and (5.26), we can extract a subsequence of $\{\boldsymbol{y}_{m}\}_{m\in\mathbb{N}}$ (not relabeling) such that
$\boldsymbol{y}_{m}\to\tilde{\boldsymbol{y}}\ \text{ in }\mathrm{L}^{2}(0,T;\mathbb{L}^{2}(\mathscr{O}_{k})),$
for all $T,k>0$. This completes the proof. ∎
Next, we show that the solution to system (3.18) satisfies the uniform-tail estimates which will help us to establish the $\mathfrak{DK}$-asymptotic compactness of $\Psi$ on unbounded domains.
Let $\Lambda$ be a smooth function such that $0\leq\Lambda(\xi)\leq 1$ for $\xi\in\mathbb{R}$ and
$\Lambda(\xi)=\begin{cases*}0,\text{ for }|\xi|\leq 1,\\ 1,\text{ for }|\xi|\geq 2.\end{cases*}$ (5.27)
Then, there exists a positive constant $C^{\ast}$ such that $|\Lambda^{\prime}(\xi)|\leq C^{\ast}$ and $|\Lambda^{\prime\prime}(\xi)|\leq C^{\ast}$ for all $\xi\in\mathbb{R}$.
Lemma 5.4.
Let (1.4) and Hypotheses 1.1-1.2 be satisfied, $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$, $\omega\in\Omega$ and $\mathrm{D}\in\mathfrak{DK}$. Then, for each $\varepsilon>0$ and for each $\xi\in[-1,0]$, there exists $k_{0}:=k_{0}(\varepsilon,\xi,\omega,\mathrm{D})\in\mathbb{N}$ such that the solution $\boldsymbol{y}(\cdot)$ of system (3.18) satisfies
$\left\|\boldsymbol{y}(\xi,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\right\|^{2}_{\mathbb{L}^{2}(\mathscr{O}^{c}_{k})}\leq\varepsilon,$ (5.28)
for all $\boldsymbol{x}\in\mathrm{D}(\theta_{s}\omega)$, for all $s\leq-\mathscr{T}=-\max\{t_{\mathrm{D}}^{\ast}(\omega),t_{\mathrm{D}}^{\ast\ast}(\omega)\}$ and for all $k\geq k_{0}$, where $\mathscr{O}_{k}=\{x\in\mathscr{O}:\,|x|<k\}$, $\mathscr{O}^{c}_{k}=\mathscr{O}\backslash\mathscr{O}_{k}$, and $t_{\mathrm{D}}^{\ast}(\omega)$ and $t_{\mathrm{D}}^{\ast\ast}(\omega)$ are the same time given in Lemma 5.2.
Proof.
Taking the divergence of the first equation of (3.18), we obtain
$-\Delta\widehat{\textbf{P}}$
$=\nabla\cdot\left[\big{(}(\boldsymbol{y}+\boldsymbol{z})\cdot\nabla\big{)}(\boldsymbol{y}+\boldsymbol{z})+\alpha\text{div}((\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2})+\beta\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))\right]-\nabla\cdot\boldsymbol{f}$
=[$=\nabla\cdot\left[\nabla\cdot\left\{\big{(}(\boldsymbol{y}+\boldsymbol{z})\otimes(\boldsymbol{y}+\boldsymbol{z})\big{)}+\alpha(\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}\right\}\right]+\beta\nabla\cdot\left[\nabla\cdot\left\{|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\right\}\right]-\nabla\cdot\boldsymbol{f}$
$=:-\Delta\widehat{\textbf{P}}_{1}-\Delta\widehat{\textbf{P}}_{2},$
in the weak sense, where
$-\Delta\widehat{\textbf{P}}_{1}$ =[$=\nabla\cdot\left[\nabla\cdot\left\{\big{(}(\boldsymbol{y}+\boldsymbol{z})\otimes(\boldsymbol{y}+\boldsymbol{z})\big{)}+\alpha(\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}\right\}\right]-\nabla\cdot\boldsymbol{f},$
$-\Delta\widehat{\textbf{P}}_{2}$ $=\beta\nabla\cdot\left[\nabla\cdot\left\{|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\right\}\right].$
Note that, by Lemma 5.2, we have for $s\leq-\mathscr{T}=\max\{t_{\mathrm{D}}^{\ast}(\omega),t_{\mathrm{D}}^{\ast\ast}(\omega)\}$
$e^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\cdot}\;\bigg{(}\nabla\cdot\left[\nabla\cdot\left\{\big{(}(\boldsymbol{y}+\boldsymbol{z})\otimes(\boldsymbol{y}+\boldsymbol{z})\big{)}+\alpha(\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}\right\}\right]-\nabla\cdot\boldsymbol{f}\bigg{)}$ $\in\mathrm{L}^{2}(s,0;\mathbb{H}^{-2}(\mathscr{O}))$
and
$e^{\frac{3\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)\cdot}\;\bigg{(}\beta\nabla\cdot\left[\nabla\cdot\left\{|\mathrm{A}(\boldsymbol{v})|^{2}\mathrm{A}(\boldsymbol{v})\right\}\right]\bigg{)}$ $\in\mathrm{L}^{\frac{4}{3}}(s,0;\mathbb{W}^{-2,\frac{4}{3}}(\mathscr{O}))$
which gives
$e^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\cdot}\;\widehat{\textbf{P}}_{1}\in\mathrm{L}^{2}(s,0;\mathbb{L}^{2}(\mathscr{O}))\;\text{ and }\;e^{\frac{3\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)\cdot}\;\widehat{\textbf{P}}_{2}\in\mathrm{L}^{\frac{4}{3}}(s,0;\mathbb{L}^{\frac{4}{3}}(\mathscr{O})).$
Also note that an application of Poincaré inequality (1.8) allow us to write
$\lambda\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)\left|\boldsymbol{y}\right|^{2}\mathrm{d}x$ $\leq\int_{\mathscr{O}}\left|\nabla\left(\Lambda\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\right)\right|^{2}\mathrm{d}x$
$\leq\frac{1}{2}\int_{\mathscr{O}}\left|\mathrm{A}\left(\Lambda\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\right)\right|^{2}\mathrm{d}x$
$=\frac{1}{2}\int_{\mathscr{O}}\left|\Lambda\left(|x|^{2}k^{-2}\right)\mathrm{A}(\boldsymbol{y})+\frac{2}{k^{2}}\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)x^{T}\boldsymbol{y}+\frac{2}{k^{2}}\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}^{T}x\right|^{2}\mathrm{d}x$
$\leq\frac{1}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x+\frac{C}{k}\int_{\mathscr{O}}\left[|\mathrm{A}(\boldsymbol{y})||\boldsymbol{y}|+|\boldsymbol{y}|^{2}\right]\mathrm{d}x$
$\leq\frac{1}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x+\frac{C}{k}\|\mathrm{A}(\boldsymbol{y})\|_{2}^{2},$
which implies
$-\frac{1}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x\leq-\lambda\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)\left|\boldsymbol{y}\right|^{2}\mathrm{d}x+\frac{C}{k}\|\mathrm{A}(\boldsymbol{y})\|_{2}^{2}.$ (5.29)
Taking the inner product of the first equation of (3.18) with $\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}$ in $\mathbb{L}^{2}(\mathscr{O})$, we have
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{y}|^{2}\mathrm{d}x$
$=\underbrace{\nu\int_{\mathscr{O}}(\Delta\boldsymbol{y})\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\mathrm{d}x}_{I_{1}(k,t)}-\underbrace{b\left(\boldsymbol{y}+\boldsymbol{z},\boldsymbol{y}+\boldsymbol{z},\Lambda^{2}\left(|x|^{2}k^{-2}\right)(\boldsymbol{y}+\boldsymbol{z})\right)}_{I_{2}(k,t)}+\underbrace{b\left(\boldsymbol{y}+\boldsymbol{z},\boldsymbol{y}+\boldsymbol{z},\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{z}\right)}_{I_{3}(k,t)}$
$\quad+\underbrace{\alpha\int_{\mathscr{O}}(\text{div}((\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}))\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\mathrm{d}x}_{I_{4}(k,t)}+\underbrace{\beta\int_{\mathscr{O}}(\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})))\Lambda^{2}\left(|x|^{2}k^{-2}\right)(\boldsymbol{y}+\boldsymbol{z})\mathrm{d}x}_{I_{5}(k,t)}$
$\quad-\underbrace{\beta\int_{\mathscr{O}}(\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})))\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{z}\mathrm{d}x}_{I_{6}(k,t)}-\underbrace{\int_{\mathscr{O}}(\nabla\widehat{\textbf{P}})\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\mathrm{d}x}_{I_{7}(k,t)}$
$\quad+\underbrace{\chi\int_{\mathscr{O}}\boldsymbol{z}\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\mathrm{d}x}_{I_{8}(k,t)}+\underbrace{\int_{\mathscr{O}}\boldsymbol{f}\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\mathrm{d}x}_{I_{9}(k,t)}.$ (5.30)
Let us now estimate each term on the right hand side of (5). Integration by parts, (1.4) (see Remark 1.1 also), and Hölder’s, Gagliardo-Nirenberg’s (2.1) and Young’s inequalities help us to obtain
$I_{1}(k,t)$ $=-\frac{\nu}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x-\frac{4\nu}{k^{2}}\int\limits_{\mathscr{O}}\Lambda\left(|x|^{2}k^{-2}\right)\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)[(x\cdot\nabla)\boldsymbol{y}\cdot\boldsymbol{y}+(\boldsymbol{y}\cdot\nabla)\boldsymbol{y}\cdot x]\mathrm{d}x$
$\leq-\frac{\nu}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x+\frac{4\sqrt{2}\nu}{k}\int\limits_{\mathscr{O}\cap\{k\leq|x|\leq\sqrt{2}k\}}\left|\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)\right|\left|\boldsymbol{y}\right|\left|\nabla\boldsymbol{y}\right|\mathrm{d}x$
$\leq-\frac{\nu}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x+\frac{C}{k}\int_{\mathscr{O}}\left|\boldsymbol{y}\right|\left|\nabla\boldsymbol{y}\right|\mathrm{d}x$
$\leq-\frac{\nu}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x+\frac{C}{k}\left(\|\boldsymbol{y}\|^{2}_{2}+\|\nabla\boldsymbol{y}\|^{2}_{2}\right)$
$\leq-\frac{\nu}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x+\frac{C}{k}\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2},$ (5.31)
$-I_{2}(k,t)$ $=\left|\frac{2}{k^{2}}\int_{\mathscr{O}}\Lambda\left(|x|^{2}k^{-2}\right)\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)x\cdot(\boldsymbol{y}+\boldsymbol{z})|\boldsymbol{y}+\boldsymbol{z}|^{2}\mathrm{d}x\right|$
$=\left|\frac{2}{k^{2}}\int\limits_{\mathscr{O}\cap\{k\leq|x|\leq\sqrt{2}k\}}\Lambda\left(|x|^{2}k^{-2}\right)\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)x\cdot(\boldsymbol{y}+\boldsymbol{z})|\boldsymbol{y}+\boldsymbol{z}|^{2}\mathrm{d}x\right|$
$\leq\frac{2\sqrt{2}}{k}\int\limits_{\mathscr{O}\cap\{k\leq|x|\leq\sqrt{2}k\}}\left|\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)\right||\boldsymbol{y}+\boldsymbol{z}|^{3}\mathrm{d}x$
$\leq\frac{C}{k}\|\boldsymbol{y}+\boldsymbol{z}\|^{3}_{3}\leq\frac{C}{k}\left[\|\boldsymbol{y}+\boldsymbol{z}\|^{4}_{4}+\|\boldsymbol{y}+\boldsymbol{z}\|^{2}_{2}\right]$
$\leq\frac{C}{k}\left[\|\boldsymbol{y}\|^{4}_{4}+\|\boldsymbol{z}\|^{4}_{4}+\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\|\boldsymbol{z}\|^{2}_{2}\right]$
$\leq\frac{C}{k}\left[\|\mathrm{A}(\boldsymbol{y})\|^{4}_{4}+\|\boldsymbol{y}\|^{4}_{2}+\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\|\boldsymbol{z}\|^{4}_{4}+\|\boldsymbol{z}\|^{2}_{2}\right]$
$\leq\frac{C}{k}\left[\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\boldsymbol{y}\|^{4}_{2}+\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\|\boldsymbol{z}\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{z}\|^{2}_{2}\right],$ (5.32)
$I_{3}(k,t)$ $\leq\int_{\mathscr{O}}|\boldsymbol{y}+\boldsymbol{z}||\nabla(\boldsymbol{y}+\boldsymbol{z})|\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{z}|\mathrm{d}x$
$\leq\|\boldsymbol{y}+\boldsymbol{z}\|_{2}\|\nabla(\boldsymbol{y}+\boldsymbol{z})\|_{4}\left(\int_{\mathscr{O}}\Lambda^{8}\left(|x|^{2}k^{-2}\right)|\boldsymbol{z}|^{4}\mathrm{d}x\right)^{\frac{1}{4}}$
$\leq C\{\|\mathrm{A}(\boldsymbol{y})\|_{2}+\|\boldsymbol{z}\|_{2}\}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|_{4}\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{z}|^{4}\mathrm{d}x\right)^{\frac{1}{4}},$ (5.33)
$I_{4}(k,t)$ $\leq\left|\frac{\alpha}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)[(\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}:\mathrm{A}(\boldsymbol{y})]\mathrm{d}x\right|$
$\qquad+\left|\frac{4\alpha}{k^{2}}\int_{\mathscr{O}}\Lambda\left(|x|^{2}k^{-2}\right)\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)(\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}:(x^{T}\boldsymbol{y})\mathrm{d}x\right|$
$\leq\frac{\nu(1-\varepsilon_{0})}{4}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x+\frac{\beta(1-\varepsilon_{0})}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{4}\mathrm{d}x$
$\quad+\frac{C}{k}\left[\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\boldsymbol{y}\|^{2}_{2}\right]$
$\leq\frac{\nu(1-\varepsilon_{0})}{4}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x+\frac{\beta(1-\varepsilon_{0})}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{4}\mathrm{d}x$
$\quad+\frac{C}{k}\left[\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}\right],$ (5.34)
$I_{5}(k,t)$ $=-\frac{\beta}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{4}\mathrm{d}x$
$\qquad-\frac{4\beta}{k^{2}}\int_{\mathscr{O}}\Lambda\left(|x|^{2}k^{-2}\right)\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)[|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}):(x^{T}(\boldsymbol{y}+\boldsymbol{z}))]\mathrm{d}x$
$=-\frac{\beta}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{4}\mathrm{d}x$
$\qquad-\frac{4\beta}{k^{2}}\int\limits_{\mathscr{O}\cap\{k\leq|x|\leq\sqrt{2}k\}}\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)[|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}):(x^{T}(\boldsymbol{y}+\boldsymbol{z}))]\mathrm{d}x$
$\leq-\frac{\beta}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{4}\mathrm{d}x+\frac{C}{k}\int\limits_{\mathscr{O}\cap\{k\leq|x|\leq\sqrt{2}k\}}|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{3}|\boldsymbol{y}+\boldsymbol{z}|\mathrm{d}x$
$\leq-\frac{\beta}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{4}\mathrm{d}x+\frac{C}{k}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{3}_{4}\|\boldsymbol{y}+\boldsymbol{z}\|_{4}$
$\leq-\frac{\beta}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{4}\mathrm{d}x+\frac{C}{k}\left[\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\boldsymbol{y}\|^{4}_{4}+\|\boldsymbol{z}\|^{4}_{4}\right]$
$\leq-\frac{\beta}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{4}\mathrm{d}x+\frac{C}{k}\left[\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\boldsymbol{y}\|^{4}_{2}+\|\boldsymbol{z}\|^{4}_{\mathbb{W}^{1,4}}\right],$ (5.35)
$-I_{6}(k,t)$ $\leq\beta\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{3}|\nabla\boldsymbol{z}|\mathrm{d}x+\frac{4\beta}{k^{2}}\int_{\mathscr{O}}\Lambda\left(|x|^{2}k^{-2}\right)\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{3}|x||\boldsymbol{z}|\mathrm{d}x$
$\leq C\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{3}_{4}\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\nabla\boldsymbol{z}|^{4}\mathrm{d}x\right)^{\frac{1}{4}}+\frac{C}{k}\int\limits_{\mathscr{O}\cap\{k\leq|x|\leq\sqrt{2}k\}}|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{3}|\boldsymbol{z}|\mathrm{d}x$
$\leq C\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{3}_{4}\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\nabla\boldsymbol{z}|^{4}\mathrm{d}x\right)^{\frac{1}{4}}+\frac{C}{k}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{3}_{4}\|\boldsymbol{z}\|_{4}$
$\leq C\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{3}_{4}\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\nabla\boldsymbol{z}|^{4}\mathrm{d}x\right)^{\frac{1}{4}}+\frac{C}{k}\left[\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\boldsymbol{z}\|^{4}_{4}\right],$ (5.36)
$-I_{7}(k,t)$ $\leq\left|\frac{4}{k^{2}}\int_{\mathscr{O}}\widehat{\textbf{P}}\;\Lambda\left(|x|^{2}k^{-2}\right)\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)(x\cdot\boldsymbol{y})\mathrm{d}x\right|$
$\leq\frac{C}{k}\int\limits_{\mathscr{O}\cap\{k\leq|x|\leq\sqrt{2}k\}}\left|\widehat{\textbf{P}}\right|\left|\boldsymbol{y}\right|\mathrm{d}x$
$\leq\frac{C}{k}\int_{\mathscr{O}}\bigg{\{}\left|(-\Delta)^{-1}\nabla\cdot\left[\nabla\cdot\big{(}(\boldsymbol{y}+\boldsymbol{z})\otimes(\boldsymbol{y}+\boldsymbol{z})\big{)}\right]\right|+\left|(-\Delta)^{-1}\nabla\cdot\left[\nabla\cdot\left\{(\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}\right\}\right]\right|$
$\qquad\qquad+\left|(-\Delta)^{-1}\nabla\cdot\left[\nabla\cdot\left\{|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\right\}\right]\right|+\left|(-\Delta)^{-1}\left[\nabla\cdot\boldsymbol{f}\right]\right|\bigg{\}}\left|\boldsymbol{y}\right|\mathrm{d}x$
$\leq\frac{C}{k}\left[\|\boldsymbol{y}+\boldsymbol{z}\|^{2}_{4}\|\boldsymbol{y}\|_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{2}_{4}\|\boldsymbol{y}\|_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{3}_{4}\|\boldsymbol{y}\|_{4}+\|\boldsymbol{f}\|_{2}\|\boldsymbol{y}\|_{2}\right]$
$\leq\frac{C}{k}\left[\|\boldsymbol{y}\|^{4}_{4}+\|\boldsymbol{z}\|^{4}_{4}+\|\boldsymbol{y}\|^{2}_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\boldsymbol{f}\|^{2}_{2}\right]$
$\leq\frac{C}{k}\left[\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\|\boldsymbol{y}\|^{4}_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\boldsymbol{f}\|^{2}_{2}+\|\boldsymbol{z}\|^{4}_{\mathbb{W}^{1,4}}\right],$ (5.37)
$I_{8}(k,t)$ $\leq\frac{\nu\lambda\varepsilon_{0}}{8}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{v}|^{2}\mathrm{d}x+C\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{f}|^{2}\mathrm{d}x,$ (5.38)
$I_{9}(k,t)$ $\leq\frac{\nu\lambda\varepsilon_{0}}{8}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{v}|^{2}\mathrm{d}x+C\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{z}|^{2}\mathrm{d}x.$ (5.39)
Combining (5)-(5.39) and using (5.29), we reach at
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{y}|^{2}\mathrm{d}x+\frac{\beta\varepsilon_{0}}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{4}\mathrm{d}x$
$\leq-\frac{\nu\lambda}{2}(1+\varepsilon_{0})\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{v}|^{2}\mathrm{d}x+\frac{C}{k}\|\mathrm{A}(\boldsymbol{v})\|^{2}_{2}+\frac{\nu\lambda\varepsilon_{0}}{4}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{v}|^{2}\mathrm{d}x$
$\quad+C\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{f}|^{2}\mathrm{d}x+C\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{z}|^{2}\mathrm{d}x$
$\quad+C\{\|\mathrm{A}(\boldsymbol{y})\|_{2}+\|\boldsymbol{z}\|_{2}\}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|_{4}\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{z}|^{4}\mathrm{d}x\right)^{\frac{1}{4}}+C\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{3}_{4}\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\nabla\boldsymbol{z}|^{4}\mathrm{d}x\right)^{\frac{1}{4}}$
$\quad+\frac{C}{k}\left[\|\boldsymbol{y}\|^{4}_{2}+\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\boldsymbol{z}\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{z}\|^{2}_{2}+\|\boldsymbol{f}\|^{2}_{2}\right],$
which implies
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{v}|^{2}\mathrm{d}x+\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{v}|^{2}\mathrm{d}x+\beta\varepsilon_{0}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{v})|^{4}\mathrm{d}x$
$\leq C\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{f}|^{2}\mathrm{d}x+C\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{z}|^{2}\mathrm{d}x+C\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{3}_{4}\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\nabla\boldsymbol{z}|^{4}\mathrm{d}x\right)^{\frac{1}{4}}$
$\quad+C\bigg{\{}\|\mathrm{A}(\boldsymbol{y})\|_{2}+\|\boldsymbol{z}\|_{2}\bigg{\}}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|_{4}\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{z}|^{4}\mathrm{d}x\right)^{\frac{1}{4}}+\frac{C}{k}\bigg{[}\|\boldsymbol{y}\|^{4}_{2}+\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}$
$\qquad+\|\boldsymbol{z}\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{z}\|^{2}_{2}+\|\boldsymbol{f}\|^{2}_{2}\bigg{]}.$ (5.40)
An application of variation of constants formula to (5.37) gives for all $\omega\in\Omega$, for all $\xi\in[-1,0]$ and for all $s\leq-\mathscr{T}=-\max\{t_{\mathrm{D}}^{\ast}(\omega),t_{\mathrm{D}}^{\ast\ast}(\omega)\}$
$e^{-\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{y}(\xi,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))|^{2}\mathrm{d}x$
$\leq e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{x}(x)-\boldsymbol{z}(x,s)|^{2}\mathrm{d}x+C\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{f}(x)|^{2}\mathrm{d}x\mathrm{d}\tau$
$\quad+C\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{z}(x,\tau)|^{2}\mathrm{d}x\mathrm{d}\tau+\frac{C}{k}\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\bigg{[}\|\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\|^{4}_{2}$
$\qquad+\|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s)))\|^{2}_{2}+\|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))+\boldsymbol{z}(\tau))\|^{4}_{4}+\|\boldsymbol{z}(\tau)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{z}(\tau)\|^{2}_{2}+\|\boldsymbol{f}\|^{2}_{2}\bigg{]}\mathrm{d}\tau$
$\quad+C\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))+\boldsymbol{z}(\tau))\|^{3}_{4}\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\nabla\boldsymbol{z}(x,\tau)|^{4}\mathrm{d}x\right)^{\frac{1}{4}}\mathrm{d}\tau$
$\quad+C\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\{\|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s)))\|_{2}+\|\boldsymbol{z}(\tau)\|_{2}\}\|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))+\boldsymbol{z}(\tau))\|_{4}$
×$\qquad\times\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{z}(x,\tau)|^{4}\mathrm{d}x\right)^{\frac{1}{4}}\mathrm{d}\tau$
$\leq e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)s}\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{x}(x)-\boldsymbol{z}(x,s)|^{2}\mathrm{d}x+C\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{f}(x)|^{2}\mathrm{d}x\mathrm{d}\tau$
$\quad+C\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{z}(x,\tau)|^{2}\mathrm{d}x\mathrm{d}\tau+\frac{C}{k}+C\left(\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\int_{\mathscr{O}\cap\{|x|\geq k\}}|\nabla\boldsymbol{z}(x,\tau)|^{4}\mathrm{d}x\mathrm{d}\tau\right)^{\frac{1}{4}}$
$\quad+C\left(\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{z}(x,\tau)|^{4}\mathrm{d}x\mathrm{d}\tau\right)^{\frac{1}{4}}$
$\to 0\text{ as }k\to\infty,$ (5.41)
where we have used the finiteness of integral obtained in (5.2)-(5.10) and Lemmas 3.5-3.6. Hence, from (5), we conclude that for any $\varepsilon>0$, for any $\omega\in\Omega$ and for any $\xi\in[-1,0]$, there exists a $k_{0}=k_{0}(\varepsilon,\xi,\omega)\in\mathbb{N}$ such that
$\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{y}(\xi,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))|^{2}\mathrm{d}x\leq\varepsilon,$
for all $k\geq k_{0}$ and $s\leq-\mathscr{T}$. This completes the proof. ∎
Let us now provide the main result of this section.
Theorem 5.5.
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Consider the MDS, $\Im=(\Omega,\mathscr{F},\mathbb{P},\theta)$ from Proposition 3.1, and the RDS $\Psi$ on $\mathbb{H}$ over $\Im$ generated by third-grade fluids equations (2.18) with additive noise satisfying Hypothesis 1.1. Then, there exists a unique random $\mathfrak{DK}$-attractor for continuous RDS $\Psi$ in $\mathbb{H}$.
Proof.
Because of [13, Theorem 2.8], it is only needed to prove that there exists a $\mathfrak{DK}$-absorbing set $\widetilde{\textbf{B}}\in\mathfrak{DK}$ and the RDS $\Psi$ is $\mathfrak{DK}$-asymptotically compact. Existence of $\mathfrak{DK}$-absorbing set $\widetilde{\textbf{B}}\in\mathfrak{DK}$: Let $\mathrm{D}$ be a random set from the class $\mathfrak{DK}$. Let $\kappa_{\mathrm{D}}(\omega)$ be the radius of $\mathrm{D}(\omega)$, that is, $\kappa_{\mathrm{D}}(\omega):=\sup\{\|x\|_{2}:x\in\mathrm{D}(\omega)\}$ for $\omega\in\Omega$.
Let $\omega\in\Omega$ be fixed. Also, let us set
$[\tilde{\kappa}_{11}(\omega)]^{2}$ $=2+2\sup_{t\leq 0}\bigg{\{}\|\boldsymbol{z}(t)\|^{2}_{2}\ e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)t}\bigg{\}}+C_{ubd}\int_{-\infty}^{0}\bigg{\{}\|\boldsymbol{z}(\tau)\|^{2}_{2}+\|\boldsymbol{z}(\tau)\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{f}\|^{2}_{2}\bigg{\}}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\mathrm{d}\tau.$ (5.42)
In view of Lemma 3.6 and Proposition 3.7, we deduce that both $\tilde{\kappa}_{11}\in\mathfrak{K}$ and also that $\tilde{\kappa}_{11}+\kappa_{12}=:\tilde{\kappa}_{13}\in\mathfrak{K}$ as well, where $\kappa_{12}$ is same as in the proof of Theorem 4.3. Therefore the random set $\widetilde{\textbf{B}}$ defined by
$\widetilde{\textbf{B}}(\omega):=\{\boldsymbol{v}\in\mathbb{H}:\|\boldsymbol{v}\|_{2}\leq\tilde{\kappa}_{13}(\omega)\}$
is such that $\widetilde{\textbf{B}}\in\mathfrak{DK}.$
Let us now show that $\widetilde{\textbf{B}}$ absorbs $\mathrm{D}$. For $\omega\in\Omega$, $\boldsymbol{x}\in\mathrm{D}(\theta_{s}\omega)$ and $s\leq-t_{\mathrm{D}}(\omega),$ from (5.2), we have
$\|\boldsymbol{y}(0,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\|_{2}\leq\tilde{\kappa}_{11}(\omega),\ \text{ for }\ \omega\in\Omega.$ (5.43)
Thus, we conclude that, for $\omega\in\Omega$
$\|\boldsymbol{v}(0,\omega;s,\boldsymbol{x})\|_{2}\leq\|\boldsymbol{y}(0,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\|_{2}+\|\boldsymbol{z}(\omega)(0)\|_{2}\leq\tilde{\kappa}_{13}(\omega).$
The above inequality implies that for $\omega\in\Omega$, $\boldsymbol{v}(0,\omega;s,\boldsymbol{x})\in\widetilde{\textbf{B}}(\omega)$, for all $s\leq-t_{\mathrm{D}}(\omega).$ This proves $\widetilde{\textbf{B}}$ absorbs $\mathrm{D}$. The RDS $\Psi$ is $\mathfrak{DK}$-asymptotically compact. In order to establish the $\mathfrak{DK}$-asymptotically compactness of $\Psi$, we use uniform-tail estimates obtained in Lemma 5.4. Let us assume that $\mathrm{D}\in\mathfrak{DK}$ and $\widetilde{\textbf{B}}\in\mathfrak{DK}$ be such that $\widetilde{\textbf{B}}$ absorbs $\mathrm{D}$. Let us fix $\omega\in\Omega$ and take a sequence of positive numbers $\{t_{m}\}^{\infty}_{m=1}$ such that $t_{1}\leq t_{2}\leq t_{3}\leq\cdots$ and $t_{m}\to\infty$. We take an $\mathbb{H}$-valued sequence $\{\boldsymbol{x}_{m}\}^{\infty}_{m=1}$ such that $\boldsymbol{x}_{m}\in\mathrm{D}(\theta_{-t_{m}}\omega),$ for all $m\in\mathbb{N}.$ Our aim is to show that the sequence $\Psi(t_{m},\theta_{-t_{m}}\omega,\boldsymbol{x}_{m})$ or $\boldsymbol{y}(0,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))$ of solutions to system (3.19) has a convergent subsequence in $\mathbb{H}$.
Lemma 5.2 implies that there exists $t_{\mathrm{D}}(\omega)\geq 1$ such that for all $s\leq-t_{\mathrm{D}}(\omega)$ and $\boldsymbol{x}\in\mathrm{D}(\theta_{s}\omega)$,
$\|\boldsymbol{y}(-1,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\|_{2}\leq C,$ (5.44)
where $C$ is a positive constant independent of $s$ and $\boldsymbol{x}.$ Since $t_{m}\to\infty$, there exists $M_{2}=M_{2}(\omega,\mathrm{D})\in\mathbb{N}$ such that $t_{m}\geq t_{\mathrm{D}}(\omega),$ for all $m\geq M_{2}$. Since $\boldsymbol{x}_{m}\in\mathrm{D}(\theta_{-t_{m}}\omega)$, (5.44) implies that for all $m\geq M_{2}$, the sequence
$\{\boldsymbol{y}(-1,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))\}_{m\geq M_{2}}\subset\mathbb{H}$ (5.45)
is bounded in $\mathbb{H}$. We infer from (5.45) and Lemma 5.3 that there exists $\xi\in(-1,0)$, $\hat{\boldsymbol{y}}\in\mathbb{H}$ and a subsequence (not relabeling) such that for every $k\in\mathbb{N}$
$\boldsymbol{y}(\xi,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))=\boldsymbol{y}(\xi,\omega;-1,\boldsymbol{y}(-1,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))-\boldsymbol{z}(-1))\to\hat{\boldsymbol{y}}\ \text{ in }\ \mathbb{L}^{2}(\mathscr{O}_{k}).$ (5.46)
as $m\to\infty$. Therefore, we infer from the proof of Lemma 3.9 that there exists a positive constant $C_{Lip}$ such that
$\|\boldsymbol{y}(0,\omega;\xi,\boldsymbol{y}(\xi,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))-\boldsymbol{z}(\xi))-\boldsymbol{y}(0,\omega;\xi,\hat{\boldsymbol{y}}-\boldsymbol{z}(\xi))\|^{2}_{2}$
$\leq C_{Lip}\|\boldsymbol{y}(\xi,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))-\hat{\boldsymbol{y}}\|^{2}_{2}.$ (5.47)
Let us now choose $\eta>0$ and fix it. Since $\hat{\boldsymbol{y}}\in\mathbb{H}$, there exists $K_{1}=K_{1}(\eta,\omega,\mathrm{D})>0$ such that for all $k\geq K_{1}$,
$\int_{\mathscr{O}\cap\{|x|\geq k\}}|\hat{\boldsymbol{y}}|^{2}\mathrm{d}x<\frac{\eta^{2}}{6C_{Lip}},$ (5.48)
where $C_{Lip}>0$ is a constant defined in (5). Also, we know from Lemma 5.4 that there exists $M_{3}=M_{3}(\xi,\eta,\omega,\mathrm{D})\in\mathbb{N}$ and $K_{2}=K_{2}(\xi,\eta,\omega,\mathrm{D})\geq K_{1}$ such that for all $m\geq M_{3}$ and $k\geq K_{2}$,
$\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{y}(\xi,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))|^{2}\mathrm{d}x<\frac{\eta^{2}}{6C_{Lip}}.$ (5.49)
From (5.46), we have that there exists $M_{4}=M_{4}(\xi,\eta,\omega,\mathrm{D})>M_{3}$ such that for all $m\geq M_{4}$,
$\int_{\mathscr{O}\cap\{|x|<K_{2}\}}|\boldsymbol{y}(\xi,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))-\hat{\boldsymbol{y}}|^{2}\mathrm{d}x<\frac{\eta^{2}}{3C_{Lip}}.$ (5.50)
Finally, we infer from (5) that
$\|\boldsymbol{y}(0,\omega;\xi,\boldsymbol{y}(\xi,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))-\boldsymbol{z}(\xi))-\boldsymbol{y}(0,\omega;\xi,\hat{\boldsymbol{y}}-\boldsymbol{z}(\xi))\|^{2}_{2}$
$\leq C_{Lip}\bigg{[}\int_{\mathscr{O}\cap\{|x|<K_{2}\}}|\boldsymbol{y}(\xi,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))-\hat{\boldsymbol{y}}|^{2}\mathrm{d}x$
$\qquad+\int_{\mathscr{O}\cap\{|x|\geq K_{2}\}}|\boldsymbol{y}(\xi,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))-\hat{\boldsymbol{y}}|^{2}\mathrm{d}x\bigg{]}$
$\leq C_{Lip}\bigg{[}\int_{\mathscr{O}\cap\{|x|<K_{2}\}}|\boldsymbol{y}(\xi,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))-\hat{\boldsymbol{y}}|^{2}\mathrm{d}x$
$\qquad+2\int_{\mathscr{O}\cap\{|x|\geq K_{2}\}}(|\boldsymbol{y}(\xi,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))|^{2}+|\hat{\boldsymbol{y}}|^{2})\mathrm{d}x\bigg{]}$
$<\eta^{2},$ (5.51)
for every $m\geq M_{3}$, where we have used (5.48)-(5.50). Since $\eta>0$ is arbitrary, we conclude the proof. ∎
6. Invariant measures
In this section, we first show the existence of invariant measures for third-grade fluids equations in $\mathbb{H}$. It is established in [27] that the existence of a compact invariant random set is a sufficient condition for the existence of invariant measures, that is, if a RDS $\Psi$ has compact invariant random set, then there exists an invariant measure for $\Psi$ ([27, Corollary 4.4]). Since, the random attractor itself is a compact invariant random set, the existence of invariant measures for the third-grade fluids equations (2.18) is a direct consequence of [27, Corollary 4.4] and Theorem 4.3. The uniqueness of invariant measures is still an open problem.
Let us define the transition operator $\{\mathrm{T}_{t}\}_{t\geq 0}$ by
$\mathrm{T}_{t}f(\boldsymbol{x})=\int_{\Omega}f(\Psi(\omega,t,\boldsymbol{x}))\mathrm{d}\mathbb{P}(\omega)=\mathbb{E}\left[f(\Psi(t,\boldsymbol{x}))\right],$ (6.1)
for all $f\in\mathscr{B}_{b}(\mathbb{H})$, where $\mathscr{B}_{b}(\mathbb{H})$ is the space of all bounded and Borel measurable functions on $\mathbb{H}$ and $\Psi$ is the RDS corresponding to system (2.18), which is defined by (4.1). The continuity of $\Psi$ (cf. Lemma 3.9), [12, Proposition 3.8] provides the following result:
Lemma 6.1.
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Then, the family $\{\mathrm{T}_{t}\}_{t\geq 0}$ is Feller, that is, $\mathrm{T}_{t}f\in\mathrm{C}_{b}(\mathbb{H})$ if $f\in\mathrm{C}_{b}(\mathbb{H})$, where $\mathrm{C}_{b}(\mathbb{H})$ is the space of all bounded and continuous functions on $\mathbb{H}$. Furthermore, for any $f\in\mathrm{C}_{b}(\mathbb{H})$, $\mathrm{T}_{t}f(\boldsymbol{x})\to f(\boldsymbol{x})$ as $t\downarrow 0$.
Analogously as in the proof of [27, Theorem 5.6], one can prove that $\Psi$ is a Markov RDS, that is, $\mathrm{T}_{t_{1}+t_{2}}=\mathrm{T}_{t_{1}}\mathrm{T}_{t_{2}}$, for all $t_{1},t_{2}\geq 0$. Since, we know by [27, Corollary 4.4] that if a Markov RDS on a Polish space has an invariant compact random set, then there exists a Feller invariant probability measure $\upmu$ for $\Psi$.
Definition 6.1.
A Borel probability measure $\upmu$ on $\mathbb{H}$ is called an invariant measure for a Markov semigroup $\{\mathrm{T}_{t}\}_{t\geq 0}$ of Feller operators on $\mathrm{C}_{b}(\mathbb{H})$ if and only if
$\mathrm{T}_{t}^{*}\upmu=\upmu,\ t\geq 0,$
where $(\mathrm{T}_{t}^{*}\upmu)(\Gamma)=\int_{\mathbb{H}}\mathrm{T}_{t}(\boldsymbol{u},\Gamma)\upmu(\mathrm{d}\boldsymbol{u}),$ for $\Gamma\in\mathscr{B}(\mathbb{H})$ and $\mathrm{T}_{t}(\boldsymbol{u},\cdot)$ is the transition probability, $\mathrm{T}_{t}(\boldsymbol{u},\Gamma)=\mathrm{T}_{t}(\chi_{\Gamma})(\boldsymbol{u}),\ \boldsymbol{u}\in\mathbb{H}$.
By the definition of random attractors, it is clear that there exists an invariant compact random set in $\mathbb{H}$. A Feller invariant probability measure for a Markov RDS $\Psi$ on $\mathbb{H}$ is, by definition, an invariant probability measure for the semigroup $\{\mathrm{T}_{t}\}_{t\geq 0}$ defined by (6.1). Hence, we have the following result on the existence of invariant measures for system (2.18) in $\mathbb{H}$.
Theorem 6.2.
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Then, there exists an invariant measure for system (2.18) in $\mathbb{H}$ with external forcing $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$ on bounded and unbounded domains, respectively.
Acknowledgments: “This work is funded by national funds through the FCT – Fundação para a Ciência e a Tecnologia, I.P., under the scope of the projects UIDB/00297/2020 (https://doi.org/10.54499/UIDB/ 00297/2020) and UIDP/00297/2020 (https://doi.org/10.54499/UIDP/00297/2020) (Center for Mathematics and Applications)”. K. Kinra would like to thank Prof. Manil T. Mohan, Department of Mathematics, Indian Institute of Technology Roorkee, Roorkee, India for introducing him into the research area of attractor theory.
Data availability: No data was used for the research described in the article.
Declarations: During the preparation of this work, the authors have not used AI tools.
Conflict of interest: The authors declare no conflict of interest.